### v1

# Universal Hitchin moduli spaces

Abstract.
We study metric aspects of the universal moduli space of solutions to Hitchin’s equations as the complex structure $J$ varies over the Teichmüller space $\mathcal{T}$ of a closed surface $\Sigma$. Our approach is gauge theoretical and builds on the theory of Kähler fibrations and the moment map interpretation of constant scalar curvature Kähler metrics. Our first main result establishes that, over the moduli space of cscK metrics, the universal moduli space of solutions to Hitchin’s equations carries a natural complex structure together with a family of pseudo-Kähler metrics forming a Kähler fibration with a Kähler Ehresmann connection.
We then investigate a second universal moduli space, constructed from the space of flat $G$-connections over $\mathcal{T}$, which admits a nontrivial $J$-dependent Kähler fibration structure discovered by Hitchin. Using symplectic reduction, we build universal moduli spaces of solutions to the harmonicity equations depending on a coupling constant $\alpha$, obtaining natural complex and pseudo-Kähler structures and an explicit Kähler potential. The main novelty here is that this moduli space is defined by a system coupling the scalar curvature with a cubic term in the Higgs field. Finally, we propose a conjectural relationship between the two resulting families of moduli spaces in the weak-coupling limit $\alpha\to 0$, inspired by the twistor geometry of Hitchin’s hyperkähler moduli space.
The first three authors are partially supported by the Spanish Ministry of Science and Innovation, through the ‘Severo Ochoa Programme for Centres of Excellence in R&D’ (CEX2023-001347-S). The first and the third authors are partially supported by MICINN under grant PID2022-141387NB-C21. The second author is partially supported by MICINN under grants PID2022-141387NB-C22 and CNS2022-135784.
1. Introduction
The moduli space $\mathcal{R}$ of representations of the fundamental group of a closed oriented surface $\Sigma$ has an interesting interplay with classical Teichmüller theory. For the case of a compact Lie group $K$, a choice of complex structure $J$ on $\Sigma$ determines an homeomorphism of the moduli space $\mathcal{R}(K)$ of $K$-representations with a moduli space of holomorphic principal $G$-bundles over the compact Riemann surface $X_{J}=(\Sigma,J)$, via the Narasimhan–Seshadri theorem [NS] and its extension to principal bundles by Ramanathan [R1, R2]. Here, $G$ denotes the complex reductive Lie group given by the complexification of $K$. This way, the character variety $\mathcal{R}(K)$ determines a holomorphic fibration over the Teichmüller space $\mathcal{T}$ of $\Sigma$, given by complex structures on $\Sigma$ modulo the action of diffeomorphisms isotopic to the identity.
A different aspect of this interplay arose forty years ago in Nigel Hitchin’s study of the self-duality equations on a Riemann surface $X_{J}$ [Hi1]
(1.1) $\begin{array}[]{l}F_{A}-[\varphi,\tau(\varphi)]=0,\\ \bar{\partial}_{J,A}\varphi=0,\end{array}$
and subsequent work by Simpson [Si3, Si4]. Here $A$ is a connection on a principal $K$-bundle over $\Sigma$ and $\varphi\in\Omega^{1,0}(X_{J},E_{K}(\mathfrak{g}))$ is the Higgs field. The corresponding gauge-theoretical moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$ has three different incarnations, as a character variety $\mathcal{R}_{c}(G)$ for a complex reductive Lie group $G$, as a moduli space space $\mathcal{M}^{\operatorname{Flat}}(G)$ of reductive flat $G$-connections, and as a moduli space of polystable $G$-Higgs bundles $\mathcal{M}(G)$. These three viewpoints on the moduli space of Hitchin’s equations determine a hyperkähler structure on $\mathcal{M}^{\operatorname{Hit}}(G)$. Interestingly, only the moduli space $\mathcal{M}(G)$ depends on the choice of a complex structure on $\Sigma$. Furthermore, the twistor space $Z\to\mathbb{CP}^{1}$ for the hyperkähler structure has generic complex structure $\mathcal{M}^{\operatorname{Flat}}(G)$, while the special points over $\mathbb{CP}^{1}$ correspond to $\mathcal{M}(G)$ and its conjugate.
In this paper we investigate metric aspects of the universal moduli space of solutions of Hitchin’s equations varying over the Teichmüller space $\mathcal{T}$. Our approach is gauge-theoretical in nature, and builds on the classical theory of Kähler fibrations [GLS, M] and in the symplectic interpretation of constant scalar curvature Kähler metrics [Do2, F]. We fix a symplectic structure $\omega$ on $\Sigma$ and consider the coupled system of equations
(1.2) $\begin{split}F_{A}-[\varphi,\tau(\varphi)]&=0,\\ \bar{\partial}_{J,A}\varphi&=0,\\ S_{g}&=\frac{2\pi\chi(\Sigma)}{V},\end{split}$
where $J$ is a complex structure on $\Sigma$ and $S_{g}$ is the scalar curvature of the metric $g=\omega(,J)$. We denote by $\mathcal{U}^{Hit}(G)$ the moduli space of solutions of (1.2) modulo unitary gauge (see Section 5.4). Our first main result can be stated as follows (cf. Theorem 5.18 and Theorem 5.22):
Theorem 1.1.
Let $X=(\Sigma,J)$ be a compact Riemann surface with genus $g(\Sigma)\geqslant 2$. Then, for any fixed total volume $V>0$ and parameter $\varepsilon\in\{-1,1\}$, there exists $\alpha_{0}>0$ such that for any $0<\alpha<\alpha_{0}$ there exists a non-empty open subset
$\mathcal{U}^{*}_{\alpha,\varepsilon}\subset\mathcal{U}^{Hit}(G),$
endowed with a complex structure $\mathbb{I}$ and a pre-symplectic structure $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$. Furthermore, the induced maps $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{Higgs}(G)$ and
(1.3) $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mu_{\curly{H}}^{-1}(0)/\curly{H}$
are holomorphic, where $\mu_{\curly{H}}^{-1}(0)/\curly{H}$ is the moduli space of constant scalar curvature Kähler metrics on $\Sigma$ with total volume $V$.
Furthermore, one has
  1. (1)
    if $\varepsilon=1$ the tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is possibly degenerate,
  2. (2)
    if $\varepsilon=-1$ the tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is non-degenerate, and defines a pseudo-Kähler structure on the moduli space.
In either case, the restriction of $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}=\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}(,\mathbb{I})$ to the fibres of (5.29) is $\alpha\mathbf{g}$, where $\mathbf{g}$ denotes the hyperkähler metric on the moduli space of solutions of Hitchin’s equations. Consequently, (5.29) has a natural structure of Kähler fibration with coupling form $\boldsymbol{\omega}^{\mathbb{I}}_{1,0}$ (see (5.22)) and Kähler Ehresmann connection.
We comment briefly on the proof. The moduli space $\mathcal{U}^{Hit}(G)$ of solutions of (1.2) modulo gauge is defined via symplectic reduction, and hence is naturally endowed with a family of (pre)-symplectic structures
(1.4) $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}=\varepsilon\boldsymbol{\omega}{J}+\alpha\boldsymbol{\sigma}^{\mathbb{I}}.$
Here $\boldsymbol{\sigma}^{\mathbb{I}}$ is a closed $2$-form restricting on the fibres to Hitchin’s symplectic structure $\omega_{\mathbf{I}}$ (see Proposition 5.6) and $\boldsymbol{\omega}{J}$ is the pull-back of the symplectic structure on the moduli space $\mu_{\curly{H}}^{-1}(0)/\curly{H}$, induced by the choice of symplectic form $\omega$ on $\Sigma$ (see (3.16)). Even though the phase space of parameters $(J,A,\psi)$, with $\psi=-i(\varphi-\tau_{h}\varphi)$, has a natural complex structure $\mathbb{I}$, the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}=\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}(,\mathbb{I})$ is not positive definite (see Lemma 3.8). This follows from an explicit formula (see Corollary 5.19)
(1.5) $\begin{split}\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a\wedge Ja)+\alpha\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{\alpha}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right),\end{split}$
where $B$ is a positive-definite invariant metric on $\mathfrak{k}=\operatorname{Lie}K$. Hence, it is not obvious a priori that $\mathcal{U}^{Hit}(G)$ inherits a complex structure compatible with $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$. The key step is to undertake a gauge fixing for solutions of the coupled Hitchin’s equations (1.2), whereby the complex structure $\mathbb{I}$ and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ descend to the moduli space. Difficulties will arise, due to the fact that $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is neither a definite pairing nor non-degenerate.
Our second main result is concerned with a universal moduli space of flat $G$-connections over $\mathcal{T}$. Even though the space of flat $G$-connections is independent of any choice of complex structure $J$ on $\Sigma$, its product with the space of complex structures on $\Sigma$ carries a non-trivial, $J$-dependent structure of Kähler fibration $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$ discovered by Hitchin [Hi1]. Building on this observation, we fix a symplectic form $\omega$ on $\Sigma$, and for any choice of coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, we consider the coupled harmonic equations
(1.6) $\begin{split}F_{A}-\frac{1}{2}[\psi,\psi]&=0,\\ d_{A}\psi&=0,\\ d_{A}^{\ast}\psi&=0,\\ S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi\right)\right)&=\frac{2\pi\chi(\Sigma)}{V},\end{split}$
where $g=\omega(,J)$ and * is the corresponding Hodge star operator. We denote by $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ the moduli space of solutions of (1.6) modulo unitary gauge (see Section 4.2). We denote by $\mathcal{U}^{\operatorname{Flat}}(G)$ the universal moduli space of reductive flat $G$-connections over $\mathcal{T}$ (see Section 4.1). Our second main result can be stated as follows (cf. Theorem 4.10 and Theorem 4.14):
Theorem 1.2.
Let $X=(\Sigma,J)$ be a compact Riemann surface with genus $g(\Sigma)\geqslant 2$. Then, for any fixed total volume $V>0$ and parameter $\varepsilon\in\{-1,1\}$, there exists $\alpha_{0}>0$ such that for any $0<\alpha<\alpha_{0}$ there exists a non-empty open subset
$\mathcal{U}^{*}_{\alpha,\varepsilon}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon},$
endowed with a complex structure $\mathbb{J}$ and a pre-symplectic structure $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$. Furthermore, the induced map
(1.7) $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)$
is holomorphic, and
  1. (1)
    if $\varepsilon=1$ the tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}=\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(,\mathbb{J})$ is possibly degenerate,
  2. (2)
    if $\varepsilon=-1$ the tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is non-degenerate, and defines a pseudo-Kähler structure on the moduli space.
In either case, $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ admits a global Kähler potential, that is, $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}=dd^{c}_{\mathbb{J}}\Phi$, where
$\Phi=\varepsilon\nu{J}+\frac{\alpha}{2}\|\psi\|^{2}_{L^{2}}$
and $\nu{J}$ is induced by the global Kähler potential in the space of complex structures compatible with the orientation $\curly{J}$ (see [F, Section 4]).
Similarly as for the universal moduli space of Hitchin’s equations (1.2), the (possibly degenerate) moduli space Kähler form and metric admit explicit formulae upon suitable gauge fixing
(1.8) $\begin{split}\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2}))\\ &+\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge J(\psi(\dot{J}_{2}))).\end{split}$
From the previous expression, we observe that, at least formally, the restriction of $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ to the fibres over $\mathcal{T}$ coincides up to scaling with Hitchin’s symplectic structure $\omega_{\mathbf{J}}$. A complete proof of this requires a better understanding of the particular gauge fixing mechanism which we use to construct the moduli space complex structure $\mathbb{J}$ (see Proposition 4.9), and we leave it as an open question.
Based on the twistor space structure for Hitchin’s original hyperkähler moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$, it is natural to speculate on a relation between the moduli spaces $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ and $\mathcal{U}^{Higgs}(G)$ in the weak coupling limit
$\alpha\to 0.$
In this limit, for instance, the equations (1.6) reduce to the coupled Hitchin equations (1.2), and we expect a suitable adiabatic limit convergence
$\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}\to\mathbf{g}^{\mathbb{I}}_{\tilde{\alpha},\varepsilon}.$
We are far from even grasping the solution of this proposal, and we leave it as an open question for future studies.
Motivation for the present work comes from the programme initiated by the first three authors in [AGG] more than 15 years ago based on the 2009 PhD Thesis of the second author [GF]. In that paper, we consider a pair $(X,E)$ consisting of a holomorphic principal $G$-bundle $E$ over a compact complex manifold $X$, where $G$ is a complex reductive group, and study certain coupled equations for a Kähler metric on $X$ and a reduction of structure group of $E$ to a maximal compact subgroup $K\subset G$. These equations, that we named Kähler–Yang–Mills equations, appear naturally as moment map equations for the action of the extended gauge group of the $K$-bundle $E_{K}$ over the underlaying smooth manifold to $X$ equipped with a symplectic form $\omega$. The extension of the gauge group of $E_{K}$ is given in this case by the group of Hamiltonian symplectomorphisms defined by $\omega$. This theory is in some sense a combination of the usual Yang–Mills theory and the Donaldson–Fujiki theory on Kähler manifolds, interpreting the constant scalar curvature as the moment map for the action of the group of Hamiltonian symplectomorphisms on the space of complex structures compatible with the symplectic structure $\omega$.
While much work remains to be done in this programme, like for example finding the algebraic general stability condition solving the Kähler–Yang–Mills equations, certain particular cases with symmetry have been understood using dimensional reduction methods [AGG2, AGGP] and symplectic reduction in stages [AGGPY]. A generalization of the Kähler–Yang–Mills equations involving Higgs fields of different kinds has been introduced in [AGG3] and remains also to be fully explored.
In early joint work with the fourth author, we extended our symplectic point of view to the problem of constructing universal moduli spaces for Hitchin’s equations and Higgs bundles. The first steps of this attempt appeared in Chapter 7 of the fourth author’s 2018 ETH PhD Thesis [T], and is the seed of the present work. In view of the current interest in the field, after a dormant period, we decided to resume this investigation. We should warn that the use of the term ‘universal’ that we make here does not necessarily imply some expected functoriality properties that such term some times entails. We have sticked, however, to this term since that is the term that we originally used in our initial work [T].
After the completion of this work, we received a copy of a manuscript [CTW], where related topics are treated. It should be very interesting to compare our approach to other analytic constructions under study [CTW, Hi4], as well as the more algebraic point of view taken by Simpson [Si2, Si3] and others [BBN, BDP, D, DF].
Acknowledgements.
The authors wish to thank Emilio Franco, Nigel Hitchin and Ignasi Mundet for helpful conversations. Special thanks are due to Richard Wentworth for stimulating discussions – in particular for spotting a mistake in the signature of the metric $\mathbf{g}^{\mathbb{I}}_{\alpha,1}$ in an earlier version of this manuscript – as well as for his lectures at ICMAT in June 2025, which encouraged us to resume this project.
2. Flat connections and Higgs bundles
2.1. Flat $\boldsymbol{G}$-connections
Through this section we consider a connected semisimple complex Lie group $G$, with Lie algebra $\mathfrak{g}$, and fix an antiholomorphic involution $\tau$ of $G$ defining a maximal compact subgroup $K:=G^{\tau}\subset G$, with Lie algebra $\mathfrak{k}$. The Killing form will be denoted by $B$. We also fix a smooth oriented compact surface $\Sigma$.
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$. The space $\curly D$} of connections on $E_{G}$ is an affine space modelled on the complex vector space $\Omega^{1}(\Sigma,E_{G}(\mathfrak{g}))$, where $E(\mathfrak{g})$ is the adjoint bundle associated to $E$ via the adjoint representation of $G$ in $\mathfrak{g}$. Hence, $\curly D$} has a natural integrable complex structure $\mathbf{J}$, defined by
(2.1) $\mathbf{J}(\dot{D})=i\dot{D},\quad\mbox{for}\;\;D\in{\curly D}\;\;\mbox{and}\;\;\dot{D}\in T_{D}{\curly D}=\Omega^{1}(\Sigma,E_{G}(\mathfrak{g})),$
induced by the complex structure of $\mathfrak{g}$ and compatible with the affine structure.
The space $\curly D$} is furthermore holomorphic symplectic, with holomorphic symplectic structure, defined by
(2.2) $\Omega_{\mathbf{J}}(\dot{D}_{1},\dot{D}_{2})=\int_{X}B(\dot{D}_{1}\wedge\dot{D}_{2}),$
which is preserved by the action of the gauge group $\curly{G}$ of $E_{G}$. We have that $\operatorname{Lie}\curly{G}=\Omega^{0}(\Sigma,E_{G}(\mathfrak{g}))$ and hence its dual $(\operatorname{Lie}\curly{G})^{\ast}$ can be identified with $\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))$. Similarly as for the case of unitary connections, studied by Atiyah–Bott [AB], there existes an equivariant complex moment map for the action of $\curly{G}$ on $\curly D$} given by
(2.3) $\mu{G}\colon{\curly D}$ $\longrightarrow$ $\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))$
(2.4) $D$ $\longmapsto$ $F_{D},$
where $F_{D}$ is the curvature of $D$. The zero moment map condition, that is when $F_{D}=0$, corresponds to the flatness of $D$. In this case one refers to the pair $(E_{G},D)$ as a flat $G$-bundle.
In order to construct a Hausdorff moduli space of flat $G$-connections modulo gauge, one needs to impose a natural stability condition arising from Geometric Invariant Theory. We say that a bundle with $G$-connection $(E_{G},D)$ is reductive (also called completely reducible) if it splits as a product of irreducible sub-bundles, where irreducible (also called simple or stable in the literature) means that there exists no nontrivial $D$-invariant sub-bundle. For a flat $G$-bundle $(E_{G},D)$, this condition can be expressed in terms of the holonomy representation: the holonomy of $D$ defines a representation
$\rho:\pi_{1}(\Sigma)\to G,$
of the fundamental group $\pi_{1}(\Sigma)$ of $\Sigma$. A representation $\rho:\pi_{1}(X)\to G$ is said to be reductive if the Zariski closure of the image of $\rho$ is a reductive group, or equivalently if $\operatorname{ad}_{G}\circ\rho$ is completely reducible, where $\operatorname{ad}_{G}:G\to\mathrm{GL}(\mathfrak{g})$ is the adjoint representation. Then, one can prove that the flat connection $D$ on $E_{G}$ is reductive if and only if its corresponding representation is reductive.
The existence of the moment map $\mu{G}$ implies that the flatness condition is invariant under the action of $\curly{G}$ on $\curly D$}, and so is the reductiveness condition, and we can thus consider the moduli space
$\mathcal{M}^{\operatorname{Flat}}(G):=\{\mbox{reductive}\;\;D\in{\curly D}\;\;\mbox{with}\;\;F_{D}=0\}/\curly{G}.$
The following important result, which we state informally, can be found in the work of Goldman [Go].
Theorem 2.1.
The moduli space of flat $G$-connections $\mathcal{M}^{\operatorname{Flat}}(G)$ on $E_{G}$ admits a natural structure of singular complex manifold, and a holomorphic symplectic structure induced by $\Omega_{\mathbf{J}}$ on its smooth locus.
2.2. Surface group representations
The complex structure on the moduli space $\mathcal{M}^{\operatorname{Flat}}(G)$ can be understood explicitly using $G$-representations of the fundamental group of the surface $\Sigma$ [AB, Hi1] (see Remark 2.3). Let $\operatorname{Hom}(\pi_{1}(\Sigma),G)$ be the set of representations of $\pi_{1}(\Sigma)$ in $G$. Since
$\pi_{1}(\Sigma)=\langle a_{1},b_{1},\dotsc,a_{g},b_{g}\;\;:\;\;\prod_{i=1}^{g}[a_{i},b_{i}]=1\rangle$
$\operatorname{Hom}(\pi_{1}(\Sigma),G)$ can be naturally identified with the subset of $G^{2g}$ consisting of $2g$-tuples $(A_{1},B_{1}\dotsc,A_{g},B_{g})$ satisfying the algebraic equation $\prod_{i=1}^{g}[A_{i},B_{i}]=1$, and hence has a natural structure as a complex algebraic variety.
The moduli space of representations of $\pi_{1}(\Sigma)$ in $G$, or $G$-character variety of $\pi_{1}(\Sigma)$ is defined as the quotient
$\mathcal{R}(G):=\operatorname{Hom}^{+}(\pi_{1}(\Sigma),G)/G,$
where $\operatorname{Hom}^{+}(\pi_{1}(\Sigma),G)\subset\operatorname{Hom}(\pi_{1}(\Sigma),G)$ is the subvariety of reductive representations, and $G$ acts by conjugation. This quotient coincides with the GIT quotient and hence
$\mathcal{R}(G)=\operatorname{Hom}(\pi_{1}(\Sigma),G)\sslash G$
is a complex algebraic variety.
Given a representation $\rho\colon\pi_{1}(\Sigma)\to G$, consider the associated flat $G$-bundle on $\Sigma$, defined as $E_{\rho}=\tilde{\Sigma}\times_{\rho}G$, where $\tilde{\Sigma}\to\Sigma$ is the universal cover and $\pi_{1}(\Sigma)$ acts on $G$ via $\rho$. This gives in fact an identification between the set of equivalence classes of representations $\operatorname{Hom}(\pi_{1}(\Sigma),G)/G$ and the set of equivalence classes of flat $G$-bundles, which in turn is parametrized by the cohomology set $H^{1}(X,G)$. We can then assign a topological invariant to a representation $\rho$ given by the characteristic class $c(\rho):=c(E_{\rho})\in\pi_{1}(G)$ corresponding to $E_{\rho}$. To define this, let $\widetilde{G}$ be the universal covering group of $G$. We have an exact sequence
$1\longrightarrow\pi_{1}(G)\longrightarrow\widetilde{G}\longrightarrow G\longrightarrow 1$
which gives rise to the (pointed sets) cohomology sequence
(2.5) $H^{1}(\Sigma,{\widetilde{G}})\longrightarrow H^{1}(\Sigma,{G})\stackrel{{\scriptstyle c}}{{\longrightarrow}}H^{2}(\Sigma,\pi_{1}(G)).$
Since $\pi_{1}(G)$ is abelian, we have
$H^{2}(\Sigma,\pi_{1}(G))\cong\pi_{1}(G),$
and $c(E_{\rho})$ is defined as the image of $E$ under the last map in (2.5). Thus the class $c(E_{\rho})$ measures the obstruction to lifting $E_{\rho}$ to a flat $\widetilde{G}-$bundle, and hence to lifting $\rho$ to a representation of $\pi_{1}(\Sigma)$ in $\widetilde{G}$.
For a fixed $c\in\pi_{1}(G)$, the moduli space of reductive representations $\mathcal{R}_{c}(G)$ with topological invariant $c$ is defined as the subvariety
(2.6) $\mathcal{R}_{c}(G):=\{\rho\in\mathcal{R}(G)\;\;:\;\;c(\rho)=c\}.$
Proposition 2.2.
Let $E_{G}$ be a smooth principal $G$-bundle over $X$ with topological class $c$. Then there is a complex analytic isomorphism
$\mathcal{M}^{\operatorname{Flat}}(G)\cong{\mathcal{R}}_{c}(G).$
Remark 2.3.
We warn the reader that, even though there is a complex analytic isomorphism between the moduli space of flat $G$-connections $\mathcal{M}^{\operatorname{Flat}}(G)$ and the $G$-character variety with fixed topological invariant $c$, these spaces carry very different algebraic structures (see e.g. [Si3]).
2.3. Harmonicity equations and hyperkähler structure
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$. Let $h\in\Omega^{0}(E_{G}(G/K))$ be a reduction of structure group of $E_{G}$ to $K$, and let $E_{K}$ be the corresponding principal $K$-bundle. From the decomposition $\mathfrak{g}=\mathfrak{k}\oplus i\mathfrak{k}$ one has that any connection $D\in{\curly D}$ on $E_{G}$ admits a decomposition
$D=A+i\psi$
where $A$ is a connection on $E_{K}$ and $\psi\in\Omega^{1}(X,E_{K}(\mathfrak{k}))$. This simple fact induces a one-to-one correspondence
(2.7) ${\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ $\longrightarrow$ ${\curly D}$
$(A,\psi)$ $\longmapsto$ $D:=A+i\psi$
where $\curly A$} denotes the space of connections on $E_{K}$. As observed by Hitchin [Hi1], the previous isomorphism naturally identifies the space of $G$-connections $\curly D$} with the cotangent space to $\curly A$}, suggesting a relation to hyperkähler geometry.
To recall this relation, we fix a complex structure $J$ on $\Sigma$ compatible with the orientation, and denote by $X=(\Sigma,J)$ the corresponding Riemann surface. Via $J$, the space $\curly A$} inherits an integrable complex structure [AB], and hence so does $\curly D$} via its identification with $T^{*}{\curly A}$ above. Explicitly, this is given by
$\mathbf{I}(a,\dot{\psi})=(Ja,-J\dot{\psi})$
where Ja=a(J)Ja=-a(J). Let us denote $\mathbf{K}:=\mathbf{I}\mathbf{J}$. Then, one can check that $\mathbf{K}$ defines a new complex structure and furthermore that the three complex structures
(2.8) $\mathbf{I}(a,\dot{\psi})$ $=(Ja,-J\dot{\psi})$
(2.9) $\mathbf{J}(a,\dot{\psi})$ $=(-\dot{\psi},a)$
(2.10) $\mathbf{K}(a,\dot{\psi})$ $=(-J\dot{\psi},-Ja),$
satisfy the quaternion relations, and define a flat hyperkähler structure on ${\curly D}\cong T^{*}{\curly A}$ with metric tensor
(2.11) $\mathbf{g}((a,\dot{\psi}),(a,\dot{\psi}))=\int_{\Sigma}B(a\wedge Ja)+\int_{\Sigma}B(\dot{\psi}\wedge J\dot{\psi}).$
The symplectic structures $\omega_{\mathbf{I}}=\mathbf{g}(\mathbf{I},)$ and $\omega_{\mathbf{K}}=\mathbf{g}(\mathbf{K},)$ combine to give the $\mathbf{J}$-holomorphic symplectic structure studied in Section 2.1
(2.12) $\Omega_{\mathbf{J}}=\omega_{\mathbf{I}}+i\omega_{\mathbf{K}}.$
A Hamiltonian action of the gauge group $\curly K$} of $E_{K}$ for the symplectic structure $\omega_{\mathbf{J}}=\mathbf{g}(\mathbf{J},)$, given by
(2.13) $\omega_{\mathbf{J}}((a_{1},\dot{\psi}_{1}),(a_{2},\dot{\psi}_{2}))=\int_{\Sigma}B(a_{1}\wedge J\dot{\psi}_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge Ja_{2}),$
was studied by Corlette [C].
Proposition 2.4.
The action of $\curly K$} on $({\curly D},\omega_{\mathbf{J}})$ is Hamiltonian, with equivariant moment map given by
$\langle\mu_{{\curly K}}^{\mathbf{J}}(D),\zeta\rangle=\int_{\Sigma}B(\zeta,d_{A}(J\psi)),$
where $\zeta\in\operatorname{Lie}{\curly K}=\Omega^{0}(\Sigma,E_{K}(\mathfrak{k}))$.
The moment maps $\mu{G}=\mu_{{\curly K}}^{\mathbf{I}}+i\mu_{{\curly K}}^{\mathbf{K}}$ and $\mu_{{\curly K}}^{\mathbf{J}}$ combine to give a hyperkähler moment map for the $\curly K$}-action on $\curly D$}
${{\boldsymbol{\mu}}}_{\curly K}=(\mu_{{\curly K}}^{\mathbf{I}},\mu_{{\curly K}}^{\mathbf{J}},\mu_{{\curly K}}^{\mathbf{K}})$
whose zero locus corresponds to solutions of the harmonicity equations
(2.14) $\begin{array}[]{l}F_{A}-\frac{1}{2}[\psi,\psi]=0,\\ d_{A}\psi=0,\\ d_{A}^{\ast}\psi=0.\end{array}$
Note that the first two are equivalent to the flatness condition for $D$, via the formula
$F_{D}=F_{A}-\frac{1}{2}[\psi,\psi]+id_{A}\psi,$
while the last equation is equivalent to the vanishing of the moment map $\mu_{{\curly K}}^{\mathbf{J}}$, due to the standard identity $\ast\psi=J\psi$.
The harmonicity condition for the reduction $h$ admits an interpretation in terms of the the holonomy representation $\rho:\pi_{1}(X)\to G$ corresponding to $D$. The reduction $h$ is equivalent to a $\pi_{1}(X)$-equivariant smooth map
$\tilde{h}:\tilde{X}\longrightarrow G/K,$
where $\tilde{X}$ is the universal covering of $X$. Here $\pi_{1}(X)$ acts on $\tilde{X}$ by Deck transformations, and via the representation $\rho$ on $G/K$. The symmetric space $G/K$ is equipped with a canonical $G$-invariant Riemannian metric determined by the Killing form $B$, and (2.14) is equivalent to the harmonicity of $\tilde{h}$ in the usual sense.
The equations (2.14) are invariant under the action of $\curly K$} and hence we can consider the moduli space
[…]
given by the infinite-dimensional hyperkähler reduction ${{\boldsymbol{\mu}}}_{\curly K}^{-1}(0)/{\curly K}$.
The following theorem, due to Donaldson [Do] for $G=\mathrm{SL}(2,\mathbb{C})$, and Corlette [C] in general, is one of the main structural results of the theory.
Theorem 2.5.
The map (2.7) induces an homeomorphism
$\mathcal{M}^{\operatorname{Harm}}(G)\cong\mathcal{M}^{\operatorname{Flat}}(G).$
Furthermore, $\mathcal{M}^{\operatorname{Harm}}(G)$ carries a natural hyperkähler structure on its smooth locus, whose generic complex structure is biholomorphic to $\mathcal{M}^{\operatorname{Flat}}(G)$.
In particular, the previous theorem states that a solution of the harmonicity equations (2.14) has an associated reductive flat connection $D=A+i\psi$. Conversely, any pair $(A,\psi)$ such that $D=A+i\psi$ is a reductive flat connection, admits an element on its $\curly{G}$-orbit, unique up to the $\curly K$}-action, solving (2.14). The non-generic complex structures on $\mathcal{M}^{\operatorname{Harm}}(G)$ correspond to the moduli space of $G$-Higgs bundles over $X$, induced by complex structure $\mathbf{I}$, and its conjugate, as we discuss in the next sections.
2.4. $\boldsymbol{G}$-Higgs bundles and Hitchin equations
Through this section we consider a connected semisimple complex Lie group $G$, with Lie algebra $\mathfrak{g}$, and fix an antiholomorphic involution $\tau$ of $G$ defining a maximal compact subgroup $K:=G^{\tau}\subset G$, with Lie algebra $\mathfrak{k}$. The Killing form will be denoted by $B$.
We also consider a compact Riemann surface $X$ with canonical line bundle $K_{X}$.
A $G$-Higgs bundle over $X$ is a pair $(E,\varphi)$ consisting of a holomorphic principal $G$-bundle $E$ over $X$ and an element $\varphi\in H^{0}(X,E(\mathfrak{g})\otimes K_{X})$, that is a holomorphic section of $E(\mathfrak{g})\otimes K_{X}$, where $E(\mathfrak{g})$ is the adjoint bundle associated to $E$ via the adjoint representation of $G$ in $\mathfrak{g}$.
There are appropriate notions of stability, semistability and polystability and we can consider $\mathcal{M}(G)$ to be the moduli space of isomorphism classes polystable $G$-Higgs bundles over $X$.
Let $h\in\Omega^{0}E(X,G/K)$ be a smooth reduction of structure group of $E$ to $K$ and let $E_{h}$ be the corresponding smooth principal $K$-bundle. Let $F_{h}\in\Omega^{2}(X,E_{h}(\mathfrak{k}))$ be the curvature of the unique connection compatible with $h$ and the holomorphic structure on $E$ (defined by the Chern–Singer correspondence). Here $E_{h}(\mathfrak{k})$ is the adjoint bundle of $E_{h}$.
Abusing notation, let
$\tau\colon\Omega^{1,0}(X,E(\mathfrak{g}))\to\Omega^{0,1}(X,E(\mathfrak{g}))$
be the conjugation defined by $h$ combined with the conjugation on complex $1$-forms on $X$. The Higgs field $\varphi$ can be viewed as a $(1,0)$-form $\varphi\in\Omega^{1,0}(X,E(\mathfrak{g}))$, then $\tau(\varphi)\in\Omega^{0,1}(X,E(\mathfrak{g}))$, and hence $\varphi,\tau(\varphi)]\in\Omega^{2}(X,E_{h}(\mathfrak{k}))$.
The proof of the following theorem is due to Hitchin [Hi1] for $G=\mathrm{SL}(2,\mathbb{C})$ and Simpson [Si1] for arbitrary $G$.
Theorem 2.6.
A reduction $h$ of structure group of $E$ to $K$ satisfies the Hitchin equation
$F_{h}-[\varphi,\tau(\varphi)]=0$
if and only if $(E,\varphi)$ is polystable.
From the point of view of moduli spaces it is convenient to fix a $C^{\infty}$ principal $K$-bundle $E_{K}$ and study the moduli space of solutions to Hitchin’s equations for a pair $(A,\varphi)$ consisting of a connection $A$ on $E_{K}$ and $\varphi\in\Omega^{1,0}(X,E_{K}(\mathfrak{g}))$:
(2.15) $\begin{array}[]{l}F_{A}-[\varphi,\tau(\varphi)]=0\\ \bar{\partial}_{A}\varphi=0.\end{array}$
Here $\bar{\partial}_{A}$ is the $(0,1)$ part of the covariant derivative $d_{A}$ defined by $A$. This is of course the Dolbeault operator defined by the holomorphic structure $J_{A}$ on $E_{G}$ corresponding to $A$ via de Chern–Singer correspondence. Here $E_{G}$ is the smooth principal $G$-bundle obtained frm $E_{K}$ by extension of structure group. The gauge group $\curly K$} of automorphisms of $E_{K}$ acts on the space of solutions and the moduli space of solutions is
[…]
Now, from Theorem 2.6 one has the following.
Theorem 2.7.
There is a homeomorphism
$\mathcal{M}(G)\cong\mathcal{M}^{\operatorname{Hit}}(G)$
To explain this correspondence we interpret the moduli space of $G$-Higgs bundles in terms of pairs $(J_{E},\varphi)$ consisting of a holomorphic structure $J_{E}$ on the smooth $G$-bundle $E_{G}$ obtained from $E_{K}$ by the extension of structure group, and $\varphi\in\Omega^{1,0}(X,E_{G}(\mathfrak{g}))$ satisfying $\bar{\partial}_{E}\varphi=0$. Such pairs are in correspondence with $G$-Higgs bundles $(E,\varphi)$, where $E$ is the holomorphic $G$-bundle defined by $J_{E}$ on $E_{G}$, and $\bar{\partial}_{E}\varphi=0$, that is $\varphi\in H^{0}(X,E(\mathfrak{g})\otimes K_{X})$. The moduli space of polystable $G$-Higgs bundles $\mathcal{M}(G)$ can now be identified with the orbit space
$\{(J_{E},\varphi)\;\;\mbox{with}\;\;\bar{\partial}_{E}\varphi=0\;\;\mbox{such that $(E,\varphi)$ is polystable}\}/\curly{G},$
where $\curly{G}$ is the gauge group of automorphisms of $E_{G}$, which is in fact the complexification of $\curly K$}. Since, by the Chern–Singer correspondence, there is a bijection between connections on $E_{K}$ and holomorphic structures on $E_{G}$, the correspondence given in Theorem 2.7 can be interpreted by saying that in the $\curly{G}$-orbit of a polystable $G$-Higgs bundle $(J_{E_{0}},\varphi_{0})$ one can find another Higgs bundle $(J_{E},\varphi)$ whose corresponding pair $(A,\varphi)$ satisfies $F_{A}-[\varphi,\tau(\varphi)]=0$, and this is unique up to gauge transformations in $\curly K$}.
2.5. Hitchin equations and hyperkähler structure
Coming back to the setup of Section 2.4, let $E_{K}$ be a smooth principal $K$-bundle over $X$, and let $E_{G}$ the principal $G$-bundle obtained by extension of structure group.
The space $\curly A$} of connections on $E_{K}$ is an affine space modelled on $\Omega^{1}(X,E_{K}(\mathfrak{k}))$, which is equipped with a symplectic structure defined by
$\omega_{\curly A}(a,b)=\int_{X}B(a\wedge b),\;\;\;\mbox{for}\;\;A\in{\curly A}\;\;\mbox{and}\;\;a,b\in T_{A}{\curly A}=\Omega^{1}(X,E_{K}(\mathfrak{k})).$
This is obviously closed since it is independent of $A\in{\curly A}$.
Now, the set $\curly C$} of holomorphic structures on $E_{G}$ is an affine space modelled on $\Omega^{0,1}(X,E_{G}(\mathfrak{g}))$, and it has a complex structure $J_{\curly C}$, induced by the complex structure of the Riemann surface, which is defined by
$J_{\curly C}(\alpha)=i\alpha,\;\;\;\mbox{for}\;\;J_{E}\in{\curly C}\;\;\mbox{and}\;\;\alpha\in T_{J_{E}}{\curly C}=\Omega^{0,1}(X,E_{G}(\mathfrak{g})).$
The Chern–Singer correspondence [Sin, W] establishes an isomorphism
(2.16) ${\curly A}$ $\longrightarrow$ ${\curly C}$
(2.17) $A$ $\longmapsto$ $J_{A}$
The corresponding tangent spaces are in bijection under the map
(2.18) $\Omega^{0,1}(X,E_{G}(\mathfrak{g}))$ $\longrightarrow$ $\Omega^{1}(X,E_{K}(\mathfrak{k}))$
(2.19) $\alpha$ $\longmapsto$ $a:=\alpha-\tau(\alpha)$
Under this identification $J_{\curly C}$ defines a complex structure $J_{\curly A}$ on $\curly A$}.
The symplectic structure $\omega_{\curly A}$ and the complex structure $J_{\curly A}$ define a Kähler structure on $\curly A$}, which is preserved by the action of the gauge group $\curly K$} of $E_{K}$. We have that $\operatorname{Lie}{\curly K}=\Omega^{0}(X,E_{K}(\mathfrak{k}))$ and hence its dual $(\operatorname{Lie}{\curly K})^{\ast}$ can be identified with $\Omega^{2}(X,E_{K}(\mathfrak{k}))$. By Atiyah–Bott [AB] one has that the moment map for the action of $\curly K$} on $\curly A$} is given by
(2.20) ${\curly A}$ $\longrightarrow$ $\Omega^{2}(X,E_{K}(\mathfrak{k}))$
(2.21) $A$ $\longmapsto$ $F_{A}.$
Now, let us denote $\Omega=\Omega^{1,0}(X,E_{G}(\mathfrak{g}))$. The linear space $\Omega$ has a natural complex structure $J_{\Omega}$ defined by multiplication by $i$, and a symplectic structure given by
$\omega_{\Omega}(\eta,\nu)=i\int_{X}B(\eta\wedge\nu^{\ast}),\;\;\;\mbox{for}\;\;\varphi\in\Omega\;\;\mbox{and}\;\;\eta,\nu\in T_{\varphi}\Omega=\Omega.$
We can now consider ${\curly A}\times\Omega$ with the symplectic structure $\omega_{\curly A}+\omega_{\Omega}$ and complex structure $J_{\curly A}+J_{\Omega}$. The action of $\curly K$} on ${\curly A}\times\Omega$ preserves these symplectic and complex structures and there is a moment map given by ([Hi1])
(2.22) $\begin{array}[]{ccc}{\curly A}\times\Omega&\longrightarrow&\Omega^{2}(X,E_{K}(\mathfrak{k}))\\ (A,\varphi)&\mapsto&F_{A}-[\varphi,\tau(\varphi)].\end{array}$
Let us denote $J_{1}:=J_{\curly A}+J_{\Omega}$. Via the identification ${\curly A}\cong{\curly C}$, we have for $\alpha\in\Omega^{0,1}(X,E_{G}(\mathfrak{g}))$ and $\eta\in\Omega^{1,0}(X,E_{G}(\mathfrak{g}))$ the following three complex structures on ${\curly A}\times\Omega$:
$\begin{array}[]{ccc}J_{1}(\alpha,\eta)&=&(i\alpha,i\eta)\\ J_{2}(\alpha,\eta)&=&(-i\tau(\eta),i\tau(\alpha))\\ J_{3}(\alpha,\eta)&=&(\tau(\eta),-\tau(\alpha)),\end{array}$
These complex structures correspond to the complex structures $\mathbf{I},\mathbf{J}$ and $\mathbf{K}$ in (2.8) via the the usual Chern–Singer correspondence.
Consequently, $J_{i}$, $i=1,2,3$, satisfy the quaternion relations, and define a hyperkähler structure on ${\curly A}\times\Omega$, with symplectic structures $\omega_{i}$, $i=1,2,3$, where $\omega_{1}=\omega_{\curly A}+\omega_{\Omega}$. The symplectic structures $\omega_{2}$ and $\omega_{3}$ combine to define the $J_{1}$-holomorphic symplectic structure $\omega_{c}:=\omega_{2}+i\omega_{3}$, given by
$\omega_{c}((a,\eta),(b,\nu))=\int_{X}B(\eta\wedge\beta-\nu\wedge\alpha),$
for $(a,\eta),(b,\nu)\in T_{A,\varphi)}({\curly A}\times\Omega)\cong\Omega^{1}(X,E_{K}(\mathfrak{k}))\times\Omega^{1,0}(X,E_{G}(\mathfrak{g}))$, where $\alpha$ and $\beta$ are the elements in $\Omega^{0,1}(X,E(\mathfrak{g})$ corresponding to $a$ and $b$ respectively under the identification (2.18).
The action of the gauge group $\curly K$} on ${\curly A}\times\Omega$ preserves the hyperkähler structure and there are moment maps given by
$\mu_{1}(A,\varphi)=F_{A}-[\varphi,\tau(\varphi)],\;\;\;\mu_{2}(A,\varphi)=\mathop{{Re}}\nolimits(\bar{\partial}_{E}\varphi),\;\;\;\mu_{3}(A,\varphi)=\mathop{{Im}}\nolimits(\bar{\partial}_{E}\varphi).$
We thus have that ${{\boldsymbol{\mu}}}^{-1}({0})/{\curly K}$ is the moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$ of solutions to Hitchin equations (2.15). In particular, if we consider the set ${{\boldsymbol{\mu}}}_{*}^{-1}({0})$ of irreducible solutions (equivalently, smooth) one has that
${{\boldsymbol{\mu}}}_{*}^{-1}({0})/{\curly K}$
is a hyperkähler manifold which, by Theorem 2.7, is homeomorphic to the subvariety of smooth points of the moduli space $\mathcal{M}(G)$, consisting of stable and simple $G$-Higgs bundles on $E_{G}$.
2.6. Non-abelian Hodge correspondence
Let us now denote by $\mathcal{M}_{c}(G)$ the moduli space of isomorphism classes of polystable $G$-Higgs bundles where the $G$-bundle has tological class $c\in\pi_{1}(G)$. We have the following.
Theorem 2.8.
There is a homeomorphism $\mathcal{R}_{c}(G)\cong\mathcal{M}_{c}(G)$.
Remark 2.9.
On the open subvarieties defined by the smooth points of $\mathcal{R}_{c}(G)$ and $\mathcal{M}_{c}(G)$, this correspondence is in fact an isomorphism of real analytic varieties.
Theorem 2.8 is proved by combining Proposition 2.2 with Theorems 2.5 and 2.7, together with the following proposition.
Proposition 2.10.
The correspondence $(A,\varphi)\mapsto(A,\psi:=-i(\varphi-\tau(\varphi)))$ defines a homeomorphism
$\mathcal{M}^{\operatorname{Hit}}(G)\cong\mathcal{M}^{\operatorname{Harm}}(G).$
One can easily see that under the affine map
$\begin{array}[]{ccc}{\curly A}\times\Omega&\longrightarrow&{\curly D}\\ (A,\varphi)&\longmapsto&A-i\varphi+i\tau(\varphi)\end{array}$
${\curly A}\times\Omega$ with complex structure $J_{2}$ corresponds to $\curly D$} with complex structure $\mathbf{J}$ (see [Hi1]).
Now, Theorems 2.7 and 2.5 can be regarded as existence theorems, establishing the non-emptiness of the hyperkähler quotient, obtained by focusing on different complex structures. For Theorem 2.7 one gives a special status to the complex structure $J_{1}$. Combining the symplectic forms determined by $J_{2}$ and $J_{3}$ one has the $J_{1}$-holomorphic symplectic form $\omega_{c}=\omega_{2}+i\omega_{3}$ on ${\curly A}\times\Omega$. The gauge group $\curly{G}={\curly K}^{\mathbb{C}}$ acts on ${\curly A}\times\Omega$ preserving $\omega_{c}$. The symplectic quotient construction can also be extended to the holomorphic situation (see e.g. [K]) to obtain the holomorphic symplectic quotient $\{(J_{E},\varphi)\;:\;\bar{\partial}_{E}\varphi=0\}/\curly{G}$. What Theorem 2.7 says is that for a class $(J_{E},\varphi)]$ in this quotient to have a representative (unique up to $K$-gauge) satisfying $\mu_{1}=0$ it is necessary and sufficient that the pair $(J_{E},\varphi)$ be polystable. This identifies the hyperkähler quotient to the set of equivalence classes of polystable $G$-Higgs bundles on $E_{G}$. If one now takes $J_{2}$ on ${\curly A}\times\Omega$ or equivalently $\curly D$} with $\mathbf{J}$ and argues in a similar way, one gets Theorem 2.5 identifying the hyperkähler quotient to the set of equivalence classes of reductive flat connections on $E_{G}$.
3. Kähler fibrations and coupled harmonic equations
3.1. Kähler fibrations
In this section we recall some basic aspects of the theory of Kähler fibrations, which we will use later. We follow closely [GLS, M].
A Kähler fibration is a holomorphic fibre bundle $\curly{X}\to\curly{J}$, with typical fiber a Kähler manifold $({\curly D},\omega_{\curly D})$, and such that transition functions between local holomorphic trivializations are contained in the group of Kähler isometries of $\curly D$}. Equivalently, $\curly{X}$ admits a smoothly varying Kähler structure on the fibres, that we shall denote $\widehat{\boldsymbol{\omega}}$. An Ehresmann connection $\Gamma$ on $\curly{X}$, given by a distribution of horizontal subspaces
$H^{\Gamma}\subset T\curly{X},$
is said to be Kähler if the associated parallel transport is by Kähler isometries of the fibres. By the fibrewise non-degeneracy of $\widehat{\boldsymbol{\omega}}$, Ehresmann connections on $\curly{X}$ correspond to real $2$-forms $\boldsymbol{\sigma}\in\Omega^{2}(\curly{X},\mathbb{R})$, which restrict to $\widehat{\boldsymbol{\omega}}$ on the fibres. Given such a $\boldsymbol{\sigma}$, the horizontal subspace of the associated connection $\Gamma^{\boldsymbol{\sigma}}$ is
(3.1) $H^{\boldsymbol{\sigma}}=\{v\in T\curly{X}\;|\;i_{v}\boldsymbol{\sigma}_{|V\curly{X}}=0\},$
where $V\curly{X}\subset T\curly{X}$ is the vertical bundle of the fibration $\curly{X}\to\curly{J}$. Conversely, given a connection $\Gamma$, we can define
$\boldsymbol{\sigma}_{\Gamma}=\widehat{\boldsymbol{\omega}}(\Gamma,\Gamma),$
where $\Gamma\colon T\curly{X}\to V\curly{X}$ is the projection induced by $\Gamma$.
Definition 3.1.
[…]
The following basic result shows that the existence of closed $(1,1)$-form restricting on the fibres to $\widehat{\boldsymbol{\omega}}$ is indeed a sufficient condition for the existence of a Kähler connection. The proof follows similarly as in [GLS, Theorem 1.2.4] and is ommited.
Theorem 3.2.
[…]
$d\boldsymbol{\sigma}=0.$
If this is the case, the horizontal subspace of the induced connection is given by $H^{\boldsymbol{\sigma}}$ in (3.1), and $\boldsymbol{\sigma}$ is a coupling form for $\Gamma^{\boldsymbol{\sigma}}$.
The existence a coupling form is a non-trivial question, related to the topology on the fibration. In the next result we recall a remarkable identity which relates the curvature of $\Gamma$ with the horizontal part of any coupling $\boldsymbol{\sigma}$ (see [GLS, Eq. (1.12)]). Recall that the curvature of an Ehresmann connection $\Gamma$ on $\curly{X}\to\curly{J}$ is the basic two-form $F_{\Gamma}\in\Omega^{2}(\curly{X},V\curly{X})$ defined by
$F_{\Gamma}(v_{1},v_{2})=-\Gamma[v_{1}^{\Gamma},v_{2}^{\Gamma}],$
where $v_{j}\in T\curly{X}$, and $v_{j}^{\Gamma}\in H^{\Gamma}$ denotes the horizontal projection with respect to $\Gamma$.
Proposition 3.3.
Assume the hypothesis of Theorem 3.2. Then,
(3.2) $\widehat{\boldsymbol{\omega}}(F_{\Gamma}(v_{1},v_{2}),\Gamma\cdot)=-d(\boldsymbol{\sigma}(v_{1}^{\Gamma},v_{2}^{\Gamma}))(\Gamma\cdot).$
In particular, if the horizontal part of the coupling form $\boldsymbol{\sigma}$ is non-constant along the fibres, then the connection $\Gamma$ cannot be flat.
When the fibre $\curly D$} is compact and simply connected, the previous result implies the existence of a natural choice of coupling form $\boldsymbol{\sigma}$, given by imposing a ‘Gysin type condition’ (see [GLS, Th. 1.4.1]). The idea is that, under this hypothesis, the vertical vector field $F_{\Gamma}(v_{1},v_{2})$ is Hamiltonian, and hence it determines up to a constant the vertical variation of the function $\boldsymbol{\sigma}(v_{1}^{\Gamma},v_{2}^{\Gamma})$. More explicitly, provided that there exists a coherent choice of fibrewise Hamiltonian function $\mu(v_{1},v_{2})$ for $F_{\Gamma}(v_{1},v_{2})$, for any pair of horizontal vector fields $v_{1},v_{2}$, that is,
$d\mu(v_{1},v_{2})=\widehat{\boldsymbol{\omega}}(F_{\Gamma}(v_{1},v_{2}),)$
the natural choice of coupling form is
(3.3) $\boldsymbol{\sigma}=\widehat{\boldsymbol{\omega}}(\Gamma,\Gamma)-\mu(v_{1},v_{2}).$
This is the case, for instance, if the connection $\Gamma$ has holonomy contained in a subgroup $K$ of Hamiltonian isommetries of the fibre $\curly D$}, and there is a $K$-equivariant moment map $\mu\colon{\curly D}\to\mathfrak{k}^{*}$.
A fundamental question in the theory of Kähler fibrations is whether there exists a Kähler metric $\omega{X}$ on $\curly{X}$ which restricts to $\widehat{\boldsymbol{\omega}}$ on the fibres. This is the case, for instance, if $\curly{J}$ is Kähler, with Kähler form $\boldsymbol{\omega}{J}$, and both $\curly{X}$ and $\curly{J}$ are compact. In this situation, and under the hypothesis of Theorem 3.2, a natural choice of Kähler form on $\curly{X}$ is the minimal coupling
$\boldsymbol{\omega}_{\alpha}=\boldsymbol{\omega}{J}+\alpha\boldsymbol{\sigma}$
for a small coupling constant $0<\alpha\ll 1$.
In the present paper, we are interested in the study of Kähler fibrations satisfying the hypothesis of Theorem 3.2. We will find situations in which the pre-symplectic manifold $(\curly{X},\boldsymbol{\sigma})$ admits a Hamiltonian action of a real Lie group $\widetilde{\curly{G}}$, preserving the holomorphic fibration structure. This will have an impact in the structure of the group of symmetries, which will typically appear as a non-trivial extension
(3.4) $1\to\curly{G}\longrightarrow{}\widetilde{\curly{G}}\longrightarrow\curly{H}\to 1,$
of a (real) subgroup $\curly{H}$ of the holomorphic automorphisms of the base $\curly{J}$ by a subgroup of Kähler isometries $\curly{G}$ of the typical fibre $\curly D$}. Previous work on this type of reductions, in the context of gauge theory, can be found in [AGG, AGG3, AGGP, AGGPY].
3.2. Kähler connection and norm squared of the Higgs field
We fix a smooth oriented compact surface $\Sigma$. We also consider a connected semisimple complex Lie group $G$, with Lie algebra $\mathfrak{g}$, and fix an antiholomorphic involution $\tau$ of $G$ defining a maximal compact subgroup $K:=G^{\tau}\subset G$, with Lie algebra $\mathfrak{k}$.
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$. Let $\curly D$} be the space of connections on $E_{G}$, equipped with the constant complex structure (2.1). Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation. Consider the space
(3.5) $\curly{X}=\curly{J}\times{\curly D}$
endowed with the product complex structure
$\mathbb{J}(\dot{J},\dot{D})=(J\dot{J},\mathbf{J}\dot{D}).$
Here we identify $T_{J}\curly{J}$ with the space of endomorphisms $\dot{J}\colon T\Sigma\to T\Sigma$ such that $\dot{J}J=-J\dot{J}$. In particular, the map $\pi_{1}\colon\curly{X}\to\curly{J}$ is holomorphic. By construction, a natural structure of Kähler fibration $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$ on $\curly{X}$ over $\curly{J}$ is given by the $\curly{J}$-dependent symplectic structure $\omega_{\mathbf{J}}(J)$ on $\curly D$}, defined in (2.13). We should emphasize that, even though the holomorphic fibration structure on $\curly{X}$ is trivial, $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$ has a non-trivial dependence on the base $\curly{J}$.
In order to give a more explicit description of $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$, we consider $h\in\Omega^{0}(E_{G}(G/K))$ a reduction of structure group of $E_{G}$ to $K$, and let $E_{K}$ be the corresponding principal $K$-bundle. From the decomposition $\mathfrak{g}=\mathfrak{k}\oplus i\mathfrak{k}$ one has the identification ${\curly D}\cong{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ in (2.7), which induces a biholomorphism
(3.6) $\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k})),$
where the right hand side is endowed with the product complex structure
(3.7) $\mathbb{J}(\dot{J},a,\psi)=(J\dot{J},-\dot{\psi},a).$
Consider the structure of Kähler fibration $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$ on $\curly{X}\to\curly{J}$ defined above, which in the present setup can be described explicitly by
(3.8) $\widehat{\boldsymbol{\omega}}_{\mathbf{J}|(J,A,\psi)}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2}))=\int_{\Sigma}B(a_{1}\wedge J\dot{\psi}_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge Ja_{2}).$
One can clearly see that $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$ has a non-trivial dependence on $J\in\curly{J}$, and hence it defines a non-trivial structure of Kähler fibration on $\curly{X}\to\curly{J}$.
The aim of this section is to prove that $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}})_{\mathbf{J}}$ admits a Kähler Ehresmann connection, following Theorem 3.2. For this, we follow a suggestion by N. Hitchin (cf. [Hi1, Section 9]) and define a real $(1,1)$-form $\boldsymbol{\sigma}^{\mathbb{J}}\in\Omega^{1,1}(\curly{X},\mathbb{R})$ by
(3.9) $\boldsymbol{\sigma}^{\mathbb{J}}=i\partial_{\mathbb{J}}\bar{\partial}_{\mathbb{J}}\|\psi\|^{2}_{L^{2}}$
where $\|\psi\|^{2}_{L^{2}}$ denotes the $J$-dependent $L^{2}$-norm of the unitary Higgs field $\psi$, namely
$\|\psi\|^{2}_{L^{2}}=\int_{\Sigma}B(\psi\wedge J\psi).$
In the next result we calculate an explicit formula for $\boldsymbol{\sigma}^{\mathbb{J}}$.
Lemma 3.4.
(3.10) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}_{|(J,A,\psi)}((\dot{J}_{1},a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge(J\dot{\psi}_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B(a_{2}\wedge(J\dot{\psi}_{1}-\psi(\dot{J}_{1})))\\ &-\int_{\Sigma}B\left(\left(J\dot{\psi}_{1}-\tfrac{1}{2}\psi(\dot{J}_{1})\right)\wedge\psi(\dot{J}_{2})\right)\\ &+\int_{\Sigma}B\left(\left(J\dot{\psi}_{2}-\tfrac{1}{2}\psi(\dot{J}_{2})\right)\wedge\psi(\dot{J}_{1})\right)\end{split}$
Proof.
We set $\nu(J,A,\psi):=\|\psi\|^{2}_{L^{2}}$ and calculate
$d\nu_{|(J,A,\psi)}(\dot{J},a,\dot{\psi})=2\int_{\Sigma}B(\dot{\psi}\wedge J\psi)-\int_{\Sigma}B(\psi\wedge\psi(\dot{J})).$
By definition of $\mathbb{J}$, we also have
$d^{c}_{\mathbb{J}}\nu_{|(J,A,\psi)}(\dot{J},a,\dot{\psi})$ $=-2\int_{\Sigma}B(a\wedge J\psi)+\int_{\Sigma}B(\psi\wedge\psi(J\dot{J}))$
$=-2\int_{\Sigma}B(a\wedge J\psi)-\int_{\Sigma}B(J\psi\wedge\psi(\dot{J})).$
The statement follows now from $dd^{c}_{\mathbb{J}}\nu=2i\partial_{\mathbb{J}}\bar{\partial}_{\mathbb{J}}\nu$. ∎
We will need the following technical lemma.
Lemma 3.5.
Given $C\in\operatorname{End}T\Sigma$ and $\psi_{1},\psi_{2}\in\Omega^{1}(X,E_{K}(\mathfrak{k}))$, the following pointwise identity holds:
$B(\psi_{1}(C)\wedge\psi_{2})+B(\psi_{1}\wedge\psi_{2}(C))=0.$
Proof.
By direct calculation at a point $x\in\Sigma$, if $C=v\otimes\gamma\in T_{x}\Sigma\otimes T^{*}_{x}\Sigma$, by dimensional reasons we have
$0$ $=\gamma\wedge i_{v}B(\psi_{1}\wedge\psi_{2})$
$=\gamma\wedge(B(\psi_{1}(v),\psi_{2})-B(\psi_{1},\psi_{2}(v)))=B(\psi_{1}(v)\gamma\wedge\psi_{2})+B(\psi_{1}\wedge\psi_{2}(v)\gamma).$
Our next result proves the existence of a natural Kähler Ehresmann connection in our Kähler fibration $\curly{X}\to\curly{J}$, with coupling form $\boldsymbol{\sigma}^{\mathbb{J}}$ and curvature given (essentially) by the unitary Higgs field $\psi$.
Proposition 3.6.
The Kähler fibration $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ admits a Kähler Ehresmann connection $\Gamma^{\mathbb{J}}\colon T\curly{X}\to V\curly{X}$, with horizontal bundle given by
$H^{\mathbb{J}}=\{v\in T\curly{X}\;|\;i_{v}\boldsymbol{\sigma}^{\mathbb{J}}_{|V\curly{X}}=0\},$
where $\sigma$ is defined by (3.9). More explicitly,
(3.11) $H^{\mathbb{J}}_{|(J,A,\psi)}=\{(\dot{J},\psi(\dot{J}),(J\psi)(\dot{J}))\;|\;\dot{J}\in T_{J}\curly{J}\}.$
Furthermore, $H^{\mathbb{J}}$ is preserved by $\mathbb{J}$ and the curvature $F_{\mathbb{J}}:=F_{\Gamma^{\mathbb{J}}}\in\Omega^{2}(\curly{X},V\curly{X})$ of $\Gamma$ is of type $(1,1)$ and given explicitly by
(3.12) $\begin{split}(F_{\mathbb{J}})_{|(J,A,\psi)}(v_{1},v_{2})&=(0,\psi([\dot{J}_{2},\dot{J}_{1}]),0).\end{split}$
for any pair of horizontal vector fields $v_{1},v_{2}\in H^{\mathbb{J}}$ covering $\dot{J}_{1},\dot{J}_{2}$, respectively.
Proof.
The existence of $\Gamma$ follows from Theorem 3.2 combined with Lemma 3.4, which implies that
(3.13) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge J\dot{\psi}_{2})+\int_{\Sigma}B(J^{2}\dot{\psi}_{1}\wedge Ja_{2})\\ &=\widehat{\boldsymbol{\omega}}_{\mathbf{J}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2})).\end{split}$
where we ommit the evaluation at the point $(J,A,\psi)$ for simplicity in the notation. Contracting now $\boldsymbol{\sigma}^{\mathbb{J}}$ with a vertical vector field
$\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}((0,a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge(J\dot{\psi}_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B\left(J\dot{\psi}_{1}\wedge(\psi(\dot{J}_{2})-a_{2})\right)\end{split}$
and therefore $(\dot{J}_{2},a_{2},\dot{\psi}_{2}))\in H^{\mathbb{J}}$ if and only if
$a_{2}=\psi(\dot{J}_{2}),\qquad\dot{\psi}_{2}=(J\psi)(\dot{J}_{2}).$
Since $\boldsymbol{\sigma}^{\mathbb{J}}$ is of type $(1,1)$ and $\mathbb{J}$ preserves $V\curly{X}$, it follows that $\mathbb{J}$ preserves $H^{\mathbb{J}}$. Finally, evaluating $\boldsymbol{\sigma}^{\mathbb{J}}$ in a pair of horizontal vector fields $v_{j}=(\dot{J}_{j},\psi(\dot{J}_{j}),(J\psi)(\dot{J}_{j}))$, we have
$\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2})&=-\int_{\Sigma}B\left(\left(\psi(\dot{J}_{1})-\tfrac{1}{2}\psi(\dot{J}_{1})\right)\wedge\psi(\dot{J}_{2})\right)+\int_{\Sigma}B\left(\left(\psi(\dot{J}_{2})-\tfrac{1}{2}\psi(\dot{J}_{2})\right)\wedge\psi(\dot{J}_{1})\right)\\ &=-\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2}))\end{split}$
where we have used that
$J((J\psi)(\dot{J}))=-J(\psi(J\dot{J}))=\psi(J\dot{J}J)=\psi(\dot{J}).$
Consequently,
$d(\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2}))((0,a,\dot{\psi}))=-\int_{\Sigma}B(\dot{\psi}(\dot{J}_{1}))\wedge\psi(\dot{J}_{2}))-\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\dot{\psi}(\dot{J}_{2})).$
On the other hand, applying Lemma 3.5,
$\widehat{\boldsymbol{\omega}}_{\mathbf{J}}((0,(J\psi)([\dot{J}_{1},\dot{J}_{2}],0),(0,a,\dot{\psi}))$ $=\int_{\Sigma}B((J\psi)([\dot{J}_{1},\dot{J}_{2}])\wedge J\dot{\psi})$
$=\int_{\Sigma}B(J(\psi([\dot{J}_{1},\dot{J}_{2}]))\wedge J\dot{\psi})$
$=\int_{\Sigma}B(\psi([\dot{J}_{1},\dot{J}_{2}])\wedge\dot{\psi})$
$=-\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\dot{\psi}(\dot{J}_{2}))+\int_{\Sigma}B(\psi(\dot{J}_{2})\wedge\dot{\psi}(\dot{J}_{1}))$
$=d(\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2}))((0,a,\dot{\psi})),$
which proves the last part of the statement, by application of Proposition 3.3. ∎
To finish this section, we provide a formula for the coupling form $\boldsymbol{\sigma}^{\mathbb{J}}$ adapted to its associated connection $\Gamma^{\mathbb{J}}$. Note that the vertical projection with respect to $\Gamma^{\mathbb{J}}$ is explicitly given by
$\Gamma^{\mathbb{J}}(\dot{J},a,\dot{\psi})=(0,a-\psi(\dot{J}),\dot{\psi}-(J\psi)(\dot{J})).$
We will also consider the symmetric tensor on $\curly{X}$ defined by
$\mathbf{g}_{\mathbb{J}}:=\boldsymbol{\sigma}^{\mathbb{J}}(,\mathbb{J}).$
By construction, this coincides with the flat hyperkähler metric (2.11) along the fibres of $\curly{X}\to\curly{J}$. As we will see shortly, $\mathbf{g}_{\mathbb{J}}$ is negative semi-definite along the horizontal directions of the connection $\Gamma^{\mathbb{J}}$. This reveals difficulties in the fundamental question of constructing a positive-definite Kähler metric on the Kähler fibration $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$.
Corollary 3.7.
For any tangent vectors $v_{j}=(\dot{J}_{j},a_{j},\dot{\psi}_{j}),v=(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$ one has
(3.14) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2})&=\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}_{\mathbb{J}}(v,v)&=\int_{\Sigma}B((a-\psi(\dot{J}))\wedge J(a-\psi(\dot{J})))\\ &+\int_{\Sigma}B((\dot{\psi}-(J\psi)(\dot{J}))\wedge J(\dot{\psi}-(J\psi)(\dot{J}))\\ &-\int_{\Sigma}B(\psi(\dot{J}))\wedge J(\psi(\dot{J}))).\end{split}$
In particular, given a horizontal vector field $v\in H^{\mathbb{J}}$ at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$, one has
(3.15) $\begin{split}\mathbf{g}_{\mathbb{J}}(v,v)=-\int_{\Sigma}B\left(\psi(\dot{J})\wedge J(\psi(\dot{J}))\right).\end{split}$
Consequently, $\mathbf{g}_{\mathbb{J}}$ is negative semi-definite along the horizontal directions of $\Gamma^{\mathbb{J}}$.
3.3. Minimal coupling and Hamiltonian action
Let $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ be the Kähler fibration defined in Section 3.2. In this section we undertake minimal couplings for the connection $\Gamma^{\mathbb{J}}$ constructed in Proposition 3.6. We will see that the corresponding (pre)symplectic structures on $\curly{X}$ admit a Hamiltonian action by a suitable extended gauge group $\widetilde{{\curly K}}$, naturally associated to a choice of symplectic structure $\omega$ on $\Sigma$. This will allow us, in Section 4, to construct pseudo-Kähler structures on the universal moduli space of solutions of the harmonicity equations (2.14) over Teichmuller space.
We start by introducing the Kähler structure we shall consider, up to sign, in the base of of the fibration $\curly{X}\to\curly{J}$. Following Donaldson [Do2] and Fujiki [F], we fix a symplectic (volume) form $\omega$ on $\Sigma$ compatible with the given orientation. The space $\curly{J}$ of complex structures $J$ on $\Sigma$ compatible with the orientation (and hence with $\omega$) is an infinite dimensional Kähler manifold, with complex structure $\mathbb{J}{J}\colon T_{J}\curly{J}\to T_{J}\curly{J}$ and Kähler form $\boldsymbol{\omega}{J}$ given, respectively, by
(3.16) $\mathbb{J}{J}\dot{J}:=J\dot{J}\;\;\text{ and }\;\;\boldsymbol{\omega}{J}(\dot{J}_{1},\dot{J}_{2}):=\frac{1}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega,$
for $\dot{J}_{1},\dot{J}_{2}\in T_{J}\curly{J}$. Note that, by dimensional reasons, any $\dot{J}\in T_{J}\curly{J}$, regarded as an endomorphism $\dot{J}\colon T\Sigma\to T\Sigma$, is symmetric with respect to the induced metric $g=\omega(\cdot,J\cdot)$
Let $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ be the Kähler fibration defined in Section 3.2. Given $\alpha>0$ a real coupling constant and $\varepsilon\in\{-1,1\}$, the family of minimal coupling symplectic structures of our interest is defined by
(3.17) $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}=\varepsilon\boldsymbol{\omega}{J}+\alpha\boldsymbol{\sigma}^{\mathbb{J}},$
where $\boldsymbol{\sigma}^{\mathbb{J}}$ is the exact $(1,1)$-form in Lemma 3.4. By construction, $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ is closed and of type $(1,1)$ with respect to the complex structure $\mathbb{J}$. Furthermore, along the fibres of $\curly{X}\to\curly{J}$ the $2$-form $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ restricts to the Kähler structure $\alpha\widehat{\boldsymbol{\omega}}$. Consider the associated symmetric tensor
(3.18) $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}:=\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(,\mathbb{J})=\varepsilon\mathbf{g}_{\curly{J}}+\alpha\mathbf{g}_{\mathbb{J}}.$
Applying Corollary 3.7, $\boldsymbol{\omega}_{\alpha}^{-1}$ is negative-definite along the horizontal subspace $H^{\mathbb{J}}\subset T\curly{X}$. The analysis for the case $\varepsilon=1$ is much more subtle, as we can see from the next result.
Lemma 3.8.
Given a horizontal vector field $v\in H^{\mathbb{J}}$ at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$, one has
(3.19) $\begin{split}\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,v)=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}\dot{J})\omega-\alpha\int_{\Sigma}B\left(\psi(\dot{J})\wedge J(\psi(\dot{J}))\right).\end{split}$
Consequently, for any $\alpha>0$ one has
  1. (1)
    For $\varepsilon=-1$, $\mathbf{g}_{\alpha}^{\varepsilon}$ is negative definite along $H^{\mathbb{J}}$. In particular, $\boldsymbol{\omega}_{\alpha}^{-1}$ is a non-degenerate symplectic structure.
  2. (2)
    For $\varepsilon=1$ and $\psi\neq 0$, $\mathbf{g}_{\alpha}^{\varepsilon}$ changes signature along the line $(J,A,\lambda\psi)\in\curly{X}$, for $\lambda\in\mathbb{R}$.
Proof.
The case $\varepsilon=-1$ is a direct consequence of Corollary 3.7. For the case $\varepsilon=1$, let $(J,A,\psi)\in\curly{X}$. Then, for any $\dot{J}\in T_{J}\curly{J}$, we have a pointwise equalities
$\operatorname{tr}(\dot{J}\dot{J})=C|\dot{J}|^{2},\qquad\Lambda_{\omega}B\left(\psi(\dot{J})\wedge J(\psi(\dot{J}))\right)=C^{\prime}|\psi|^{2}|\dot{J}|^{2},$
for suitable positive constants $C,C^{\prime}\in\mathbb{R}$, where the $|\cdot|$ denotes tensorial norms with respect to $B$ and $g=\omega(,J)$. Hence, for $\psi\neq 0$ and
$\lambda^{2}_{0}=\frac{C}{\alpha C^{\prime}|\psi|^{2}},$
the tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is:
  • positive definite at points $(J,A,\lambda\psi)\in\curly{X}$, for $|\lambda|<\lambda_{0}$,
  • positive definite along vertical directions and negative definite along the horizontal bundle $H^{\mathbb{J}}$, at points $(J,A,\lambda\psi)\in\curly{X}$, for $|\lambda|>\lambda_{0}$,
  • null at the horizontal bundle $H^{\mathbb{J}}$ at the points $(J,A,\lambda_{0}\psi)\in\curly{X}$.
The result follows now from the previous cases.
Our next goal is to prove that $(\curly{X},\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon})$ admits a Hamiltonian action by a suitable extended gauge group $\widetilde{{\curly K}}$, determined by $\omega$ and the reduction $E_{K}\subset E_{G}$. Consider the group $\curly{H}$ of Hamiltonian symplectomorphisms of $(\Sigma,\omega)$. The natural group of symmetries of our theory is the (Hamiltonian) extended gauge group $\widetilde{{\curly K}}$. By definition, $\widetilde{{\curly K}}$ is the group of automorphisms of $E_{K}$ which cover elements of the group $\curly{H}$. There is a canonical short exact sequence of Lie groups
(3.20) $1\to{\curly K}\longrightarrow{}\widetilde{{\curly K}}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\curly{H}\to 1,$
where $p$ maps each $g\in\widetilde{{\curly K}}$ into the Hamiltonian symplectomorphism $p(g)\in\curly{H}$ that it covers, and so its kernel $\curly K$} is the gauge group of $E_{K}$, that is, the normal subgroup of $\widetilde{{\curly K}}$ consisting of automorphisms of $E_{K}$ covering the identity map on $\Sigma$.
Similarly as in [AGG, Section 2.2], the group $\widetilde{{\curly K}}$ acts on $(\curly{X},\mathbb{J},\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon})$ preserving $\mathbb{J}$ and $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$, and covering the $\curly{H}$-action on $\curly{J}$ by push-forward. This last action preserves the Kähler structure on $\curly{J}$, and furthermore is Hamiltonian, with equivariant moment map $\mu\colon\curly{J}\to(\operatorname{Lie}\curly{H})^{*}$ given by
(3.21) $\langle\mu(J),\eta_{f}\rangle=-\int_{\Sigma}fS_{J}\omega,$
for $f\in C^{\infty}_{0}(\Sigma)$, identified with an element $\eta_{f}$ in $\operatorname{Lie}\curly{H}$, and $S_{J}$ is the scalar curvature of the metric $g=\omega(,J)$ defined by $\omega$ and $J$. Here, $C^{\infty}_{0}(\Sigma)$ denotes the space of smooth functions on $\Sigma$ such that $\int_{\Sigma}f\omega=0$. This fact was shown by Donaldson [Do2] and Fujiki [F] independently for symplectic manifolds of arbitrary dimensions, and it seems that was known to Quillen in the case of surfaces.
Proposition 3.9.
The action of $\widetilde{{\curly K}}$ on $(\curly{X},\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon})$ is Hamiltonian, with equivariant moment map given by
(3.22) $\begin{split}\langle\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}(J,A,\psi),\zeta\rangle&=\alpha\int_{\Sigma}B(A\zeta+2\psi(Jp\zeta),d_{A}(J\psi))\\ &-\int_{\Sigma}f\left(\varepsilon S_{J}-\alpha\Lambda_{\omega}d\left(B\left(\Lambda_{\omega}F_{A},J\psi\right)\right)\right)\omega,\end{split}$
where $A\zeta\in\Omega^{0}(X,E_{K}(\mathfrak{k}))$ denotes the vertical part of the $K$-equivariant vector field $\zeta$ on the total space of $E_{K}$, with respect to the connection $A$.
Proof.
Since $\sigma=\tfrac{1}{2}dd^{c}_{\mathbb{J}}\nu$, with $\nu(J,A,\psi):=\|\psi\|^{2}_{L^{2}}$ (see Lemma 3.4), and $d^{c}_{\mathbb{J}}\nu$ is preserved by the $\widetilde{{\curly K}}-$action, there exists an equivariant moment map given by
$\langle\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}(J,A,\psi),\zeta\rangle=-\varepsilon\int_{\Sigma}fS_{J}\omega-\frac{\alpha}{2}d^{c}_{\mathbb{J}}\nu(\zeta\cdot(J,A,\psi)),$
where $\zeta\cdot(J,A,\psi)$ is the the infinitesimal action of $\zeta$ on $(J,A,\psi)$. Using the complex connection $D=A+i\psi$ it follows that
$\zeta\cdot D$ $=-d_{D}(D\zeta)-i_{y}F_{D}$
$=-\left(d_{A}(A\zeta)+i[\psi,A\zeta]+id_{A}(\psi(y))-[\psi,\psi(y)]\right)-i_{y}\left(F_{A}-\tfrac{1}{2}[\psi\wedge\psi]+id_{A}\psi\right)$
$=-d_{A}(A\zeta)-i_{y}F_{A}+[\psi,\psi(y)]+\tfrac{1}{2}i_{y}[\psi\wedge\psi]-i\left(d_{A}(\psi(y))+i_{y}(d_{A}\psi)+[\psi,A\zeta]\right)$
$=-d_{A}(A\zeta)-i_{y}F_{A}-i\left(d_{A}(\psi(y))+i_{y}(d_{A}\psi)+[\psi,A\zeta]\right),$
where $y:=p\zeta$, and therefore, by the proof of Lemma 3.4 and integration by parts,
$d^{c}_{\mathbb{J}}\nu(\zeta\cdot(J,A,\psi))$ $=2\int_{\Sigma}B\left(\left(d_{A}(A\zeta)+i_{y}F_{A}\right)\wedge J\psi\right)-\int_{\Sigma}B(J\psi\wedge\psi(L_{y}J))$
$=-2\int_{\Sigma}\left(B(A\zeta,d_{A}(J\psi))+B\left(F_{A},J\psi(y)\right)+2B(\psi(Jy),d_{A}(J\psi))\right)$
$=-2\int_{\Sigma}B(A\zeta+2\psi(Jy),d_{A}(J\psi))-2\int_{\Sigma}fd\left(B\left(\Lambda_{\omega}F_{A},J\psi\right)\right).$
For the second equality, we have used that, for any vector field $y$ on $\sigma$, one has $L_{y}J=2i\bar{\partial}y^{1,0}-2i\partial y^{0,1}$ and therefore
$B(J\psi\wedge\psi(L_{y}J))$ $=2B(\psi^{1,0}\wedge\psi(\bar{\partial}y^{1,0}))+cc.$
$=2B(\psi^{1,0}\wedge(i_{y^{1,0}}(\bar{\partial}_{A}\psi^{1,0})+\bar{\partial}_{A}(i_{y^{1,0}}\psi)))+cc.$
$=-2\bar{\partial}(B(\psi^{1,0},\psi(y^{1,0})))+4B(\bar{\partial}_{A}\psi^{1,0},\psi(y^{1,0}))+cc.$
The $\widetilde{{\curly K}}-$action produces a ‘coupling term’ in the base $\Sigma$ (cf. Remark 4.2), which combines with the scalar curvature of the metric $g=\omega(,J)$. In particular, similarly as in [AGG], zeros of the moment map $\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}$ are given by solutions of the coupled system of equations
(3.23) $\begin{split}d_{A}^{*}\psi&=0\\ \varepsilon S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
where $\chi(\Sigma)$ is the Euler characteristic of $\Sigma$ and $V=\int_{\Sigma}\omega$ is the total volume. In the next section we combine these equations with the flatness condition $F_{D}=0$ for the complex connection, in order to introduce the universal Hitchin moduli space.
4. Universal moduli spaces
4.1. Universal moduli space of flat $\boldsymbol{G}$-connections
We fix a smooth oriented compact surface $\Sigma$. We also consider a connected semisimple complex Lie group $G$, with Lie algebra $\mathfrak{g}$, and fix an antiholomorphic involution $\tau$ of $G$ defining a maximal compact subgroup $K:=G^{\tau}\subset G$, with Lie algebra $\mathfrak{k}$.
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$. Let $\curly D$} be the space of connection on $E_{G}$, equipped with the constant complex structure (2.1). Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation. Consider the Kähler fibration $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ defined in Section 3.2. Our aim in this section is to investigate the existence of a Kähler structure on $\curly{X}$ and quotients thereof, combining the minimal coupling for the connection $\Gamma^{\mathbb{J}}$ constructed in Proposition 3.6, as considered in Section 3.3, with symplectic reduction. Difficulties will arise from Corollary 3.7, which shows that the associated symmetric tensor $\mathbf{g}_{\mathbb{J}}:=\sigma(,\mathbb{J})$ is negative semi-definite along horizontal directions.
In order to have some control in our construction, we need to impose the integrability condition $F_{D}=0$. For this, we first construct a universal moduli space of flat $G$-connections, with a holomorphic fibration structure over Teichmüller space. The first step is to show that the flatness condition for $D\in{\curly D}$ is compatible with the natural group of symmetries in the present context: there is an extended complex gauge group of symmetries, acting on $\curly{X}$ and preserving the complex structure, defined by an extension
(4.1) $1\to\curly{G}\longrightarrow{}\widetilde{\curly{G}}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\operatorname{Diff}_{0}(\Sigma)\to 1.$
Here, $\widetilde{\curly{G}}$ is the subgroup of $G$-equivariant diffeomorphisms of $E_{G}$ which project to $\operatorname{Diff}_{0}(\Sigma)$, the component of the identity in the group of diffeomorphisms of $\Sigma$, and $p$ maps each $g\in\widetilde{\curly{G}}$ into the diffeomorphism $p(g)\in\operatorname{Diff}_{0}(\Sigma)$ that it covers. Note that the kernel of $p$ is the gauge group of $E_{G}$, that is, the normal subgroup of $\widetilde{\curly{G}}$ consisting of automorphisms of $E_{G}$ covering the identity map on $\Sigma$. Note also that, for any choice of symplectic structure $\omega$ on $\Sigma$ and reduction $h\in\Omega^{0}(E_{G}(G/K))$ of structure group of $E_{G}$ to $K$, there is a natural group homomorphism $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$ induced by the inclusion map.
More explicitly, the $\widetilde{\curly{G}}-$action on $\curly{X}$ is given by
$g(J,D)=(p(g)_{*}J,gD),$
which is compatible with the holomorphic fibration structure, in the sense that
$\pi_{1}(g(J,D))=p(g)_{*}J.$
In different words, the extension (4.1) is given by restriction of the natural map from holomorphic automorphisms of $\curly{X}$ which preserve the fibration structure, to holomorphic automorphisms of the base $\curly{J}$. It is important to observe that even though $\operatorname{Diff}_{0}(\Sigma)$ acts by holomorphic automorphisms of the base $\curly{J}$, this is a real infinite-dimensional Lie group, which does not admit a complexification (see e.g. [Do3]).
The holomorphic symplectic structure $\Omega_{\mathbf{J}}$ on $\curly D$} is constant on $\curly{J}$ (see (2.1)), and hence it makes sense to consider the holomorphic presymplectic structure on $\curly{X}$, defined by pull-back
$\Omega_{\mathbb{J}}=\pi_{2}^{*}\Omega_{\mathbf{J}}.$
Similarly as in the previous section, $\Omega_{\mathbb{J}}$ induces a structure of holomorphic symplectic fibration on $\curly{X}\to\curly{J}$ and a holomorphic Ehresmann connection, which in this case is simply the trivial connection. We will return to this structure when studying the universal Higgs bundle moduli space in Section 5. The following result is an immediate consequence of the proof of [AGG, Proposition 1.6], but we sketch the proof in the present setup for the convenience of the reader.
Proposition 4.1.
The action of $\widetilde{\curly{G}}$ on $(\curly{X},\Omega_{\mathbb{J}})$ is Hamiltonian, with equivariant moment map given by
$\langle\mu_{\widetilde{\curly{G}}}(J,D),\zeta\rangle=\int_{\Sigma}B(F_{D},D\zeta),$
where $D\zeta\in\Omega^{0}(X,E_{G}(\mathfrak{g}))$ denotes the vertical part of the $G$-equivariant vector field $\zeta$ on the total space of $E_{G}$, with respect to the connection $D$.
Proof.
Taking variations on the formula for $\mu_{\widetilde{\curly{G}}}$ and integrating by parts, we have
$\langle d\mu_{\widetilde{\curly{G}}}(\dot{J},\dot{D}),\zeta\rangle$ $=\int_{\Sigma}B(d_{D}(\dot{D}),D\zeta)+\int_{\Sigma}B(F_{D},i_{p\zeta}\dot{D})$
$=\int_{\Sigma}B(\dot{D}\wedge d_{D}(D\zeta))-\int_{\Sigma}B(i_{p\zeta}F_{D}\wedge\dot{D})$
$=-\int_{\Sigma}B((d_{D}(D\zeta)+i_{p\zeta}F_{D})\wedge\dot{D}).$
On the other hand, the infinitesimal action of $\zeta$ on $D$ is given by
$\zeta\cdot D=-d_{D}(D\zeta)-i_{p\zeta}F_{D}$
and the result follows. ∎
It is remarkable that, unlike in the moment calculation in Proposition 3.9, in the present setup the extension of the complex gauge group by diffeomorphisms does not produce any additional ‘coupling term’ in the base $\Sigma$ (cf. Remark 4.2). One can intepret this fact as a functorial property of flat connections under the action of diffeomorphisms on a surface. This allows us to define the universal moduli space of flat $G$-connections, as follows, and to impose the flatness condition for $D$ as an integrability condition in the next sections.
Let us denote ${\curly D}^{*}\subset{\curly D}$ the complex subspace of reductive connections, and set
$\curly{X}^{*}=\curly{J}\times{\curly D}^{*}.$
It is not difficult to see that $\curly D$$^{*}$, and hence $\curly{X}^{*}$, is preserved by the $\widetilde{\curly{G}}-$action. Therefore, considering the pull-back of $\Omega_{\mathbb{J}}$ to $\curly{X}^{*}$, there is an induced moment map away from the singularities of $\curly{X}^{*}$, which, by abuse of notation, we denote simply by
$\mu_{\widetilde{\curly{G}}|\curly{X}^{*}}\colon\curly{X}^{*}\to\operatorname{Lie}\widetilde{\curly{G}}^{*}$.
From the previous discussion, it is natural to define the universal moduli space of flat $G$-connections on $E_{G}$ as the complex symplectic quotient
$\mathcal{U}^{\operatorname{Flat}}(G):=\{(J,D)\;\in\;\curly{X}^{*}\;\;\mbox{with}\;\;F_{D}=0\}/\widetilde{\curly{G}}=\mu_{\widetilde{\curly{G}}|\curly{X}^{*}}^{-1}(0)/\widetilde{\curly{G}}$.
By construction, $\mathcal{U}^{\operatorname{Flat}}(G)$ fibers over the Teichmüller space $\mathcal{T}$ of complex structures modulo diffeomorphisms isotopic to the identity
(4.2) $\mathcal{U}^{\operatorname{Flat}}(G)\to\mathcal{T}:=\curly{J}/\operatorname{Diff}_{0}(\Sigma).$
This fibration is naturally holomorphic, and $\Omega_{\mathbb{J}}$ induces a structure of holomorphic symplectic fibration with flat Ehresmann connection.
Remark 4.2.
For a $G$-bundle on a higher dimensional symplectic manifold $(M^{2n},\omega)$, consider the subgroup $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$ given by the preimage by $p$ of the group of Hamiltonian symplectomorphisms on $M$. Then, the proof of [AGG, Proposition 1.6] shows that the induced $\widetilde{{\curly K}}-$action on $\curly D$} is Hamiltonian, with equivariant moment map
$\langle\boldsymbol{\mu}_{\widetilde{{\curly K}}}(J,D),\zeta\rangle=\frac{1}{(n-1)!}\int_{M}B(F_{D},D\zeta)\wedge\omega^{n-1}-\frac{1}{2(n-2)!}\int_{M}fB(F_{D}\wedge F_{D})\wedge\omega^{n-2},$
where $f\in C^{\infty}(M,\mathbb{R})$ is the unique smooth function satisfying $i_{y}\omega=df$ and $\int_{M}f\omega^{n}=0$, for $p(\zeta)=y$. In this case, there is a ‘coupling term’ in the base corresponding to the function
$\frac{(n-1)}{2}\frac{B(F_{D}\wedge F_{D})\wedge\omega^{n-2}}{\omega^{n}}.$
4.2. Universal moduli space of the coupled harmonic equations
In this section we combine the results of the previous two sections in order to define universal moduli spaces of solutions of the harmonicity equations (2.14), varying over Teichmuller space. As in the previous sections, we consider a fixed principal $K$-bundle $E_{K}$ over a smooth compact oriented surface $\Sigma$ with fixed symplectic form $\omega$.
Definition 4.3.
We say that a triple $(J,A,\psi)\in\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ is a solution of the coupled harmonic equations with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, if the following conditions are satisfied
(4.3) $\begin{split}F_{A}-\frac{1}{2}[\psi,\psi]&=0,\\ d_{A}\psi&=0,\\ d_{A}^{\ast}\psi&=0,\\ S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi\right)\right)&=\frac{2\pi\chi(\Sigma)}{V},\end{split}$
where $g=\omega(,J)$ and * is the corresponding Hodge star operator.
As in Section 4.1, consider ${\curly D}^{*}\subset{\curly D}$ the complex subspace of reductive flat connections. Then, $\curly{X}^{*}:=\curly{J}\times{\curly D}^{*}$ is formally a complex submanifold of $\curly{X}$ preserved by the $\widetilde{{\curly K}}-$action (see Section 3.3), and inherits a minimal coupling structure $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ and moment map $\mu_{\widetilde{{\curly K}}|\curly{X}^{*}}$. We define the universal Hitchin moduli space as the symplectic quotient
$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}:=\mu_{\widetilde{{\curly K}}|\curly{X}^{*}}^{-1}(0)/\widetilde{{\curly K}}$,
By construction, $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ parametrizes solutions of the coupled harmonic equations (4.3) with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, modulo the $\widetilde{{\curly K}}-$action, and inherits a natural pre-symplectic structure on its smooth locus. This structure is furthermore symplectic in the case $\varepsilon=-1$, by Lemma 3.8. There are natural maps (cf. (5.3))
$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)\to\mathcal{T}.$
Hence, in particular, $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ can be regarded as a fibration over the Teichmüller space $\mathcal{T}$. Observe that, since the symmetric $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is not positive definite (see Lemma 3.8), it is not obvious a priori that $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ inherits a complex structure compatible with the (pre)symplectic structure induced by $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$. The main goal of this section is to study sufficient conditions under which this natural condition for the moduli space holds, furthermore proving that the map $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)$ is holomorphic.
The first step is to undertake a gauge fixing for solutions of the coupled harmonic equations (4.3), whereby the complex structure (3.7) and the symmetric tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ descend to the moduli space. Difficulties will arise, due to the fact that $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is neither a definite pairing nor non-degenerate.
Consider a solution $(J,A,\psi)$ of the coupled harmonic equations (4.3) with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$. We start by characterizing the tangent space to $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ at $(J,A,\psi)]$. An infinitesimal variation of the triple $(J,A,\psi)$ is given by
(4.4) $(\dot{J},a,\dot{\psi})\in\mathcal{S}^{1}:=T_{J}\curly{J}\oplus\Omega^{1}(X,E_{K}(\mathfrak{k}))\oplus\Omega^{1}(X,E_{K}(\mathfrak{k})).$
As above, we identify $T_{J}\curly{J}$ with the space of endomorphisms $\dot{J}\colon T\Sigma\to T\Sigma$ such that $\dot{J}J=-J\dot{J}$. The proof of the following lemma follows from a straightforward calculation.
Lemma 4.4.
The linearization of the coupled harmonic equations (4.3) at $(J,A,\psi)$ is given by
(4.5) $\begin{split}d_{A}a-[\dot{\psi},\psi]&=0,\\ d_{A}\dot{\psi}+[a,\psi]&=0,\\ d_{A}(J\dot{\psi})+[a,J\psi]-d_{A}(\psi(\dot{J}))&=0,\\ \varepsilon\delta S(\dot{J})-\alpha*d(B(\Lambda_{\omega}d_{A}a,*\psi))-\alpha*d(B(\Lambda_{\omega}F_{A},*\dot{\psi}))+\alpha*d(B(\Lambda_{\omega}F_{A},\psi(\dot{J})))&=0,\end{split}$
where $\delta S\colon T_{J}\curly{J}\to C_{0}^{\infty}(\Sigma,\mathbb{R})$ is the linearization of the scalar curvature.
We denote by ${\bf L}_{\alpha}^{\varepsilon}(\dot{J},a,\dot{\psi})$ the differential operator defined by the left-hand side of equations (5.17). We turn next to the study of the infinitesimal action, in order to define a complex. From the proof of Proposition 3.9, using the connection $A$ we can identify elements $\zeta\in\operatorname{Lie}\widetilde{{\curly K}}$ with pairs
$\zeta\cong(f,u)\in\mathcal{S}^{0}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(X,E_{K}(\mathfrak{k})),$
and the infinitesimal action $\mathbf{P}(f,u):=(f,u)\cdot(J,A,\psi)$ at $(J,A,\psi)$ is
(4.6) $\mathbf{P}(f,u)=-(L_{\eta_{f}}J,d_{A}u+i_{\eta_{f}}F_{A},d_{A}(i_{\eta_{f}}\psi)+[\psi,u]).$
Define the vector space
$\mathcal{S}^{2}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus(\Omega^{2}(X,E_{K}(\mathfrak{k}))^{\oplus 3},$
so that ${\bf L}_{\alpha}^{\varepsilon}(\dot{J},a,\dot{\psi})\in\mathcal{S}^{2}$ (where we read equations (5.17) from bottom to top), and consider the complex of linear differential operators
(4.7) ${(\mathcal{S}^{*})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{S}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}}$${\mathcal{S}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L}_{\alpha}^{\varepsilon}}$${\mathcal{S}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
The cohomology $H^{1}(\mathcal{S}^{*}):=\frac{\operatorname{Ker}{\bf L}_{\alpha}^{\varepsilon}}{\operatorname{Im}\,\mathbf{P}}$ can be formally identified with the tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$. Our next result shows that the moduli space $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ is finite dimensional. Notice that both $\mathbf{P}$ and ${\bf L}_{\alpha}^{\varepsilon}$ are multi-degree differential operators, and hence we shall use the use the generalized notion of ellipticity provided by Douglis and Nirenberg [DN]. For the general theory of linear multi-degree elliptic differential operators we refer to [LM].
Lemma 4.5.
The sequence (4.7) is an elliptic complex of multi-degree linear differential operators. Consequently, the cohomology groups $H^{j}(\mathcal{S}^{*})$, with $j=0,1,2$, are finite-dimensional.
Proof.
We consider the decompositions
$\mathbf{P}=\left(\begin{array}[]{cccc}\mathbf{P}_{11}&\mathbf{P}_{12}\\ \mathbf{P}_{21}&\mathbf{P}_{22}\\ \mathbf{P}_{31}&\mathbf{P}_{32}\\ \end{array}\right)\qquad{\bf L}_{\alpha}^{\varepsilon}=\left(\begin{array}[]{cccc}\mathbf{L}_{11}&\mathbf{L}_{12}&\mathbf{L}_{13}\\ \mathbf{L}_{21}&\mathbf{L}_{22}&\mathbf{L}_{23}\\ \mathbf{L}_{31}&\mathbf{L}_{32}&\mathbf{L}_{33}\\ \mathbf{L}_{41}&\mathbf{L}_{42}&\mathbf{L}_{43}\end{array}\right),$
where to clarify the notation, we notice e.g. that
$\mathbf{P}_{11}(f)$ $=-L_{\eta_{f}}J,$ $\mathbf{P}_{12}$ $=0,$ $\mathbf{P}_{21}(f)$ $=-i_{\eta_{f}}F_{A}$
$\mathbf{P}_{22}(u)$ $=-d_{A}u,$ $\mathbf{P}_{31}(f)$ $=-d_{A}(i_{\eta_{f}}\psi),$ $\mathbf{P}_{32}(u)$ $=-[\psi,u].$
Then, the tuples $\mathbf{t}=(2,2)$ and $\mathbf{s}=(0,1,0)$ form a system of orders for $\mathbf{P}$, since
$o(\mathbf{P}_{11})$ $=2\leqslant 2-0,$ $o(\mathbf{P}_{12})$ $=0\leqslant 2-0,$ $o(\mathbf{P}_{21})$ $=1\leqslant 2-1$
$o(\mathbf{P}_{22})$ $=1\leqslant 2-1,$ $o(\mathbf{P}_{31})$ $=2\leqslant 2-0,$ $o(\mathbf{P}_{32})$ $=0\leqslant 2-0.$
and the associated leading symbol is
$\sigma_{\mathbf{P}}(v)(f,u)=(\sigma_{\mathbf{P}_{1}}(v),\sigma_{\mathbf{P}_{2}}(v),\sigma_{\mathbf{P}_{3}}(v)),$
with
$\sigma_{\mathbf{P}_{1}}(v)$ $=-fJv\otimes\omega^{-1}v,$
$\sigma_{\mathbf{P}_{2}}(v)$ $=-v\otimes u-fi_{\omega^{-1}v}F_{A},$
$\sigma_{\mathbf{P}_{3}}(v)$ $=fv\otimes\psi(Jg^{-1}v).$
Similarly, $\mathbf{t}=(2,2,2)$ and $\mathbf{s}=(0,1,1,1)$ form a system of orders for ${\bf L}_{\alpha}^{\varepsilon}$, since
$o(\mathbf{L}_{11})$ $=2\leqslant 2-0,$ $o(\mathbf{L}_{12})$ $=2\leqslant 2-0,$ $o(\mathbf{L}_{13})$ $=1\leqslant 2-0$
$o(\mathbf{L}_{21})$ $=1\leqslant 2-1,$ $o(\mathbf{L}_{22})$ $=0\leqslant 2-1,$ $o(\mathbf{L}_{23})$ $=1\leqslant 2-1,$
$o(\mathbf{L}_{31})$ $=0\leqslant 2-1,$ $o(\mathbf{L}_{32})$ $=0\leqslant 2-1,$ $o(\mathbf{L}_{33})$ $=1\leqslant 2-1,$
$o(\mathbf{L}_{41})$ $=0\leqslant 2-1,$ $o(\mathbf{L}_{42})$ $=1\leqslant 2-1,$ $o(\mathbf{L}_{43})$ $=0\leqslant 2-1,$
and the associated leading symbol is
$\sigma_{{\bf L}_{\alpha}^{\varepsilon}}(v)(\dot{J},a,\dot{\psi})=(\sigma_{\mathbf{L}_{1}}(v),\sigma_{\mathbf{L}_{2}}(v),\sigma_{\mathbf{L}_{3}}(v),\sigma_{\mathbf{L}_{4}}(v)),$
with
$\sigma_{\mathbf{L}_{1}}(v)$ $=\varepsilon*v\wedge v(\dot{J})-\alpha*v\wedge B(\Lambda_{\omega}(v\wedge a),*\psi),$
$\sigma_{\mathbf{L}_{2}}(v)$ $=-v\wedge\psi(\dot{J})+v\wedge J\dot{\psi},$
$\sigma_{\mathbf{L}_{3}}(v)$ $=v\wedge\dot{\psi},$
$\sigma_{\mathbf{L}_{4}}(v)$ $=v\wedge a.$
We prove next that the associated sequence of symbols is exact. Assuming $\sigma_{\mathbf{P}}(v)(f,u)=0$, it follows from $\sigma_{\mathbf{P}_{1}}(v)=0$ that $f=0$. Hence, $\sigma_{\mathbf{P}_{2}}(v)=0$ implies $u=0$, and $\sigma_{\mathbf{P}}(v)$ is injective.
To prove that $\sigma_{{\bf L}_{\alpha}^{\varepsilon}}(v)$ is surjective, we take $w\in\mathcal{S}^{2}$, which we can assume to be of the form
$w=(\nu,u_{2}\omega,u_{3}\omega,u_{4}\omega).$
By dimensional reasons $v\wedge\omega=0$, and hence
$u_{4}\omega=v\wedge a,\qquad\textrm{ for }a\in\Omega^{1}(X,E_{K}(\mathfrak{k}).$
We can choose $\dot{J}$ in the line spanned by $g^{-1}v\otimes Jv$ such that
$\varepsilon*v\wedge v(\dot{J})=\nu+\alpha*v\wedge B(\Lambda_{\omega}(v\wedge a),*\psi).$
Finally, with these choices, $\sigma_{{\bf L}_{\alpha}^{\varepsilon}}(v)(\dot{J},a,\dot{\psi})=w$ is equivalent to the equation
$v^{0,1}\wedge\dot{\psi}^{1,0}=(u_{2}+iu_{3})\omega+iv\wedge\psi(\dot{J}).$
Again, by dimensional reasons $v^{0,1}\wedge((u_{2}+iu_{3})\omega+iv\wedge\psi(\dot{J}))=0$, and therefore $\sigma_{{\bf L}_{\alpha}^{\varepsilon}}(v)$ is surjective.
To finish, assume that $\sigma_{{\bf L}_{\alpha}^{\varepsilon}}(v)(\dot{J},a,\dot{\psi})=0$. Then, $\sigma_{\mathbf{L}_{4}}(v)=0$ implies that there exists $x\in\Omega^{0}(X,E_{K}(\mathfrak{k})$ such that
$a=v\otimes x.$
From this, we obtain
$B(\Lambda_{\omega}(v\wedge a),*\psi)=0$
and therefore $\sigma_{\mathbf{L}_{1}}(v)=0$ implies $v\wedge v(\dot{J})=0$. By [F], there exists a smooth function $f\in C^{\infty}_{0}(\Sigma,\mathbb{R})$ such that
$\dot{J}=-fJv\otimes\omega^{-1}v$
Combined with $\sigma_{\mathbf{L}_{2}}(v)=0$, we obtain
$v\wedge J\dot{\psi}=v\wedge\psi(\dot{J})=-f\psi(\omega^{-1}v)v\wedge Jv=f\psi(Jg^{-1}v)v\wedge Jv.$
Using now that $\dot{\psi}\wedge v=0$, it follows that
$\dot{\psi}=fv\otimes\psi(Jg^{-1}v).$
Finally, writting $\omega=\lambda v\wedge Jv$ and
$F_{A}=tv\wedge Jv\textrm{ for }t\in\Omega^{0}(X,E_{K}(\mathfrak{k}),$
it follows that
$a+fi_{\omega^{-1}v}F_{A}=v\otimes(x+\lambda^{-1}ft),$
and therefore we conclude
$(\dot{J},a,\dot{\psi})=\sigma_{{\bf P}}(v)(f,-x-\lambda^{-1}ft).$
Our strategy to build a complex structure induced by (3.7) on the moduli space is to work orthogonally to the image of the infinitesimal action operator $\mathbf{P}$ in (4.7) with respect to the indefinite pairing $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ in (3.18). The existence of this complex structure will automatically yield a symmetric tensor of type $(1,1)$, since the two-form $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ in (3.17) is well defined on the cohomology $H^{1}(\mathcal{S}^{*})$ by Proposition 3.9. Consider the $L^{2}$-pairing on $\mathcal{S}^{0}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(X,E_{K}(\mathfrak{k})$, the domain of the operator $\mathbf{P}$, induced by $\omega$ and $B$:
(4.8) $\begin{split}\langle(f,u),(f,u)\rangle&{}=\int_{\Sigma}f^{2}\omega+\int_{\Sigma}B(u,u)\omega.\end{split}$
Notice that, regarded as a pairing on $\operatorname{Lie}\widetilde{{\curly K}}$ (see Section 3.3), this is $A$-dependent, and hence it can be regarded as a family of pairings varying over the configuration space $\curly{X}$ in Proposition 3.9. Consider the map $\mu\colon\curly{X}\to\mathcal{S}^{0}$ defined by (cf. (3.22))
$\mu(J,A,\psi)=(-\varepsilon S_{J}+\alpha*d(B(\Lambda_{\omega}F_{A},*\psi)),\alpha d_{A}(J\psi)).$
Consider the operator $\widetilde{\mathbf{P}}\colon\mathcal{S}^{0}\to\mathcal{S}^{1}$, defined by
(4.9) $\widetilde{\mathbf{P}}(f,u)=-(L_{\eta_{f}}J,d_{A}(u+2\psi(J\eta_{f}))+i_{\eta_{f}}F_{A},d_{A}(i_{\eta_{f}}\psi)+[\psi,(u+2\psi(J\eta_{f})]),$
which we regard as a modified infinitesimal action (see (4.10)). Observe that, by definition, $\operatorname{Im}\widetilde{\mathbf{P}}=\operatorname{Im}\mathbf{P}$.
Lemma 4.6.
The following operator provides a formal adjoint of $\widetilde{\mathbf{P}}$ for the pairings (4.8) and (3.18)
$\widetilde{\mathbf{P}}^{*}=\delta\mu\circ\mathbb{J}\colon\mathcal{S}^{1}\to\mathcal{S}^{0}.$
More explicitly, setting $\widetilde{\mathbf{P}}^{*}=\widetilde{\mathbf{P}}^{*}_{1}\oplus\widetilde{\mathbf{P}}^{*}_{2}$, we have:
$\widetilde{\mathbf{P}}^{*}_{1}(\dot{J},a,\dot{\psi})$ $=-\varepsilon\delta S(J\dot{J})+\alpha*d\left(B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)-B(\Lambda_{\omega}F_{A},*a)+B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right),$
$\widetilde{\mathbf{P}}^{*}_{2}(\dot{J},a,\dot{\psi})$ $=\alpha*\left(d_{A}(Ja)-[\dot{\psi},J\psi]-d_{A}(\psi(J\dot{J}))\right).$
Proof.
The proof follows from a straightforward calculation using the formal properties of the moment map in Proposition 3.9. Let $(f,u)\in\mathcal{S}^{0}$ and define
$\widetilde{\zeta}=u+2\psi(J\eta_{f})+A^{\perp}(\eta_{f}).$
Notice that
(4.10) $\widetilde{\zeta}\cdot(J,A,\Psi)=\widetilde{\mathbf{P}}(f,u).$
Then, setting $v=(\dot{J},a,\dot{\psi})$, we have
$\begin{split}\langle\widetilde{\mathbf{P}}^{*}v,(f,u)\rangle=\langle d\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}(\mathbb{J}v),\zeta\rangle=\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(\zeta\cdot(J,A,\psi),\mathbb{J}v)=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(\widetilde{\mathbf{P}}(f,u),v)=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,\widetilde{\mathbf{P}}(f,u)).\end{split}$
Consider now the differential operator
[…]
The key condition on the solution $(J,A,\psi)$ of (4.3) which we need to assume in order to construct the complex structure on the moduli space is the vanishing of the kernel of $\mathcal{L}$. Notice that, unlike in the standard cases in gauge theory in which the parameter space metric is positive definite, $\operatorname{ker}\mathcal{L}$ does not relate in general to automorphisms of the triple $(J,\psi,A)$, but rather to null vectors with respect to $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$. We build on the following technical result, which shows in particular that the subspace of null vectors in $\operatorname{Im}\mathbf{P}$ is finite-dimensional. Using the $L^{2}$ norm (4.8), we extend the domain of $\mathcal{L}$ to an appropriate Sobolev completion.
Proposition 4.7.
The operator $\mathcal{L}$ is Fredholm with zero index. Furthermore, elements $(f,u)\in\operatorname{ker}\mathcal{L}$ are smooth.
Proof.
We observe that $\widetilde{\mathbf{P}}=\mathbf{P}\circ\mathbf{T}$, where $\mathbf{T}\colon\mathcal{S}^{0}\to\mathcal{S}^{0}$ is the invertible operator
(4.12) $\mathbf{T}(f,u)=(f,u+2\psi(J\eta_{f})).$
Consider the operator $\widetilde{\mathcal{L}}:=\widetilde{\mathbf{P}}^{*}\circ\mathbf{P}\colon\mathcal{S}^{0}\to\mathcal{S}^{0}$, explicitly given by
$\widetilde{\mathcal{L}}_{1}(f,u)$ $=\varepsilon\delta S(JL_{\eta_{f}}J)-\alpha*d\left(B\left(\Lambda_{\omega}([F_{A},i_{\eta_{f}}\psi]+d_{A}[\psi,u]),*\psi\right)\right)$
$+\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*(d_{A}u+i_{\eta_{f}}F_{A})\right)-B\left(\Lambda_{\omega}F_{A},\psi(JL_{\eta_{f}}J)\right)\right),$
$\widetilde{\mathcal{L}}_{2}(f,u)$ $=-\alpha*\left(d_{A}(J(d_{A}u+i_{\eta_{f}}F_{A}))-[(d_{A}(i_{\eta_{f}}\psi)+[\psi,u]),J\psi]-d_{A}(\psi(JL_{\eta_{f}}J))\right).$
Then, the tuples $\mathbf{t}=(4,4)$ and $\mathbf{s}=(0,2)$ form a system of orders for $\widetilde{\mathcal{L}}$, since
$o(\widetilde{\mathcal{L}}_{11})$ $=4\leqslant 4-0,$ $o(\widetilde{\mathcal{L}}_{12})$ $=2\leqslant 4-0,$
$o(\widetilde{\mathcal{L}}_{21})$ $=2\leqslant 4-2,$ $o(\widetilde{\mathcal{L}}_{22})$ $=2\leqslant 4-2.$
and the associated leading symbol is
$\sigma_{\widetilde{\mathcal{L}}_{1}}(v)(f,u)$ $=-\varepsilon*v\wedge v(J\omega^{-1}v)Jvf$
$=-\varepsilon|v|^{4}f,$
$\sigma_{\widetilde{\mathcal{L}}_{2}}(v)(f,u)$ $=-\alpha*\left(v\wedge J(v\otimes u+fi_{\omega^{-1}}F_{A})-fv\wedge i_{\omega^{-1}}\psi+fv\wedge\psi(Jv\otimes\omega^{-1}v)\right)$
$=-\alpha|v|^{2}u-\alpha*\left(v\wedge Ji_{\omega^{-1}}F_{A}+v\wedge i_{\omega^{-1}}\psi-v\wedge\psi(Jv\otimes\omega^{-1}v)\right)f,$
which is clearly invertible. Consequently, by the general theory of linear multi-degree elliptic differential operators (see [LM]) it follows that $\widetilde{\mathcal{L}}$ is Fredholm and elements in $\operatorname{ker}\widetilde{\mathcal{L}}$ are smooth. The result follows now from $\operatorname{ker}\mathcal{L}=\mathbf{T}^{-1}\operatorname{ker}\widetilde{\mathcal{L}}$ and the fact that $\mathcal{L}$ is self-adjoint. ∎
Assuming that $\operatorname{ker}\mathcal{L}$ is trivial, in the next result we obtain a natural gauge fixing via a $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$-orthogonal decomposition
(4.13) $\mathcal{S}^{1}=\operatorname{Im}\;\widetilde{\mathbf{P}}\oplus(\operatorname{Im}\;\widetilde{\mathbf{P}})^{\perp_{\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}}}.$
Lemma 4.8.
Assume that $\operatorname{ker}\mathcal{L}=\{0\}$. Then, there exists an orthogonal decomposition (5.20) for the pairing $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$. Consequently, for any element $v\in\mathcal{S}^{1}$ there exists a unique $\Pi v\in\operatorname{Im}\;\widetilde{\mathbf{P}}$ such that $(\dot{J},a,\dot{\psi})=v-\Pi v$ solves the linear equations
$\begin{split}\alpha d_{A}(Ja)-\alpha[\dot{\psi},J\psi]-\alpha d_{A}(\psi(J\dot{J}))&=0\\ \varepsilon\delta S(J\dot{J})+\alpha*d\left(B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)-B(\Lambda_{\omega}F_{A},*a)+B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right)&=0.\end{split}$
Proof.
Notice first that from the non-degeneracy of $B$, the pairing given in (4.8) is non-degenerate. Thus
$\ker\,{\bf{\widetilde{P}}}^{*}=(\operatorname{Im}\,\widetilde{\mathbf{P}})^{\perp_{\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}}}.$
If $v\in\operatorname{Im}\;\widetilde{\mathbf{P}}\cap(\operatorname{Im}\;\widetilde{\mathbf{P}})^{\perp_{\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}}}$, then $v=\widetilde{\mathbf{P}}(y)$ for $y\in\mathcal{S}^{0}$. But then $\widetilde{\mathbf{P}}^{*}\circ\widetilde{\mathbf{P}}(y)=0$ and, by $\operatorname{ker}\mathcal{L}=\{0\}$, $v=0$. Thus
(4.14) $\operatorname{Im}\;\widetilde{\mathbf{P}}\cap(\operatorname{Im}\;\widetilde{\mathbf{P}})^{\perp_{\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}}}=\{0\}.$
Let $v\in\mathcal{S}^{1}$. The condition
$v-{\bf\widetilde{P}}(y)\in(\operatorname{Im}\;\widetilde{\mathbf{P}})^{\perp_{\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}}}$
for some $y\in\mathcal{S}^{0}$ is equivalent to
(4.15) ${\bf\widetilde{P}}^{*}(v)={\bf{\widetilde{P}}}^{*}\circ\widetilde{\mathbf{P}}(y).$
But by Proposition 4.7 and condition $\operatorname{ker}\mathcal{L}=\{0\}$, $\widetilde{\mathbf{P}}^{*}\circ{\bf\widetilde{P}}$ is surjective. Then, by elliptic regularity, one can solve (4.15) for $y\in\mathcal{S}^{0}$. The orthogonal decomposition follows. The last statement of the lemma comes from the expression of $\bf{\widetilde{P}}^{*}$ in Lemma 5.14. ∎
The above lemma suggests to define the space of harmonic representatives of the complex (4.7), as follows:
$\mathcal{H}^{1}(\mathcal{S}^{*})=\ker\bf{L}^{\varepsilon}_{\alpha}\cap\ker{\bf\widetilde{P}}^{*}$.
Our next result provides our gauge fixing mechanism for the linearization of coupled harmonic equations (4.3).
Proposition 4.9.
Assume $\operatorname{ker}\mathcal{L}=\{0\}$ and $\alpha\neq 0$. Then, the inclusion $\mathcal{H}^{1}(\mathcal{S}^{*})\subset\ker\bf{L}^{\varepsilon}_{\alpha}$ induces an isomorphism
$\mathcal{H}^{1}(\mathcal{S}^{*})\simeq H^{1}(\mathcal{S}^{*}).$
More precisely, any class in the cohomology $H^{1}(\mathcal{S}^{*})$ of the complex (4.7) admits a unique representative $(\dot{J},a,\dot{\psi})$ solving the linear equations
(4.16) $\begin{split}d_{A}a-[\dot{\psi},\psi]&=0,\\ d_{A}\dot{\psi}+[a,\psi]&=0,\\ d_{A}(J\dot{\psi})+[a,J\psi]-d_{A}(\psi(\dot{J}))&=0,\\ d_{A}(Ja)-[\dot{\psi},J\psi]-d_{A}(\psi(J\dot{J}))&=0\\ \varepsilon\delta S(\dot{J})-\alpha*d\left(B(\Lambda_{\omega}d_{A}a,*\psi)+B(\Lambda_{\omega}F_{A},*\dot{\psi})-B(\Lambda_{\omega}F_{A},\psi(\dot{J}))\right)&=0,\\ \varepsilon\delta S(J\dot{J})-\alpha*d\left(-B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)+B(\Lambda_{\omega}F_{A},*a)-B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right)&=0,\end{split}$
Proof.
The correspondence between $H^{1}(\mathcal{S}^{*})$ and the space of solutions of (4.16) follows from Lemma 4.4 and Lemma 4.8. ∎
We are ready to prove our main result, which shows that the gauge fixing in Proposition 4.9 enables us to descend the complex structure $\mathbb{J}$ in $\curly{X}$ and the symmetric tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$, to an open subset of the moduli space $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$, via the symplectic reduction in Proposition 3.9. Define
$\mathcal{U}^{*}_{\alpha,\varepsilon}=\{[(J,A,\psi)]\;|\;\operatorname{ker}\mathcal{L}=\{0\}\}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}.$
Theorem 4.10.
The set $\mathcal{U}^{*}_{\alpha,\varepsilon}$ is open in $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$, and for any smooth point $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$ the tangent space to $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ at $(J,A,\psi)]$, identified with the space of solutions of the gauge fixed linear equations (4.16), inherits a complex structure $\mathbb{J}$ and a symmetric tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ such that $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(\mathbb{J},)$, given respectively by (3.7) and (3.18), and where $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ stands for the restriction of (3.17). Furthermore,
  1. (1)
    if $\varepsilon=1$ the tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is possibly degenerate,
  2. (2)
    if $\varepsilon=-1$ the tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is non-degenerate, and defines a pseudo-Kähler structure on the moduli space.
In either case, $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ admits a global Kähler potential, that is, $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}=dd^{c}_{\mathbb{J}}\Phi$, where
$\Phi=\varepsilon\nu{J}+\frac{\alpha}{2}\|\psi\|^{2}_{L^{2}}$
and $\nu{J}$ is induced by the global Kähler potential in the space of complex structures compatible with the orientation $\curly{J}$ (see [F, Section 4]).
Proof.
The fact that $\mathcal{U}^{*}_{\alpha,\varepsilon}$ is open follows from upper semicontinuity of the dimension of the kernel for elliptic operators. Given now a smooth point $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$, the tangent space is identified with $H^{1}(\mathcal{S}^{*})$ and inherits a complex structure by Proposition 4.9, using that $\mathbb{J}$ in (3.7) preserves (4.16). The existence the of the symmetric tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is a direct consequence of Lemma 3.4 and Proposition 3.9, while the listed signature properties follow from Lemma 3.8. The formula for the Kähler potential is a direct consequence of (3.13), while a explicit formula for $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ follows from Lemma 3.4. ∎
To finish this section, we provide an explicit formula for the (pre)symplectic structure $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$. The proof is straightforward from the previous discussion and Lemma 3.4.
Corollary 4.11.
Let $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$ be a smooth point and take $v_{1},v_{2}$ tangent vectors of $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ at $(J,A,\psi)]$, identified with solutions $(\dot{J}_{j},a_{j},\dot{\psi}_{j})$ of the gauge fixed linear equations (4.16). Then, one has
(4.17) $\begin{split}\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2}))\\ &+\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge J(\psi(\dot{J}_{2}))),\end{split}$
4.3. Comparison with $\boldsymbol{\mathcal{U}^{\operatorname{Flat}}(G)}$ and existence
This aim of this section is twofold. Firstly, we establish a comparison between the moduli spaces $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ and the universal moduli space of flat $G$-connections $\mathcal{U}^{\operatorname{Flat}}(G)$ constructed in Section 4.1. As we will see, for any $\alpha>0$ and $\varepsilon\neq 0$, Theorem 4.10 induces a natural holomorphic map
$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)$
and hence a holomorphic map into Teichmuller space $\mathcal{T}$. Secondly, we will prove that for genus of the surface $\Sigma$ bigger than zero, the open $\mathcal{U}^{*}_{\alpha,\varepsilon}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ is non-empty for sufficiently small values of $\alpha$.
Consider a point $(J,D)]\in\mathcal{U}^{\operatorname{Flat}}(G)$, regarded as the $\widetilde{\curly{G}}-$orbit of $(J,A,\psi)\in\curly{X}^{*}$ solving the equations
(4.18) $\begin{split}F_{A}-\frac{1}{2}[\psi,\psi]&=0,\\ d_{A}\psi&=0.\end{split}$
The tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{\operatorname{Flat}}(G)$ can be formally identified with the cohomology of the complex of linear differential operators
(4.19) ${(\mathcal{C}^{*})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{C}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}^{c}}$${\mathcal{C}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L^{c}}}$${\mathcal{C}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
where we have
$\mathcal{C}^{0}=\operatorname{Lie}\widetilde{\curly{G}}\cong\Omega^{0}(T\Sigma)\oplus\Omega^{0}(X,E_{G}(\mathfrak{g})),\qquad\mathcal{C}^{1}=\mathcal{S}^{1},\qquad\mathcal{C}^{2}=\Omega^{2}(X,E_{G}(\mathfrak{g}))$
and
$\mathbf{P}^{c}(y,u_{0}+iu_{1})$ $=-\left(L_{y}J,d_{A}u_{0}+i_{y}F_{A}-[\psi,u_{1}],d_{A}u_{1}+d_{A}(\psi(y))+[\psi,u_{0}]\right),$
$\mathbf{L}^{c}(\dot{J},a,\dot{\psi})$ $=d_{A}a-[\dot{\psi},\psi]+i\left(d_{A}\dot{\psi}+[a,\psi]\right).$
Lemma 4.12.
The sequence (5.24) is an elliptic complex of degree-one linear differential operators. Consequently, the cohomology groups $H^{j}(\mathcal{C}^{*})$, with $j=0,1,2$, are finite-dimensional.
Proof.
The leading symbol of $\sigma_{\mathbf{P}^{c}}$ is
$\sigma_{\mathbf{P}^{c}}(v)(y,u)=(-Jv\otimes y,-v\otimes u_{0},-v\otimes u_{1}-v\otimes\psi(y)).$
Similarly, the associated leading symbol of ${\bf L}^{c}$ is
$\sigma_{{\bf L}^{c}}(v)(\dot{J},a,\dot{\psi})=v\wedge(a+i\dot{\psi}).$
The symbol $\sigma_{\mathbf{P}^{c}}(v)$ is obviously injective, while $\sigma_{{\bf L}^{c}}(v)$ is surjective for dimensional reasons. Assume now that $\sigma_{{\bf L}^{c}}(v)(\dot{J},a,\dot{\psi})=0$. Then, $a+i\dot{\psi}=v\otimes(u_{0}+iu_{1}^{\prime})$. Again, by dimensional reasons $\dot{J}=-Jv\otimes y$ for some vector $y$, and hence
$(\dot{J},a,\dot{\psi})=\sigma_{\mathbf{P}^{c}}(v)(y,u_{0}+i(u_{1}^{\prime}-\psi(y))).$
Applying Theorem 2.5, a solution $(J,A,\psi)$ of the coupled harmonic equations (4.3) induces a reductive flat $G$-connection $D=A+i\psi\in{\curly D}^{*}$. This fact, jointly with the natural inclusion $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$, leads to a continuous map
(4.20) $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)\colon[(J,A,\psi)]\mapsto[(J,A+i\Psi)].$
Building on Theorem 4.10, our next goal is to prove that this induces a holomorphic map $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)$.
Lemma 4.13.
Let $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$. Then, (5.25) induces a complex linear map
(4.21) ${H^{1}(\mathcal{S}^{*})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathcal{C}^{*}).}$
where the complex structure on $H^{1}(\mathcal{S}^{*})$ is the one induced by Proposition 4.9.
Proof.
Using that $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$ we can identify
$H^{1}(\mathcal{S}^{*})\simeq\mathcal{H}^{1}(\mathcal{S}^{*}),$
where the right hand side is given by solutions of the gauge-fixed system of linear equations (4.16). Then, the map
$\mathcal{H}^{1}(\mathcal{S}^{*})\to H^{1}(\mathcal{C}^{*})\colon[(\dot{J},a,\dot{\psi})]\mapsto[(\dot{J},a+i\dot{\psi})]$
is complex $\mathbb{C}$-linear, since both complex structures are induced by $\mathbb{I}$ in (3.7). ∎
To finish this section, we address the question of non-emptyness of the moduli space $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$, in genus $g(\Sigma)\geqslant 2$. We change our perspective on the equations (4.3): we fix a compact Riemann surface $X=(\Sigma,J)$ and consider a flat $G$-bundle $(E_{G},D)$. In this setup, consider the coupled equations
(4.22) $\begin{split}d_{A_{h}}^{*}\psi_{h}&=0\\ \varepsilon S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi_{h}\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V}\end{split}$
for pairs $(g,h)$, where $g$ is a Kähler metric on $X$ with total volume $V$ and $h\in\Omega^{0}(E_{G}(G/K))$ is a reduction of structure group of $E_{G}$ to $K$. Here, we denote
$D=A_{h}+i\psi_{h}$
the natural decomposition of $D$ with respect to $h$. A solution $(g,h)$ of the equations (4.22) determines then a solution of (4.3), for the same value of parameter $\varepsilon\in\{-1,1\}$ and coupling constant $\alpha$, given by the triple $(J,A_{h},\psi_{h})$.
Theorem 4.14.
Let $(E_{G},D)$ be an irreducible flat $G$-bundle over a compact Riemann surface $X=(\Sigma,J)$ with genus $g(\Sigma)\geqslant 2$. Then, for any fixed total volume $V>0$ and parameter $\varepsilon\in\{-1,1\}$, there exists $\alpha_{0}>0$ such that for any $0<\alpha<\alpha_{0}$ there exists a solution $(g_{\alpha},h_{\alpha})$ of the equations (4.22) with
$(J,A_{h_{\alpha}},\psi_{h_{\alpha}})]\in\mathcal{U}^{*}_{\alpha,\varepsilon}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}.$
Furthermore, the induced map $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)$ is holomorphic.
Proof.
The first part of the proof follows by a perturbation argument. Let $g$ be the unique constant scalar curvature Kähler metric on $X$ with total volume $V$, and $h$ the unique harmonic reduction on $(E_{G},D)$ (which exists by Theorem 2.5). Consider the non-linear operator
$\mathbf{Q}_{\alpha}\colon C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(X,E_{K}(\mathfrak{k}))\to C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{2}(X,E_{K}(\mathfrak{k}))$
defined by
$\mathbf{Q}_{\alpha}(f,u)=(S_{g_{f}}-\varepsilon^{-1}\alpha*d\left(B\left(\Lambda_{\omega_{f}}F_{A_{h_{u}}},*\psi_{h_{u}}\right)\right)-2\pi\chi(\Sigma)V^{-1},\operatorname{Ad}(e^{-iu})d_{A_{h_{u}}}J\psi_{h_{u}})$
where
$\omega_{f}=\omega+2i\partial\bar{\partial}f,\qquad h_{u}=e^{iu}h.$
Notice that $\mathbf{Q}_{\alpha}$ is a multidegree elliptic differential operator, $\mathbf{Q}_{0}(0,0)=0$, and that the linearization of $\mathbf{Q}_{0}$ at $(0,0)$ is
$\delta_{0}\mathbf{Q}_{0}(f,u)=-(\delta S(JL_{\eta_{f}}J),d_{A}(Jd_{A_{h}}u)-[[\psi_{h},u],J\psi]).$
The operator $f\mapsto\delta S(JL_{\eta_{f}}J)$ is the (real) Lichnerowicz operator acting on functions: this is an elliptic self-adjoint semipositive differential operator of order 4, whose kernel is given by the Hamiltonian functions of Killing Hamiltonian vector fields. In particular, by our assumption $g(\Sigma)\geqslant 2$, it is invertible. On the other hand, the operator
$u\mapsto*\left(d_{A}(Jd_{A_{h}}u)-[[\psi_{h},u],J\psi]\right)$
is self-adjoint and has kernel given by the infinitesimal unitary gauge transformations preserving $D$. Hence, since $D$ is irreducible by hypothesis, this operator is also invertible. By an standard implicit function theorem argument, taking Sobolev completions in the domain and target of $\mathbf{Q}_{\alpha}$, it follows that there exists $\alpha_{0}>0$ such that for any $0<\alpha<\alpha_{0}$ there exists a smooth solution $(g_{\alpha},h_{\alpha})$ of the equations (4.22), that is, with
$(J,A_{h_{\alpha}},\psi_{h_{\alpha}})]\in\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}.$
To finish the first part of the proof, we need to show that $(J,A_{h_{\alpha}},\psi_{h_{\alpha}})]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$ for sufficiently small $\alpha$. Following the notation in Lemma 5.14, for $(g_{\alpha},A_{h_{\alpha}},\psi_{h_{\alpha}})$ define the one-parameter family of multidegree differential operators
$\mathcal{N}^{\alpha}=(\varepsilon^{-1}\widetilde{\mathbf{P}}_{1}^{*},\alpha^{-1}\widetilde{\mathbf{P}}_{2}^{*})\circ\widetilde{\mathbf{P}}\colon\mathcal{S}^{0}\to\mathcal{S}^{0}.$
Notice that $\mathcal{N}^{0}$ is well-defined, and furthermore
$\ker\mathcal{N}^{\alpha}=\ker\mathcal{L}_{(g_{\alpha},A_{h_{\alpha}},\psi_{h_{\alpha}})}$
for any $\alpha\neq 0$, where $\mathcal{L}_{(g_{\alpha},A_{h_{\alpha}},\psi_{h_{\alpha}})}=\widetilde{\mathbf{P}}^{*}\circ\widetilde{\mathbf{P}}$ is the operator associated to the solution $(J,A_{h_{\alpha}},\psi_{h_{\alpha}})$ of (4.3) with respect to the symplectic structure $\omega_{\alpha}=g_{\alpha}(J,)$. Decomposing
$\mathcal{N}^{0}=\mathcal{N}^{0}_{0}\oplus\mathcal{N}^{0}_{1}$
it follows that
$\mathcal{N}^{0}_{0}(f,u)=\delta S(JL_{\eta_{f}}J).$
Hence, arguing as before, $(f,u)\in\ker\mathcal{N}^{0}$ implies that $f=0$, and hence it follows that
$\mathcal{N}^{0}(f,u)=-(0,*d_{A_{h}}(Jd_{A_{h}}u)-*[[\psi_{h},u],J\psi_{h}])=0.$
Again, by irreducibility of the $G$-connection $D$, this implies $u=0$, and hence we conclude that $\ker\mathcal{N}^{0}=\{0\}$. Arguing as in the proof of Proposition 4.7 and Theorem 4.10, $\ker\mathcal{N}^{\alpha}$ is upper semicontinuous, and hence for sufficiently small $\alpha$ we have
$\ker\mathcal{N}^{\alpha}=\{0\}.$
The last part of the statement follows from that fact that (5.25) is induced by a $C^{\infty}$-Frechet map, combined with Lemma 5.21.
Remark 4.15.
We observe that, even though the proof of Theorem 4.10 works in the case $g(\Sigma)=1$, the hypothesis of existence of an irreducible flat $G$-connection is never satisfied in this case [FGN]. We thank Emilio Franco for this observation.
5. Universal G-Higgs bundle moduli
5.1. The universal Higgs field
We fix a smooth oriented compact surface $\Sigma$. We also consider a connected semisimple complex Lie group $G$, with Lie algebra $\mathfrak{g}$, and fix an antiholomorphic involution $\tau$ of $G$ defining a maximal compact subgroup $K:=G^{\tau}\subset G$, with Lie algebra $\mathfrak{k}$.
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$. Consider $h\in\Omega^{0}(E_{G}(G/K))$ a reduction of structure group of $E_{G}$ to $K$, and let $E_{K}$ be the corresponding principal $K$-bundle. Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation. Consider the space
$\curly{X}=\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k})).$
In this section we are interested in the geometry of the submanifold of $\curly{X}$ given by universal Higgs fields,
$\curly{X}^{Higgs}=\{(J,A,\psi)\in\curly{X}\;|\;\bar{\partial}_{J,A}\varphi=0\}$
where the Higgs field
$\varphi(J,A,\psi):=\psi^{1,0_{J}}$
is regarded as a $\Omega^{1}(\Sigma,E_{G}(\mathfrak{g}))$-valued function on the parameter space $\curly{X}$. Our first goal is to prove that $\curly{X}$ admits a complex structure which is compatible with the integrability condition $\bar{\partial}_{J,A}\varphi=0$, making $\curly{X}^{Higgs}\subset\curly{X}$ a complex submanifold. For this, consider the complex structure
(5.1) $\mathbb{I}(\dot{J},a,\dot{\psi})=(J\dot{J},Ja,-J\dot{\psi}+\psi(\dot{J})),$
for $(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$. Notice that this complex structure corresponds to the natural fibrewise cotangent complex structure in the holomorphic fibration $\curly{J}\times{\curly A}\to\curly{J}$. In particular, the map $\pi_{1}\colon\curly{X}\to\curly{J}$ is holomorphic, and the complex structure along the fibres coincides with the $J$-dependent complex structure $\mathbf{I}$ in (2.8). Using the Chern-Singer correspondence, it is not difficult to give holomorphic coordinates on the space $\curly{J}\times{\curly A}\cong\curly{J}\times{\curly C}$, and hence on $(\curly{X},\mathbb{I})$. In the next result we give an independent proof of the vanishing of the Nijenhuis tensor $N_{\mathbb{I}}$ of $\mathbb{I}$.
Lemma 5.1.
The complex structure $\mathbb{I}$ is formally integrable, that is, its Nijenhuis tensor vanishes
$N_{\mathbb{I}}=0.$
Proof.
The complex structures on the fibers and base of the fibration $(\curly{X},\mathbb{I})\to(\curly{J},\mathbb{J}{J})$ are well-known to be integrable. Hence, it suffices to calculate the mixed component of the Nijenhuis tensor
$N_{\mathbb{I}}(v_{1},v_{2})$ $=[\mathbb{I}v_{1},\mathbb{I}v_{2}]-\mathbb{I}[\mathbb{I}v_{1},v_{2}]-\mathbb{I}[v_{1},\mathbb{I}v_{2}]-[v_{1},v_{2}],$
that is, for $v_{1}=(\dot{J},0,0)$, and $v_{2}=(0,a,\dot{\psi})$. With this choice, the last term vanishes identically, and we have
$N_{\mathbb{I}}(v_{1},v_{2})$ $=[(J\dot{J},0,0),(0,Ja,-J\dot{\psi})]-\mathbb{I}[(J\dot{J},0,0),(0,a,\dot{\psi})]$
$-\mathbb{I}[(\dot{J},0,0),(0,Ja,-J\dot{\psi})],$
$=(0,-a(J\dot{J}),\dot{\psi}(J\dot{J}))-\mathbb{I}(0,-a(\dot{J}),\dot{\psi}(\dot{J}))$
$=(0,-a(J\dot{J}+\dot{J}J),\dot{\psi}(J\dot{J}+\dot{J}J))=0.$
In our next result we show that $\curly{X}^{Higgs}\subset(\curly{X},\mathbb{I})$ is fomally a complex submanifold.
Proposition 5.2.
The tangent space of $\curly{X}^{Higgs}$ at $(J,A,\psi)$, is given by triples $(\dot{J},a,\dot{\psi})\in T_{J}\curly{J}\oplus\Omega^{1}(X,E_{K}(\mathfrak{k}))\oplus\Omega^{1}(X,E_{K}(\mathfrak{k}))$ satisfying
(5.2) $d_{A}\dot{\psi}^{1,0_{J}}+[a,\varphi]-\frac{i}{2}d_{A}(\psi(\dot{J}))=0.$
Consequently, $\curly{X}^{Higgs}$ is fomally a complex submanifold of $(\curly{X},\mathbb{I})$.
Proof.
By dimensional reasons, the integrability condition $\bar{\partial}_{J,A}\psi^{1,0_{J}}=0$ is equivalent to
$\frac{1}{2}d_{A}\left(\psi(\operatorname{Id}-iJ)\right)=0.$
Taking variations on the parameters, we obtain (5.2). Now, given $(\dot{J},a,\dot{\psi})$ solving (5.2), we have
$d_{A}(-J\dot{\psi}+\psi(\dot{J}))^{1,0_{J}}+[Ja,\varphi]-\frac{i}{2}d_{A}\left(\psi(J\dot{J})\right)$ $=id_{A}\dot{\psi}^{1,0_{J}}+i[a,\varphi]+d_{A}\left(\psi^{0,1_{J}}(\dot{J})\right)$
$-\frac{1}{2}d_{A}\left(\psi^{0,1_{J}}(\dot{J})\right)+\frac{1}{2}d_{A}\left(\psi^{1,0_{J}}(\dot{J})\right)$
$=i\left(d_{A}\dot{\psi}^{1,0_{J}}+[a,\varphi]-\frac{i}{2}d_{A}\left(\psi(\dot{J})\right)\right)=0$
and hence the tangent space to $\curly{X}^{Higgs}$ is preserved by $\mathbb{I}$. ∎
Our next goal is to define a universal moduli space of Higgs bundles, with a holomorphic fibration structure over Teichmüller space. The first step is to show that the integrability condition for $\varphi$ is compatible with the group $\widetilde{\curly{G}}$ defined in (4.1). We observe first that there is a holomorhic action of $\widetilde{\curly{G}}$ on $\curly{J}\times{\curly A}$, given by
$g(J,A)=(p(g)_{*}J,A_{g}),$
where $A_{g}$ is the defined via the Chern-Singer correspondence
$A_{g}=A_{h,g_{*}\bar{\partial}_{J,A}}.$
This action extends to the fibrewise cotangent bundle, defining a holomorphic $\widetilde{\curly{G}}-$action on $(\curly{X},\mathbb{I})$ and preserving $\curly{X}^{Higgs}$. Let us denote $\curly{X}^{ps}\subset\curly{X}^{Higgs}$ the subset of triples $(J,A,\psi)$ such that $(E,\varphi)$ is a polystable $G$-Higgs bundle over the compact Riemann surface $X=(\Sigma,J)$, where $E$ denotes the holomorphic principal $G$-bundle over $X$ corresponding to the connection $A$. The $\widetilde{\curly{G}}-$action on $\curly{X}$ preserves the subsets
$\curly{X}^{ps}\subset\curly{X}^{Higgs}\subset\curly{X}$
and we define the universal moduli space of $G$-Higgs bundles on $E_{G}$ as the quotient
$\mathcal{U}^{Higgs}(G):=\curly{X}^{ps}/\widetilde{\curly{G}}$.
By construction, $\mathcal{U}^{Higgs}(G)$ fibers over the Teichmüller space $\mathcal{T}$ of complex structures modulo diffeomorphisms isotopic to the identity
(5.3) $\mathcal{U}^{Higgs}(G)\to\mathcal{T}:=\curly{J}/\operatorname{Diff}_{0}(\Sigma).$
This fibration is naturally holomorphic with respect to the complex structure $\mathbb{I}$.
Remark 5.3.
Despite our efforts, we have not been able to find a complex symplectic interpretation of the universal moduli space of $G$-Higgs bundles. Further insight on this is provided in ongoing work by N. Hitchin [Hi4].
5.2. Kähler fibration for the universal Higgs field
Consider the structure of Kähler fibration $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ on $(\curly{X},\mathbb{I})\to\curly{J}$ defined fibrewise by the symplectic structure $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ on ${\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$. More explicitly,
(5.4) $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2}))=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2}).$
The aim of this section is to prove that such structure admits a Kähler Ehresmann connection compatible with Hitchin’s equations (2.15). The situation here is opposite to that on Section 3.2: the holomorphic fibration defined by $\mathbb{I}$ is non-trivial, whereas $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ defines a trivial symplectic fibration, constant along $\curly{X}$. Nonetheless, the combination of these two structures defines a non-trivial structure of Kähler fibration on $\curly{X}\to\curly{J}$, different from the one considered in Section 3.2.
In order to achieve our goal in this section, we will apply Theorem 3.2. As we have seen in Lemma 5.2, the integrability condition
$\bar{\partial}_{J,A}\psi^{1,0_{J}}=0$
cuts a holomorphic submanifold $\curly{X}^{Higgs}$ of $(\curly{X},\mathbb{I})$. Given that $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ is constant on $\curly{X}$, it provides a candidate for coupling two-form $\boldsymbol{\sigma}$ on $(\curly{X},\widehat{\boldsymbol{\omega}}_{\mathbf{I}})$, inducing the trivial Ehresmann connection on $\curly{X}\to\curly{J}$, regarded as the product $\curly{X}=\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k})$. However, one can readily check that the horizontal subspace of this connection is not tangent to the submanifold $\curly{X}^{Higgs}$, and hence is not appropiate for our goals. Motivated by this, we consider the Ehresmann connection $\Gamma^{\mathbb{I}}$ with horizontal subspace
$H^{\mathbb{I}}=\Big\{\left(\dot{J},0,-\tfrac{1}{2}\psi(J\dot{J})\right)\;|\;\dot{J}\in T_{J}\curly{J}\Big\}.$
Lemma 5.4.
The Ehresmann connection $\Gamma^{\mathbb{I}}$ is preserved by $\mathbb{I}$, and
$\Gamma^{\mathbb{I}}\circ\mathbb{I}=\mathbf{I}\circ\Gamma^{\mathbb{I}},$
where $\mathbf{I}$ denotes the complex structure along the fibres. Furthermore, its curvature $F_{\mathbb{I}}\in\Omega^{2}(\curly{X},V\curly{X})$ is of type $(1,1)$ and given explicitly by
(5.5) $\begin{split}(F_{\mathbb{I}})_{|(J,A,\psi)}(v_{1},v_{2})&=\frac{1}{4}\left(0,0,\psi([\dot{J}_{1},\dot{J}_{2}])\right),\end{split}$
for any pair of horizontal vector fields $v_{1},v_{2}\in H^{\sigma}$ covering $\dot{J}_{1},\dot{J}_{2}$, respectively.
Proof.
We take $v=(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$ and notice that
$\Gamma^{\mathbb{I}}(\dot{J},a,\dot{\psi})=\left(0,a,\dot{\psi}+\tfrac{1}{2}\psi(J\dot{J})\right),$
which implies that
$\Gamma^{\mathbb{I}}(\mathbb{I}v)$ $=\left(0,Ja,-J\dot{\psi}+\psi(\dot{J})-\tfrac{1}{2}\psi(\dot{J})\right)=\mathbf{I}\left(0,a,\dot{\psi}-\tfrac{1}{2}\psi(\dot{J}J)\right)=\mathbf{I}\Gamma^{\mathbb{I}}(v).$
In particular, this implies that $\Gamma^{\mathbb{I}}$ is preserved by $\mathbb{I}$. To calculate the formula for the curvature, we choose horizontal vectors
$v_{j}=\left(\dot{J},0,-\tfrac{1}{2}\psi(J\dot{J}_{j})\right)\in H^{\mathbb{I}}_{(J,A,\psi)},$
with $j=1,2$, and vector fields $Y_{j}$ such that $Y_{j}(J,a,\psi)=v_{j}$. By taking coordinates in $\curly{J}$, we can choose $Y_{j}$ such that the $T\curly{J}$ component of the Lie bracket $Y_{1},Y_{2}]$ vanishes at $(J,A,\psi)$. Then, we calculate
$F_{\mathbb{I}}(v_{1},v_{2})$ $=-[Y_{1},Y_{2}]_{|(J,A,\psi)}$
$=-\frac{1}{2}\frac{d}{dt}_{|t=0}\left(0,0,-\psi_{t,1}(J_{t,1}Y_{2})+\psi_{t,2}(J_{t,2}Y_{1}))\right)$
$=-\frac{1}{2}\left(0,0,\frac{1}{2}\psi(J\dot{J}_{1}J\dot{J}_{2})-\psi(\dot{J}_{1}\dot{J}_{2})-\frac{1}{2}\psi(J\dot{J}_{2}J\dot{J}_{1})+\psi(\dot{J}_{2}\dot{J}_{1})\right)$
$=\frac{1}{4}\left(0,0,\psi([\dot{J}_{1},\dot{J}_{2}])\right).$
The fact that $F_{\mathbb{I}}$ is of type $(1,1)$ follows from this formula and the definition of $\mathbb{I}$. ∎
As a direct consequence of the equality $\Gamma^{\mathbb{I}}\circ\mathbb{I}=\mathbf{I}\circ\Gamma^{\mathbb{I}}$, it follows that $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}(\Gamma^{\mathbb{I}},\Gamma^{\mathbb{I}})$ defines a $(1,1)$-form on $(\curly{X},\mathbb{I})$ such that the associated Ehresmann connection equals $\Gamma^{\mathbb{I}}$. In the next result we calculate an explicit formula for this $(1,1)$-form, the proof being straightforward and hence omitted.
Lemma 5.5.
Given $(J,A,\psi)\in\curly{X}$ and any pair of tangent vectors $v_{j}=(\dot{J}_{j},a_{j},\dot{\psi}_{j})\in T_{(J,A,\psi)}\curly{X}$, one has
(5.6) $\begin{split}\widehat{\boldsymbol{\omega}}_{\mathbf{I}}(\Gamma^{\mathbb{I}}v_{1},\Gamma^{\mathbb{I}}v_{2})&=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{1}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{1}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right)\end{split}$
Our next result proves that the Ehresmann connection $\Gamma^{\mathbb{I}}$ is Kähler, by calculation of an explicit coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$ via equation (3.3). Define a basic two-from $\mu\in\Omega^{2}(\curly{X})$ by
$\mu\left((\dot{J}_{1},a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2})\right)=\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right).$
Proposition 5.6.
The Ehresmann connection $\Gamma^{\mathbb{I}}$ is Kähler. Furthermore, the following formula gives a coupling form for $\Gamma^{\mathbb{I}}$
$\boldsymbol{\sigma}^{\mathbb{I}}:=\widehat{\boldsymbol{\omega}}_{\mathbf{I}}(\Gamma^{\mathbb{I}},\Gamma^{\mathbb{I}})-\mu.$
More explicitly,
(5.7) $\begin{split}\boldsymbol{\sigma}^{\mathbb{I}}_{|(J,A,\psi)}(v_{1},v_{2})&=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B\left(\left(\dot{\psi}_{1}+\tfrac{1}{2}\psi(J\dot{J}_{1})\right)\wedge\left(\dot{\psi}_{2}+\tfrac{1}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right)\\ &=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{1}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{1}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{1}{2}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right),\end{split}$
for $v_{j}=(\dot{J}_{j},a_{j},\dot{\psi}_{j})\in T_{(J,A,\psi)}\curly{X}$, $j=1,2$.
Proof.
Notice that $\mu$ is of type $(1,1)$ for $\mathbb{I}$, and hence so is $\boldsymbol{\sigma}^{\mathbb{I}}$. By Theorem 3.2, it suffices to prove that $\boldsymbol{\sigma}^{\mathbb{I}}$ is closed. Using the abstract formula for the coupling form in (3.3), this last fact reduces to the identity
$d\mu(v_{1},v_{2})=\widehat{\boldsymbol{\omega}}_{\mathbf{I}}(F_{\mathbb{I}}(v_{1},v_{2}),)$
evaluated on vertical vector fields. Now, by direct calculation
$d\mu(v_{1},v_{2})(0,a,\dot{\psi})$ $=\frac{1}{4}\int_{\Sigma}B\left(\dot{\psi}(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right)+\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\dot{\psi}(\dot{J}_{2})\right)$
$=\frac{1}{4}\int_{\Sigma}B\left(\dot{\psi}\wedge\psi([\dot{J}_{1},\dot{J}_{2}])\right)$
$=-\widehat{\boldsymbol{\omega}}_{\mathbf{I}}((0,a,\dot{\psi}),F_{\mathbb{I}}(v_{1},v_{2})).$
To finish this section, we provide a formula for the the symmetric tensor on $\curly{X}$ associated to the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$, explicitly given by
$\mathbf{g}_{\mathbb{I}}:=\boldsymbol{\sigma}^{\mathbb{I}}(,\mathbb{I}).$
By construction, this coincides with the flat hyperkähler metric (2.11) along the fibres of $\curly{X}\to\curly{J}$. Similarly as in Corollary 3.7, $\mathbf{g}_{\mathbb{I}}$ is negative semi-definite along the horizontal directions of the connection $\Gamma^{\mathbb{I}}$.
Corollary 5.7.
For any tangent vector $v=(\dot{J},a,\dot{\psi}),\in T_{(J,A,\psi)}\curly{X}$ one has
(5.8) $\begin{split}\mathbf{g}_{\mathbb{I}}(v,v)&=\int_{\Sigma}B(a\wedge Ja)+\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{1}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{1}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right)\end{split}$
In particular, given a horizontal vector field $v\in H^{\mathbb{I}}$ at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$, one has
(5.9) $\begin{split}\mathbf{g}_{\mathbb{I}}(v,v)=-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J(\psi(\dot{J}))\right).\end{split}$
Consequently, $\mathbf{g}_{\mathbb{I}}$ is negative semi-definite along the horizontal directions of $\Gamma^{\mathbb{I}}$.
Proof.
Formula (5.8) follows from (5.7), using that $\Gamma^{\mathbb{I}}\circ\mathbb{I}=\mathbf{I}\circ\Gamma^{\mathbb{I}}$ (see Lemma 5.4).
5.3. Hamiltonian action and coupled Hitchin’s equations
We follow the notation of the previous section. Our aim in this section is to investigate a natural Hamiltonian action for minimal couplings associated to the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$ for the connection $\Gamma^{\mathbb{I}}$, in the Kähler fibration $(\curly{X},\mathbb{I},\widehat{\boldsymbol{\omega}}_{\mathbf{I}})$ over the space of complex structure $\curly{J}$.
We fix a symplectic form $\omega$ on $\Sigma$, compatible with the orientation. Given $\alpha>0$ a real coupling constant and $\varepsilon\in\{-1,1\}$, the family of minimal coupling symplectic structures of our interest is defined by
(5.10) $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}=\varepsilon\boldsymbol{\omega}{J}+\alpha\boldsymbol{\sigma}^{\mathbb{I}},$
where $\boldsymbol{\sigma}^{\mathbb{I}}$ is the closed $(1,1)$-form in Proposition 5.6 and $\boldsymbol{\omega}{J}$ is as in (3.16). Notice that $\boldsymbol{\omega}{J}$ depends on the choice of $\omega$. By construction, $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$ is closed and of type $(1,1)$ with respect to the complex structure $\mathbb{I}$. Furthermore, along the fibres of $\curly{X}\to\curly{J}$ the $2$-form $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$ restricts to the Kähler structure $\alpha\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$. Consider the associated symmetric tensor
(5.11) $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}:=\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}(,\mathbb{I})=\varepsilon\mathbf{g}_{\curly{J}}+\alpha\mathbf{g}_{\mathbb{I}}.$
The following result is the analogue of Lemma 3.8 in the present setup. The proof is analogue, by application in this case of Corollary 5.7, and is ommitted.
Lemma 5.8.
Given a horizontal vector field $v\in H^{\mathbb{I}}$ at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$, one has
(5.12) $\begin{split}\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(v,v)=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}\dot{J})\omega-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J(\psi(\dot{J}))\right).\end{split}$
Consequently, for any $\alpha>0$ one has
  1. (1)
    For $\varepsilon=-1$, $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is negative definite along $H^{\mathbb{I}}$. In particular, $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,-1}$ is a non-degenerate symplectic structure.
  2. (2)
    For $\varepsilon=1$ and $\psi\neq 0$, $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ changes signature along the line $(J,A,\lambda\psi)\in\curly{X}$, for $\lambda\in\mathbb{R}$.
Our next goal is to prove that $(\curly{X},\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon})$ admits a Hamiltonian action by the extended gauge group $\widetilde{{\curly K}}$, determined by $\omega$ and the reduction $E_{K}\subset E_{G}$. For this, we will need some refined information about the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$. Consider the closed two form $\pi^{*}_{2}\omega_{\mathbf{I}}$ on $\curly{X}$, given by the pull-back of $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ via the canonical projection $\pi_{2}\colon\curly{X}\to{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$. Explicitly, for $v_{j}=(\dot{J}_{j},a_{j},\dot{\psi}_{j})\in T_{(J,A,\psi)}\curly{X}$, $j=1,2$, one has
$\pi^{*}_{2}\omega_{\mathbf{I}}(v_{1},v_{2})=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2}).$
The two-form $\pi^{*}_{2}\omega_{\mathbf{I}}$ restricts on the fibres to $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$, and hence, by the general theory of symplectic fibrations (see [GLS, Section 1.6]), one expects that there exists a one-form $\lambda\in\Omega^{1}(\curly{X})$ such that
(5.13) $\boldsymbol{\sigma}^{\mathbb{I}}=\pi^{*}_{2}\omega_{\mathbf{I}}+d\lambda.$
In the next result we prove that this is the case in the present situation, providing a refined version of the previous formula.
Lemma 5.9.
Consider the one-form $\lambda\in\Omega^{1}(\curly{X})$ defined by
$\lambda(v)=\frac{1}{4}\int_{\Sigma}B\left(\psi(J\dot{J})\wedge\psi\right)$
for $v=(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$. Then, formula (5.13) holds. Furthemore, considering the $L^{2}$-norm of the unitary Higgs field
$\nu(J,A,\psi)=\|\psi\|^{2}_{L^{2}}:=\int_{\Sigma}B(\psi\wedge J\psi),$
one has
(5.14) $\boldsymbol{\sigma}^{\mathbb{I}}=\pi^{*}_{2}\omega_{{\curly A}}+\tfrac{i}{2}\partial_{\mathbb{I}}\bar{\partial}_{\mathbb{I}}\nu$
where
$\omega_{{\curly A}}((a_{1},\dot{\psi}_{1}),(a_{2},\dot{\psi}_{2}))=\int_{\Sigma}B(a_{1}\wedge a_{2}).$
Proof.
By the proof of Lemma 3.4 and (5.1), we have
$d^{c}_{\mathbb{I}}\nu_{|(J,A,\psi)}(\dot{J},a,\dot{\psi})$ $=2\int_{\Sigma}B((J\dot{\psi}-\psi(\dot{J}))\wedge J\psi)+\int_{\Sigma}B(\psi\wedge\psi(J\dot{J}))$
$=2\int_{\Sigma}B(\dot{\psi}\wedge\psi)-2\int_{\Sigma}B(\psi(\dot{J}J)\wedge\psi)+\int_{\Sigma}B(\psi\wedge\psi(J\dot{J}))$
$=2\int_{\Sigma}B(\dot{\psi}\wedge\psi)+\int_{\Sigma}B(\psi(J\dot{J})\wedge\psi)$
$=2\int_{\Sigma}B(\dot{\psi}\wedge\psi)+4\lambda.$
We now calculate
$4d\lambda((\dot{J}_{1},a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))$ $=\int_{\Sigma}B(\dot{\psi}_{1}(J\dot{J}_{2})\wedge\psi)-\int_{\Sigma}B(\dot{\psi}_{2}(J\dot{J}_{1})\wedge\psi)$
$+\int_{\Sigma}B(\psi(J\dot{J}_{2})\wedge\dot{\psi}_{1})-\int_{\Sigma}B(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2})$
$+\int_{\Sigma}B(\psi(\dot{J}_{1}\dot{J}_{2})\wedge\psi)-\int_{\Sigma}B(\psi(\dot{J}_{2}\dot{J}_{1})\wedge\psi)$
$=-2\int_{\Sigma}B(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2}))-2\int_{\Sigma}B(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2})$
$-2\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2}))$
$=4\boldsymbol{\sigma}^{\mathbb{I}}-4\pi^{*}_{2}\omega_{\mathbf{I}},$
which proves the first of the statement. As for the final part, from the previous formulae we conclude
$dd^{c}_{\mathbb{I}}\nu$ $=2i\partial_{\mathbb{I}}\bar{\partial}_{\mathbb{I}}\nu$
$=-4\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})+4d\lambda$
$=-4\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})+4\boldsymbol{\sigma}^{\mathbb{I}}-4\pi^{*}_{2}\omega_{\mathbf{I}}$
$=-4\pi^{*}_{2}\omega_{{\curly A}}+4\boldsymbol{\sigma}^{\mathbb{I}}.$
Consider the extended gauge group $\widetilde{{\curly K}}$ of $h$ and $\omega$. Similarly as in [AGG, Section 2.2], the group $\widetilde{{\curly K}}$ acts on $(\curly{X},\mathbb{I},\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon})$ preserving $\mathbb{I}$ and $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$, and covering the $\curly{H}$-action on $\curly{J}$ by push-forward. We are ready to prove the main result of this section. We use the same notation as in Proposition 3.9.
Proposition 5.10.
The action of $\widetilde{{\curly K}}$ on $(\curly{X},\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon})$ is Hamiltonian, with equivariant moment map given by
$\langle\mu_{\widetilde{{\curly K}}}^{\mathbb{I}}(J,A,\psi),\zeta\rangle$ $=\alpha\int_{\Sigma}B\left(F_{A}-\tfrac{1}{2}[\psi\wedge\psi],A\zeta\right)$
$-\int_{\Sigma}f\left(\varepsilon S_{J}\omega+\alpha d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-\alpha d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)$
Proof.
By Lemma 5.9 we have that $\boldsymbol{\sigma}^{\mathbb{I}}=\pi^{*}_{2}\omega_{\mathbf{I}}+d\lambda$, and since $\lambda$ is preserved by the $\widetilde{{\curly K}}-$action, there exists an equivariant moment map for $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$ given by
$\langle\mu_{\widetilde{{\curly K}}}^{\mathbb{I}}(J,A,\psi),\zeta\rangle=-\varepsilon\int_{\Sigma}fS_{J}\omega+\alpha\operatorname{Re}\int_{\Sigma}B(F_{D},D\zeta)-\alpha\lambda(\zeta\cdot(J,A,\psi)),$
where $D=A+i\psi$ and we have used that $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ and Proposition 4.1. The proof follows from the formulae
$\operatorname{Re}\int_{\Sigma}B(F_{D},D\zeta)$ $=\int_{\Sigma}B\left(F_{A}-\tfrac{1}{2}[\psi\wedge\psi],A\zeta\right)-B\left(d_{A}\psi,\psi(y)\right)$
$=\int_{\Sigma}B\left(F_{A}-\tfrac{1}{2}[\psi\wedge\psi],A\zeta\right)-B\left(\Lambda_{\omega}(d_{A}\psi),\psi\right)\wedge i_{y}\omega$
$=\int_{\Sigma}B\left(F_{A}-\tfrac{1}{2}[\psi\wedge\psi],A\zeta\right)+\int_{\Sigma}df\wedge\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right),$
$=\int_{\Sigma}B\left(F_{A}-\tfrac{1}{2}[\psi\wedge\psi],A\zeta\right)-\int_{\Sigma}fd\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right),$
where $\zeta$ covers the Hamiltonian vector field with Hamiltonian $f\in C^{\infty}_{0}(\Sigma)$, and (see the proof of Proposition 3.9):
$\lambda(\zeta\cdot(J,A,\psi))$ $=-\frac{1}{4}\int_{\Sigma}B(\psi(JL_{y}J)\wedge\psi)$
$=-\int_{\Sigma}B(\psi(Jy),d_{A}(J\psi))$
$=-\int_{\Sigma}B(\psi,\Lambda_{\omega}(d_{A}(J\psi)))\wedge i_{Jy}\omega$
$=\int_{\Sigma}i_{Jy}\omega\wedge B(\psi,\Lambda_{\omega}(d_{A}(J\psi)))$
$=\int_{\Sigma}d^{c}f\wedge B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi)))\right)$
$=-\int_{\Sigma}fd^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)$
Similarly as in Proposition 3.9, the $\widetilde{{\curly K}}-$action produces a coupling term in the base $\Sigma$ (cf. Remark 4.2), which combines with the scalar curvature of the metric $g=\omega(,J)$. In particular, similarly as in [AGG], zeros of the moment map $\mu_{\widetilde{{\curly K}}}^{\mathbb{I}}$ are given by solutions of the coupled system of equations
(5.15) $\begin{split}F_{A}-\tfrac{1}{2}[\psi\wedge\psi]&=0\\ \varepsilon S_{J}+\alpha\Lambda_{\omega}\left(d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
In the next section we combine these equations with the integrability condition $\bar{\partial}_{A}\varphi=0$ for the Higgs field, in order to introduce the universal moduli space of Hitchin’s equations.
5.4. Metric structure of the universal moduli of Hitchin’s equations
In this section we combine the results of the previous two sections in order to define a universal moduli spaces of solutions of Hitchin’s equations (2.15), varying over (a cover of) Teichmuller space. As in the previous sections, we consider a fixed principal $K$-bundle $E_{K}$ over a smooth compact oriented surface $\Sigma$ with fixed symplectic form $\omega$.
Definition 5.11.
We say that a triple $(J,A,\psi)\in\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ is a solution of the coupled Hitchin’s equations, if the following conditions are satisfied
(5.16) $\begin{split}F_{A}-[\varphi,\tau(\varphi)]&=0,\\ \bar{\partial}_{J,A}\varphi&=0,\\ S_{g}&=\frac{2\pi\chi(\Sigma)}{V},\end{split}$
where $g=\omega(,J)$ and $\varphi=\psi^{1,0_{J}}$.
As in Section 5.1, consider $\curly{X}^{Higgs}\subset\curly{X}$ the subset defined by the condition $\bar{\partial}_{J,A}\varphi=0$. By Proposition 5.2, $\curly{X}^{Higgs}$ is formally a complex submanifold of $(\curly{X},\mathbb{I})$ preserved by the $\widetilde{{\curly K}}-$action. For any choice of coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, $\curly{X}^{Higgs}$ inherits a minimal coupling structure $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$ and moment map $\mu_{\widetilde{{\curly K}}^{\mathbb{I}}|\curly{X}^{Higgs}}$, as in Proposition 5.10. We define the universal moduli space of solutions of Hitchin’s equations as the symplectic quotient
$\mathcal{U}^{Hit}(G)_{\alpha}^{\varepsilon}:=(\mu_{\widetilde{{\curly K}}^{\mathbb{I}}|\curly{X}^{Higgs}})^{-1}(0)/\widetilde{{\curly K}}$.
By the explicit formula for the moment in Proposition 5.10, the underlying space to $\mathcal{U}^{Hit}(G)_{\alpha}^{\varepsilon}$ is independent of the choice of parameters, as the coupling terms in the scalar equation in (5.15) vanishes identically upon imposing the integrability condition $\bar{\partial}_{J,A}\varphi=0$. We shall denote this moduli space simply by $\mathcal{U}^{Hit}(G)$. Hence, for any $\alpha>0$ and $\varepsilon\in\{-1,1\}$, the moduli space $\mathcal{U}^{Hit}(G)$ parametrizes solutions of the equations (5.16) modulo the $\widetilde{{\curly K}}-$action, and the only difference is the induced presymplectic structure on the moduli space. This structure is furthermore symplectic in the case $\varepsilon=-1$, by Lemma 5.8. By construction, there are natural maps (cf. (5.3))
$\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)\to\mathcal{T}.$
Hence, in particular, $\mathcal{U}^{Hit}(G)$ can be regarded as a fibration over the Teichmüller space $\mathcal{T}$. Similarly as in Section 4.2, since the symmetric $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is not positive definite (see Lemma 5.8), it is not obvious a priori that $\mathcal{U}^{Hit}(G)$ inherits a complex structure compatible with the (pre)symplectic structure induced by $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$. The main goal of this section is to study sufficient conditions under which this natural condition for the moduli space holds, furthermore proving that the map $\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)$ is holomorphic.
The analysis is analogue to that in Sections 4.2 and 3.3, and hence we just outline the argument. The first step is to undertake a gauge fixing for solutions of the coupled Hitchin’s equations (5.16), whereby the complex structure (5.1) and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ descend to the moduli space. We observe that this will produce a priori different complex structures on $\mathcal{U}^{Hit}(G)$.
We fix a solution $(J,A,\psi)$ of the coupled Hitchin’s equations (5.16).
Lemma 5.12.
The linearization of the coupled Hitchin’s equations (5.16) at $(J,A,\psi)$ is given by
(5.17) $\begin{split}d_{A}a-[\dot{\psi},\psi]&=0,\\ d_{A}\dot{\psi}+[a,\psi]&=0,\\ d_{A}(J\dot{\psi})+[a,J\psi]-d_{A}(\psi(\dot{J}))&=0,\\ \varepsilon\delta S(\dot{J})&=0,\end{split}$
where $\delta S\colon T_{J}\curly{J}\to C_{0}^{\infty}(\Sigma,\mathbb{R})$ is the linearization of the scalar curvature.
We denote by ${\bf L}(\dot{J},a,\dot{\psi})$ the differential operator defined by the left-hand side of equations (5.17). Using the same notation as in (4.7), we define a complex of linear differential operators
(5.18) ${(\mathcal{S}^{*}_{0})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{S}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}}$${\mathcal{S}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L}}$${\mathcal{S}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
The cohomology $H^{1}(\mathcal{S}^{*}_{0}):=\frac{\operatorname{Ker}{\bf L}}{\operatorname{Im}\,\mathbf{P}}$ can be formally identified with the tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{Hit}(G)$. Our next result shows that the moduli space $\mathcal{U}^{Hit}(G)$ is finite dimensional.
Lemma 5.13.
The sequence (5.18) is an elliptic complex of multi-degree linear differential operators. Consequently, the cohomology groups $H^{j}(\mathcal{S}^{*}_{0})$, with $j=0,1,2$, are finite-dimensional.
In order to construct a complex structure induced by (5.1) on the moduli space, we work orthogonal to the image of the infinitesimal action operator $\mathbf{P}$ in (5.18) with respect to the indefinite pairing $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$. The existence of this complex structure will automatically yield a symmetric tensor of type $(1,1)$, since the two-form $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$ is well defined on the cohomology $H^{1}(\mathcal{S}^{*}_{0})$ by Proposition 5.10. Consider the $L^{2}$-pairing on $\mathcal{S}^{0}$ define by (4.8). Consider the map $\mu\colon\curly{X}\to\mathcal{S}^{0}$ defined by (cf. Proposition 5.10)
$\mu^{\mathbb{I}}(J,A,\psi)=\left(\mu^{\mathbb{I}}_{0},\mu^{\mathbb{I}}_{1}\right).$
where
$\mu^{\mathbb{I}}_{0}$ $=-\left(\varepsilon S_{J}\omega+\alpha d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-\alpha d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)$
$\mu^{\mathbb{I}}_{0}$ $=\alpha\left(F_{A}-\tfrac{1}{2}[\psi\wedge\psi]\right).$
Lemma 5.14.
The following operator provides a formal adjoint of the infinitesimal action $\mathbf{P}$ for the pairings (4.8) and (5.11)
$\mathbf{P}^{*}_{\alpha,\varepsilon}=\delta\mu^{\mathbb{I}}\circ\mathbb{I}\colon\mathcal{S}^{1}\to\mathcal{S}^{0}.$
Consider now the differential operator
[…]
The key condition on the solution $(J,A,\psi)$ of (5.16) which we need to assume in order to construct the complex structure on the moduli space is the vanishing of the kernel of $\mathcal{L}_{0}^{\alpha,\varepsilon}$. Notice that, unlike in the standard cases in gauge theory in which the parameter space metric is positive definite, $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}$ does not relate in general to automorphisms of the triple $(J,\psi,A)$, but rather to null vectors with respect to $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$.
Proposition 5.15.
The operator $\mathcal{L}_{0}^{\alpha,\varepsilon}$ is Fredholm with zero index. Furthermore, elements $(f,u)\in\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}$ are smooth.
Assuming that $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}$ is trivial, in the next result we obtain a natural gauge fixing via a $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$-orthogonal decomposition
(5.20) $\mathcal{S}^{1}=\operatorname{Im}\;\mathbf{P}\oplus(\operatorname{Im}\;\mathbf{P})^{\perp_{\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}}}.$
Lemma 5.16.
Assume that $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}$. Then, there exists an orthogonal decomposition (5.20) for the pairing $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$. Consequently, for any element $v\in\mathcal{S}^{1}$ there exists a unique $\Pi v\in\operatorname{Im}\;\mathbf{P}$ such that $(\dot{J},a,\dot{\psi})=v-\Pi v$ solves the linear equation
$\begin{split}\mathbf{P}^{*}_{\alpha,\varepsilon}(\dot{J},a,\dot{\psi})=\delta\mu^{\mathbb{I}}\circ\mathbb{I}(\dot{J},a,\dot{\psi})=0.\end{split}$
The above lemma suggests to define the space of harmonic representatives of the complex (4.7), as follows:
$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})=\ker\bf{L}\cap\ker{\bf P}^{*}_{\alpha,\varepsilon}.$
It is important to observe that, on the subspace $\ker\bf{L}$, one has an equality
${\bf P}^{*}_{\alpha,\varepsilon}(v)={\bf L}\circ\mathbb{I}(v),\qquad\textrm{ for any }v\in\ker\bf{L}.$
Consequently, we have
$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})=\ker\bf{L}\cap\ker{\bf L}\circ\mathbb{I}.$
Our next result provides our gauge fixing mechanism for the linearization of the coupled Hitchin’s equations (5.16).
Proposition 5.17.
Assume $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}$. Then, the inclusion $\mathcal{H}^{1}(\mathcal{S}^{*}_{0})\subset\ker\bf{L}$ induces an isomorphism
$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})\simeq H^{1}(\mathcal{S}^{*}_{0}).$
More precisely, any class in the cohomology $H^{1}(\mathcal{S}^{*}_{0})$ of the complex (4.7) admits a unique representative $(\dot{J},a,\dot{\psi})$ solving the linear equations
(5.21) ${\bf L}(\dot{J},a,\dot{\psi})=0,\qquad{\bf L}\circ\mathbb{I}(\dot{J},a,\dot{\psi})=0.$
We are ready to prove our main result, which shows that the gauge fixing in Proposition 5.17 enables us to descend the complex structure $\mathbb{I}$ in $\curly{X}$ and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$, to an open subset of the moduli space $\mathcal{U}^{Hit}(G)$, via the symplectic reduction in Proposition 5.10. Define
$\mathcal{U}^{*}_{\alpha,\varepsilon}=\{[(J,A,\psi)]\;|\;\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}\}\subset\mathcal{U}^{Hit}(G).$
Theorem 5.18.
For any coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, the set $\mathcal{U}^{*}_{\alpha,\varepsilon}$ is open in $\mathcal{U}^{Hit}(G)$. For any smooth point $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$ the tangent space to $\mathcal{U}^{Hit}(G)$ at $(J,A,\psi)]$, identified with the space of solutions of the gauge fixed linear equations (5.21), inherits a complex structure $\mathbb{I}$, independent of $\alpha$ and $\varepsilon$, and a symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ such that $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}=\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(\mathbb{I},)$, given respectively by (5.1) and (5.8), and where $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$ stands for the restriction of (5.10). Furthermore,
  1. (1)
    if $\varepsilon=1$ the tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is possibly degenerate,
  2. (2)
    if $\varepsilon=-1$ the tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is non-degenerate, and defines a pseudo-Kähler structure on the moduli space.
To finish this section, we provide an explicit formula for the (pre)symplectic structure $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$. The proof is straightforward from the previous discussion and Proposition 5.6.
Corollary 5.19.
Let $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$ be a smooth point and take $v_{1},v_{2}$ tangent vectors of $\mathcal{U}^{Hit}(G)$ at $(J,A,\psi)]$, identified with solutions $(\dot{J}_{j},a_{j},\dot{\psi}_{j})$ of the gauge fixed linear equations (5.21). Then, one has
(5.22) $\begin{split}\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a_{1}\wedge a_{2})-\alpha\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{\alpha}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{\alpha}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{\alpha}{2}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right),\\ \mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a\wedge Ja)+\alpha\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{\alpha}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right),\end{split}$
5.5. Comparison with $\boldsymbol{\mathcal{U}^{Higgs}(G)}$ and existence
In this section we establish a comparison between the moduli space $\mathcal{U}^{Hit}(G)$ and the universal moduli space of $G$-Higgs bundles $\mathcal{U}^{Higgs}(G)$ constructed in Section 5.1. As we will see, Theorem 5.18 induces a natural holomorphic map
$\mathcal{U}^{Hit}(G)\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{Higgs}(G)$
and hence a holomorphic map into Teichmuller space $\mathcal{T}$. We will also prove that for genus of the surface $\Sigma$ bigger than zero, the open $\mathcal{U}^{*}\subset\mathcal{U}^{Hit}(G)$ is non-empty for sufficiently small values of $\alpha$. Our analysis is very similar to that in Section 4.3, and hence we omit the details.
Consider a point $(J,A,\psi)]\in\mathcal{U}^{Higgs}(G)$, regarded as the $\widetilde{\curly{G}}-$orbit of $(J,A,\psi)\in\curly{X}$ solving the equations
(5.23) $\begin{split}\bar{\partial}_{J,A}\psi^{1,0_{J}}=0\end{split}$
The tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{Higgs}(G)$ can be formally identified with the cohomology of the complex of linear differential operators
(5.24) ${(\mathcal{C}_{0}^{*})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{C}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}^{c}_{0}}$${\mathcal{C}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L^{c}_{0}}}$${\mathcal{C}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
where we have
$\mathcal{C}^{0}=\operatorname{Lie}\widetilde{\curly{G}}\cong\Omega^{0}(T\Sigma)\oplus\Omega^{0}(X,E_{G}(\mathfrak{g})),\qquad\mathcal{C}^{1}=\mathcal{S}^{1},\qquad\mathcal{C}^{2}=\Omega^{2}(X,E_{G}(\mathfrak{g}))$
and
$\mathbf{P}^{c}_{0}(y,u_{0}+iu_{1})$ $=-\left(L_{y}J,d_{A}u_{0}+Jd_{A}u_{1}+i_{y}F_{A},[u_{0},\psi]+[u_{1},J\psi]+i_{y}d_{A}\psi\right),$
$\mathbf{L}^{c}_{0}(\dot{J},a,\dot{\psi})$ $=d_{A}\dot{\psi}^{1,0_{J}}+[a,\psi^{1.0_{J}}]-\frac{i}{2}d_{A}(\psi(\dot{J})).$
Lemma 5.20.
The sequence (5.24) is an elliptic complex of degree-one linear differential operators. Consequently, the cohomology groups $H^{j}(\mathcal{C}^{*})$, with $j=0,1,2$, are finite-dimensional.
Applying Theorem 2.7, a solution $(J,A,\psi)$ of the coupled Hitchin’s equations (5.16) induces a polystable $G$-Higgs bundle $(E,\varphi)\cong(\bar{\partial}_{J,A},\psi^{1,0_{J}})$. This fact, jointly with the natural inclusion $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$, leads to a continuous map
(5.25) $\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)\colon[(J,A,\psi)]\mapsto[(J,A,\psi)].$
By definition of $\mathcal{U}^{Hit}(G)$, there is a continuous diagram of moduli spaces
(5.26) ${\mathcal{U}^{Hit}(G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{U}^{Higgs}(G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mu_{\curly{H}}^{-1}(0)/\curly{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}}$
where $\mu_{\curly{H}}^{-1}(0)/\curly{H}$ is the moduli space of constant scalar curvature Kähler metrics with total volume $V$, modulo $\omega$-Hamiltonian diffeomorphisms. The fibre of $\mu_{\curly{H}}^{-1}(0)/\curly{H}\to\mathcal{T}$ can be identified with $H^{1}(\Sigma,\mathbb{R})$ [F]. Building on Theorem 5.18, our next goal is to prove that this induces a holomorphic map $\mathcal{U}^{Hit}(G)\supset\mathcal{U}^{*}\to\mathcal{U}^{Higgs}(G)$.
Lemma 5.21.
Let $(J,A,\psi)]\in\mathcal{U}^{*}_{0}$. Then, (5.25) induces a complex linear map
(5.27) ${H^{1}(\mathcal{S}^{*}_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathcal{C}^{*}_{0}).}$
where the complex structure on $H^{1}(\mathcal{S}^{*}_{0})$ is the one induced by Proposition 5.17.
To finish this section, we address the question of non-emptyness of the moduli space $\mathcal{U}^{Hit}(G)$, in genus $g(\Sigma)\geqslant 2$. For this, we fix a compact Riemann surface $X=(\Sigma,J)$ and consider a $G$-Higgs bundle $(E,\varphi)$. In this setup, consider the coupled equations
(5.28) $\begin{split}F_{h}-[\varphi\wedge\tau_{h}\varphi]&=0\\ \varepsilon S_{g}&=\varepsilon\frac{2\pi\chi(\Sigma)}{V}\end{split}$
for pairs $(g,h)$, where $g$ is a Kähler metric on $X$ with total volume $V$ and $h\in\Omega^{0}(E_{G}(G/K))$ is a reduction of structure group of $E$ to $K$. A solution $(g,h)$ of the equations (5.28) determines then a solution of (5.16), given by the triple $(J,A_{h},-i(\varphi-\tau_{h}\varphi))$. We are ready to present our last main result.
Theorem 5.22.
Let $(E,\varphi)$ be stable $G$-Higgs bundle over a compact Riemann surface $X=(\Sigma,J)$ with genus $g(\Sigma)\geqslant 2$. Then, for any fixed total volume $V>0$ and parameter $\varepsilon\in\{-1,1\}$, there exists $\alpha_{0}>0$ such that for any $0<\alpha<\alpha_{0}$ there exists a solution $(g_{\alpha},h_{\alpha})$ of the equations (5.16) with
$(J,A_{h_{\alpha}},\psi_{h_{\alpha}})]\in\mathcal{U}^{*}_{\alpha,\varepsilon}\subset\mathcal{U}^{Hit}(G).$
Furthermore, the induced maps $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{Higgs}(G)$ and
(5.29) $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mu_{\curly{H}}^{-1}(0)/\curly{H}$
are holomorphic, where $\mu_{\curly{H}}^{-1}(0)/\curly{H}$ is the moduli space of constant scalar curvature Kähler metrics with total volume $V$.
The restriction of the $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ in Corollary 5.19 to the fibres of (5.29) is $\alpha\mathbf{g}$, where $\mathbf{g}$ denotes the hyperkähler metric on the moduli space of solutions of Hitchin’s equations. Consequently, (5.29) has a natural structure of Kähler fibration with coupling form $\boldsymbol{\omega}_{1,0}$ (see (5.22)) and Kähler Ehresmann connection.
Proof.
The proof is analogue of that of Theorem 4.14, and is therefore omitted. The only difference is that we can control the non-degeneracy of $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ along the fibres over the moduli space of constant scalar curvature Kähler metrics $\mathcal{N}=\mu_{\curly{H}}^{-1}(0)/\curly{H}$. To see this, we observe that if $(\dot{J},a,\dot{\psi})$ is in the tangent to the fibre over $\mathcal{N}$, then due to the gauge fixing in Proposition 5.17, we must have $\dot{J}=0$, and the statement follows from Corollary 5.19. Hence, assuming that $\dot{J}=L_{J\eta_{f}}J$, and since the kernel of the Lichnerowicz operator is given by the Hamiltonian functions of Killing holomorphic vector field, $f$ must be constant by our assumption $g(\Sigma)\geqslant 2$. ∎
Remark 5.23.
Even though the proof of Theorem 5.18 works in the case $g(\Sigma)=1$, the hypothesis of existence of a stable $G$-Higgs bundle is never satisfied in this case [FGN]. We thank Emilio Franco for this observation.

### v2

@@
 The moduli space $\mathcal{R}$ of representations of the fundamental group of a closed oriented surface $\Sigma$ has an interesting interplay with classical Teichmüller theory.
-For the case of a compact Lie group $K$, a choice of complex structure $J$ on $\Sigma$ determines an homeomorphism of the moduli space $\mathcal{R}(K)$ of $K$-representations with a moduli space of holomorphic principal $G$-bundles over the compact Riemann surface $X_{J}=(\Sigma,J)$, via the Narasimhan–Seshadri theorem [NS] and its extension to principal bundles by Ramanathan [R1, R2].
+For the case of a compact Lie group $K$, a choice of complex structure $J$ on $\Sigma$ determines a homeomorphism of the moduli space $\mathcal{R}(K)$ of $K$-representations with a moduli space of holomorphic principal $G$-bundles over the compact Riemann surface $X_{J}=(\Sigma,J)$, via the Narasimhan–Seshadri theorem [NS] and its extension to principal bundles by Ramanathan [Ra1, Ra2].
 Here, $G$ denotes the complex reductive Lie group given by the complexification of $K$.
@@
 A different aspect of this interplay arose forty years ago in Nigel Hitchin’s study of the self-duality equations on a Riemann surface $X_{J}$ [Hi1]
-(1.1) $\begin{array}[]{l}F_{A}-[\varphi,\tau(\varphi)]=0,\\ \bar{\partial}_{J,A}\varphi=0,\end{array}$
-and subsequent work by Simpson [Si3, Si4].
+(1.1) $\begin{split}F_{A}-[\varphi,\tau(\varphi)]&=0,\\ \bar{\partial}_{J,A}\varphi&=0,\end{split}$
+and subsequent work by Hitchin [Hi2, Hi3] and Simpson [Si1, Si2, Si3, Si4].
 Here $A$ is a connection on a principal $K$-bundle over $\Sigma$ and $\varphi\in\Omega^{1,0}(X_{J},E_{K}(\mathfrak{g}))$ is the Higgs field.
-The corresponding gauge-theoretical moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$ has three different incarnations, as a character variety $\mathcal{R}_{c}(G)$ for a complex reductive Lie group $G$, as a moduli space space $\mathcal{M}^{\operatorname{Flat}}(G)$ of reductive flat $G$-connections, and as a moduli space of polystable $G$-Higgs bundles $\mathcal{M}(G)$.
+The corresponding gauge-theoretical moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$ has three different incarnations, as a character variety $\mathcal{R}_{c}(G)$ for the complex reductive Lie group $G$, as a moduli space $\mathcal{M}^{\operatorname{Flat}}(G)$ of reductive flat $G$-connections, and as a moduli space of polystable $G$-Higgs bundles $\mathcal{M}^{Higgs}(G)$.
 These three viewpoints on the moduli space of Hitchin’s equations determine a hyperkähler structure on $\mathcal{M}^{\operatorname{Hit}}(G)$.
-Interestingly, only the moduli space $\mathcal{M}(G)$ depends on the choice of a complex structure on $\Sigma$.
-Furthermore, the twistor space $Z\to\mathbb{CP}^{1}$ for the hyperkähler structure has generic complex structure $\mathcal{M}^{\operatorname{Flat}}(G)$, while the special points over $\mathbb{CP}^{1}$ correspond to $\mathcal{M}(G)$ and its conjugate.
+Interestingly, only the moduli space $\mathcal{M}^{Higgs}(G)$ depends on the choice of a complex structure on $\Sigma$.
+Furthermore, the twistor space $Z\to\mathbb{CP}^{1}$ for the hyperkähler structure has generic complex structure $\mathcal{M}^{\operatorname{Flat}}(G)$, while the special points over $\mathbb{CP}^{1}$ correspond to $\mathcal{M}^{Higgs}(G)$ and its conjugate.
 In this paper we investigate metric aspects of the universal moduli space of solutions of Hitchin’s equations varying over the Teichmüller space $\mathcal{T}$.
-Our approach is gauge-theoretical in nature, and builds on the classical theory of Kähler fibrations [GLS, M] and in the symplectic interpretation of constant scalar curvature Kähler metrics [Do2, F].
+Our approach is gauge-theoretical in nature, and builds on the classical theory of Kähler fibrations [GLS, Mu] and on the symplectic interpretation of constant scalar curvature Kähler metrics [Do2, Fu].
 We fix a symplectic structure $\omega$ on $\Sigma$ and consider the coupled system of equations
@@
 where $J$ is a complex structure on $\Sigma$ and $S_{g}$ is the scalar curvature of the metric $g=\omega(,J)$.
-We denote by $\mathcal{U}^{Hit}(G)$ the moduli space of solutions of (1.2) modulo unitary gauge (see Section 5.4).
+Let $\mathcal{U}^{Hit}(G)$ be the moduli space of solutions of (1.2) modulo unitary gauge (see Section 5.4), and $\mathcal{U}^{Higgs}(G)$ be the universal moduli space of $G$-Higgs bundles (see Section 5.1).
 Our first main result can be stated as follows (cf.
-Theorem 5.18 and Theorem 5.22):
+Theorem 5.18, and Theorem 5.22).
 Theorem 1.1.
@@
 Furthermore, the induced maps $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{Higgs}(G)$ and
-(1.3) $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mu_{\curly{H}}^{-1}(0)/\curly{H}$
+(1.3) $\mathcal{U}^{*}_{\alpha,\varepsilon}\longrightarrow\mu_{\curly{H}}^{-1}(0)/\curly{H}$
 are holomorphic, where $\mu_{\curly{H}}^{-1}(0)/\curly{H}$ is the moduli space of constant scalar curvature Kähler metrics on $\Sigma$ with total volume $V$.
@@
 This follows from an explicit formula (see Corollary 5.19)
-(1.5) $\begin{split}\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a\wedge Ja)+\alpha\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{\alpha}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right),\end{split}$
+(1.5) $\begin{split}\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}\dot{J})\omega\\ &+\alpha\int_{\Sigma}B(a\wedge Ja)+\alpha\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\right)\\ &-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right),\end{split}$
 where $B$ is a positive-definite invariant metric on $\mathfrak{k}=\operatorname{Lie}K$.
 Hence, it is not obvious a priori that $\mathcal{U}^{Hit}(G)$ inherits a complex structure compatible with $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$.
-The key step is to undertake a gauge fixing for solutions of the coupled Hitchin’s equations (1.2), whereby the complex structure $\mathbb{I}$ and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ descend to the moduli space.
+The key step is to undertake a gauge fixing for solutions of the coupled Hitchin equations (1.2), whereby the complex structure $\mathbb{I}$ and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ descend to the moduli space.
 Difficulties will arise, due to the fact that $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is neither a definite pairing nor non-degenerate.
@@
 endowed with a complex structure $\mathbb{J}$ and a pre-symplectic structure $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$. Furthermore, the induced map
-(1.7) $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)$
+(1.7) $\mathcal{U}^{*}_{\alpha,\varepsilon}\longrightarrow\mathcal{U}^{\operatorname{Flat}}(G)$
 is holomorphic, and
@@
 $\Phi=\varepsilon\nu{J}+\frac{\alpha}{2}\|\psi\|^{2}_{L^{2}}$
-and $\nu{J}$ is induced by the global Kähler potential in the space of complex structures compatible with the orientation $\curly{J}$ (see [F, Section 4]).
-Similarly as for the universal moduli space of Hitchin’s equations (1.2), the (possibly degenerate) moduli space Kähler form and metric admit explicit formulae upon suitable gauge fixing
-(1.8) $\begin{split}\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2}))\\ &+\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge J(\psi(\dot{J}_{2}))).\end{split}$
+and $\nu{J}$ is induced by the global Kähler potential in the space of complex structures compatible with the orientation $\curly{J}$ (see [Fu, Section 4]).
+Similarly to the universal moduli space of Hitchin’s equations (1.2), the (possibly degenerate) moduli space Kähler form and metric admit explicit formulae upon suitable gauge fixing
+(1.8) $\begin{split}\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &+\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge J(\psi(\dot{J}_{2}))).\end{split}$
 From the previous expression, we observe that, at least formally, the restriction of $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ to the fibres over $\mathcal{T}$ coincides up to scaling with Hitchin’s symplectic structure $\omega_{\mathbf{J}}$.
-A complete proof of this requires a better understanding of the particular gauge fixing mechanism which we use to construct the moduli space complex structure $\mathbb{J}$ (see Proposition 4.9), and we leave it as an open question.
+A complete proof of this requires a deeper understanding of the particular gauge fixing mechanism which we use to construct the moduli space complex structure $\mathbb{J}$ (see Proposition 4.9), and we leave it as an open question.
 Based on the twistor space structure for Hitchin’s original hyperkähler moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$, it is natural to speculate on a relation between the moduli spaces $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ and $\mathcal{U}^{Higgs}(G)$ in the weak coupling limit
@@
 $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}\to\mathbf{g}^{\mathbb{I}}_{\tilde{\alpha},\varepsilon}.$
-We are far from even grasping the solution of this proposal, and we leave it as an open question for future studies.
-Motivation for the present work comes from the programme initiated by the first three authors in [AGG] more than 15 years ago based on the 2009 PhD Thesis of the second author [GF].
+We are far from even grasping the solution to this problem, and we leave it as an open question for future studies.
+Motivation for the present work comes from the programme initiated by the first three authors in [AGG1] more than 15 years ago based on the 2009 PhD Thesis of the second author [Ga].
 In that paper, we consider a pair $(X,E)$ consisting of a holomorphic principal $G$-bundle $E$ over a compact complex manifold $X$, where $G$ is a complex reductive group, and study certain coupled equations for a Kähler metric on $X$ and a reduction of structure group of $E$ to a maximal compact subgroup $K\subset G$.
@@
 In early joint work with the fourth author, we extended our symplectic point of view to the problem of constructing universal moduli spaces for Hitchin’s equations and Higgs bundles.
-The first steps of this attempt appeared in Chapter 7 of the fourth author’s 2018 ETH PhD Thesis [T], and is the seed of the present work.
-In view of the current interest in the field, after a dormant period, we decided to resume this investigation.
-We should warn that the use of the term ‘universal’ that we make here does not necessarily imply some expected functoriality properties that such term some times entails.
-We have sticked, however, to this term since that is the term that we originally used in our initial work [T].
+The first steps of this attempt appeared in Chapter 7 of the fourth author’s 2018 ETH PhD Thesis [Tr], and is the seed of the present work.
+In view of the current interest in the subject, after a dormant period, we decided to resume this investigation.
+We should warn that the use of the term ‘universal’ that we make here does not necessarily imply any expected functoriality properties that such term some times entails.
+We have sticked, however, to this term since that is the one that we originally used in our initial work [Tr].
 After the completion of this work, we received a copy of a manuscript [CTW], where related topics are treated.
-It should be very interesting to compare our approach to other analytic constructions under study [CTW, Hi4], as well as the more algebraic point of view taken by Simpson [Si2, Si3] and others [BBN, BDP, D, DF].
+It should be very interesting to compare our approach to other analytic constructions under study [CTW, Hi4], as well as the more algebraic point of view taken by Simpson [Si2, Si3] and others [BBN, BDP, Da, DF].
 Acknowledgements.
 The authors wish to thank Emilio Franco, Nigel Hitchin and Ignasi Mundet for helpful conversations.
-Special thanks are due to Richard Wentworth for stimulating discussions – in particular for spotting a mistake in the signature of the metric $\mathbf{g}^{\mathbb{I}}_{\alpha,1}$ in an earlier version of this manuscript – as well as for his lectures at ICMAT in June 2025, which encouraged us to resume this project.
+Special thanks are due to Richard Wentworth for stimulating discussions — in particular for spotting a mistake in the signature of the metric $\mathbf{g}^{\mathbb{I}}_{\alpha,1}$ in an earlier version of this manuscript — as well as for his lectures at ICMAT in June 2025, which encouraged us to resume this project.
 2. Flat connections and Higgs bundles
@@
 The space $\curly D$} is furthermore holomorphic symplectic, with holomorphic symplectic structure, defined by
-(2.2) $\Omega_{\mathbf{J}}(\dot{D}_{1},\dot{D}_{2})=\int_{X}B(\dot{D}_{1}\wedge\dot{D}_{2}),$
+(2.2) $\Omega_{\mathbf{J}}(\dot{D}_{1},\dot{D}_{2})=\int_{\Sigma}B(\dot{D}_{1}\wedge\dot{D}_{2}),$
 which is preserved by the action of the gauge group $\curly{G}$ of $E_{G}$.
 We have that $\operatorname{Lie}\curly{G}=\Omega^{0}(\Sigma,E_{G}(\mathfrak{g}))$ and hence its dual $(\operatorname{Lie}\curly{G})^{\ast}$ can be identified with $\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))$.
-Similarly as for the case of unitary connections, studied by Atiyah–Bott [AB], there existes an equivariant complex moment map for the action of $\curly{G}$ on $\curly D$} given by
-(2.3) $\mu{G}\colon{\curly D}$ $\longrightarrow$ $\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))$
-(2.4) $D$ $\longmapsto$ $F_{D},$
-where $F_{D}$ is the curvature of $D$. The zero moment map condition, that is when $F_{D}=0$, corresponds to the flatness of $D$. In this case one refers to the pair $(E_{G},D)$ as a flat $G$-bundle.
+As in the case of unitary connections, studied by Atiyah–Bott [AB], there exists an equivariant complex moment map for the action of $\curly{G}$ on $\curly D$} given by
+(2.3) $\begin{array}[]{rcl}\mu{G}\colon{\curly D}&\longrightarrow&\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))\\ D&\longmapsto&F_{D},\end{array}$
+where $F_{D}$ is the curvature of $D$. The zero moment map condition, that is, $F_{D}=0$, corresponds to the flatness of $D$. In this case one refers to the pair $(E_{G},D)$ as a flat $G$-bundle.
 In order to construct a Hausdorff moduli space of flat $G$-connections modulo gauge, one needs to impose a natural stability condition arising from Geometric Invariant Theory.
-We say that a bundle with $G$-connection $(E_{G},D)$ is reductive (also called completely reducible) if it splits as a product of irreducible sub-bundles, where irreducible (also called simple or stable in the literature) means that there exists no nontrivial $D$-invariant sub-bundle.
-For a flat $G$-bundle $(E_{G},D)$, this condition can be expressed in terms of the holonomy representation: the holonomy of $D$ defines a representation
-$\rho:\pi_{1}(\Sigma)\to G,$
-of the fundamental group $\pi_{1}(\Sigma)$ of $\Sigma$.
-A representation $\rho:\pi_{1}(X)\to G$ is said to be reductive if the Zariski closure of the image of $\rho$ is a reductive group, or equivalently if $\operatorname{ad}_{G}\circ\rho$ is completely reducible, where $\operatorname{ad}_{G}:G\to\mathrm{GL}(\mathfrak{g})$ is the adjoint representation.
-Then, one can prove that the flat connection $D$ on $E_{G}$ is reductive if and only if its corresponding representation is reductive.
+If the connection $D$ is flat, the holonomy map defines a representation $\rho$ of the fundamental group $\pi_{1}(\Sigma)$ of $\Sigma$ in $G$.
+The representation $\rho:\pi_{1}(\Sigma)\to G$ is said to be reductive if the Zariski closure of the image of $\rho$ is a reductive group, or equivalently if $\operatorname{ad}_{G}\circ\rho$ is completely reducible, where $\operatorname{ad}_{G}:G\to\mathrm{GL}(\mathfrak{g})$ is the adjoint representation.
+We say that $(E_{G},D)$ is reductive if $\rho$ is reductive.
 The existence of the moment map $\mu{G}$ implies that the flatness condition is invariant under the action of $\curly{G}$ on $\curly D$}, and so is the reductiveness condition, and we can thus consider the moduli space
@@
 Theorem 2.1.
-The moduli space of flat $G$-connections $\mathcal{M}^{\operatorname{Flat}}(G)$ on $E_{G}$ admits a natural structure of singular complex manifold, and a holomorphic symplectic structure induced by $\Omega_{\mathbf{J}}$ on its smooth locus.
+The moduli space of flat $G$-connections $\mathcal{M}^{\operatorname{Flat}}(G)$ on $E_{G}$ admits a natural structure of a (possibly singular) complex manifold, and a holomorphic symplectic structure induced by $\Omega_{\mathbf{J}}$ on its smooth locus.
 2.2. Surface group representations
@@
 Since
-$\pi_{1}(\Sigma)=\langle a_{1},b_{1},\dotsc,a_{g},b_{g}\;\;:\;\;\prod_{i=1}^{g}[a_{i},b_{i}]=1\rangle$
+$\pi_{1}(\Sigma)=\big\langle a_{1},b_{1},\dotsc,a_{g},b_{g}\;\;:\;\;\prod_{i=1}^{g}[a_{i},b_{i}]=1\big\rangle,$
 $\operatorname{Hom}(\pi_{1}(\Sigma),G)$ can be naturally identified with the subset of $G^{2g}$ consisting of $2g$-tuples $(A_{1},B_{1}\dotsc,A_{g},B_{g})$ satisfying the algebraic equation $\prod_{i=1}^{g}[A_{i},B_{i}]=1$, and hence has a natural structure as a complex algebraic variety.
@@
 Given a representation $\rho\colon\pi_{1}(\Sigma)\to G$, consider the associated flat $G$-bundle on $\Sigma$, defined as $E_{\rho}=\tilde{\Sigma}\times_{\rho}G$, where $\tilde{\Sigma}\to\Sigma$ is the universal cover and $\pi_{1}(\Sigma)$ acts on $G$ via $\rho$.
-This gives in fact an identification between the set of equivalence classes of representations $\operatorname{Hom}(\pi_{1}(\Sigma),G)/G$ and the set of equivalence classes of flat $G$-bundles, which in turn is parametrized by the cohomology set $H^{1}(X,G)$.
+This gives in fact an identification between the set of equivalence classes of representations $\operatorname{Hom}(\pi_{1}(\Sigma),G)/G$ and the set of equivalence classes of flat $G$-bundles, which in turn is parametrized by the cohomology set $H^{1}(\Sigma,G)$.
 We can then assign a topological invariant to a representation $\rho$ given by the characteristic class $c(\rho):=c(E_{\rho})\in\pi_{1}(G)$ corresponding to $E_{\rho}$.
@@
 For a fixed $c\in\pi_{1}(G)$, the moduli space of reductive representations $\mathcal{R}_{c}(G)$ with topological invariant $c$ is defined as the subvariety
-(2.6) $\mathcal{R}_{c}(G):=\{\rho\in\mathcal{R}(G)\;\;:\;\;c(\rho)=c\}.$
+(2.5) $\mathcal{R}_{c}(G):=\{\rho\in\mathcal{R}(G)\;\mid\;c(\rho)=c\}.$
 Proposition 2.2.
-Let $E_{G}$ be a smooth principal $G$-bundle over $X$ with topological class $c$. Then there is a complex analytic isomorphism
+Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$ with topological class $c$. Then there is a complex analytic isomorphism
 $\mathcal{M}^{\operatorname{Flat}}(G)\cong{\mathcal{R}}_{c}(G).$
 Remark 2.3.
-We warn the reader that, even though there is a complex analytic isomorphism between the moduli space of flat $G$-connections $\mathcal{M}^{\operatorname{Flat}}(G)$ and the $G$-character variety with fixed topological invariant $c$, these spaces carry very different algebraic structures (see e.g.
+We warn the reader that, even though there is a real analytic isomorphism between the moduli space of flat $G$-connections $\mathcal{M}^{\operatorname{Flat}}(G)$ and the $G$-character variety with fixed topological invariant $c$, these spaces carry very different complex structures (see e.g.
 [Si3]).
@@
 $D=A+i\psi$
-where $A$ is a connection on $E_{K}$ and $\psi\in\Omega^{1}(X,E_{K}(\mathfrak{k}))$. This simple fact induces a one-to-one correspondence
-(2.7) ${\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ $\longrightarrow$ ${\curly D}$
-$(A,\psi)$ $\longmapsto$ $D:=A+i\psi$
+where $A$ is a connection on $E_{K}$ and $\psi\in\Omega^{1}(X,E_{K}(\mathfrak{k}))$. We thus have a one-to-one correspondence
+(2.6) $\begin{array}[]{rcl}{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))&\longrightarrow&{\curly D}\\ (A,\psi)&\longmapsto&D:=A+i\psi,\end{array}$
 where $\curly A$} denotes the space of connections on $E_{K}$.
@@
 where Ja=a(J)Ja=-a(J). Let us denote $\mathbf{K}:=\mathbf{I}\mathbf{J}$. Then, one can check that $\mathbf{K}$ defines a new complex structure and furthermore that the three complex structures
-(2.8) $\mathbf{I}(a,\dot{\psi})$ $=(Ja,-J\dot{\psi})$
-(2.9) $\mathbf{J}(a,\dot{\psi})$ $=(-\dot{\psi},a)$
-(2.10) $\mathbf{K}(a,\dot{\psi})$ $=(-J\dot{\psi},-Ja),$
+(2.7) $\begin{split}\mathbf{I}(a,\dot{\psi})&=(Ja,-J\dot{\psi}),\\ \mathbf{J}(a,\dot{\psi})&=(-\dot{\psi},a),\\ \mathbf{K}(a,\dot{\psi})&=(-J\dot{\psi},-Ja),\end{split}$
 satisfy the quaternion relations, and define a flat hyperkähler structure on ${\curly D}\cong T^{*}{\curly A}$ with metric tensor
@@
 (2.10) $\omega_{\mathbf{J}}((a_{1},\dot{\psi}_{1}),(a_{2},\dot{\psi}_{2}))=\int_{\Sigma}B(a_{1}\wedge J\dot{\psi}_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge Ja_{2}),$
-was studied by Corlette [C].
+was studied by Corlette [Co].
 Proposition 2.4.
@@
 whose zero locus corresponds to solutions of the harmonicity equations
-(2.14) $\begin{array}[]{l}F_{A}-\frac{1}{2}[\psi,\psi]=0,\\ d_{A}\psi=0,\\ d_{A}^{\ast}\psi=0.\end{array}$
+(2.11) $\begin{split}F_{A}-\frac{1}{2}[\psi,\psi]&=0,\\ d_{A}\psi&=0,\\ d_{A}^{\ast}\psi&=0.\end{split}$
 Note that the first two are equivalent to the flatness condition for $D$, via the formula
@@
 while the last equation is equivalent to the vanishing of the moment map $\mu_{{\curly K}}^{\mathbf{J}}$, due to the standard identity $\ast\psi=J\psi$.
-The harmonicity condition for the reduction $h$ admits an interpretation in terms of the the holonomy representation $\rho:\pi_{1}(X)\to G$ corresponding to $D$.
-The reduction $h$ is equivalent to a $\pi_{1}(X)$-equivariant smooth map
-$\tilde{h}:\tilde{X}\longrightarrow G/K,$
-where $\tilde{X}$ is the universal covering of $X$.
-Here $\pi_{1}(X)$ acts on $\tilde{X}$ by Deck transformations, and via the representation $\rho$ on $G/K$.
+The harmonicity condition for the reduction $h$ admits an interpretation in terms of the the holonomy representation $\rho:\pi_{1}(\Sigma)\to G$ corresponding to $D$.
+The reduction $h$ is equivalent to a $\pi_{1}(\Sigma)$-equivariant smooth map
+$\tilde{h}:\tilde{\Sigma}\longrightarrow G/K,$
+where $\tilde{\Sigma}$ is the universal covering of $\Sigma$.
+Here $\pi_{1}(\Sigma)$ acts on $\tilde{\Sigma}$ by Deck transformations, and via the representation $\rho$ on $G/K$.
 The symmetric space $G/K$ is equipped with a canonical $G$-invariant Riemannian metric determined by the Killing form $B$, and (2.11) is equivalent to the harmonicity of $\tilde{h}$ in the usual sense.
@@
 given by the infinite-dimensional hyperkähler reduction ${{\boldsymbol{\mu}}}_{\curly K}^{-1}(0)/{\curly K}$.
-The following theorem, due to Donaldson [Do] for $G=\mathrm{SL}(2,\mathbb{C})$, and Corlette [C] in general, is one of the main structural results of the theory.
+The following theorem, due to Donaldson [Do1] for $G=\mathrm{SL}(2,\mathbb{C})$, and Corlette [Co] in general, is one of the main structural results of the theory.
 Theorem 2.5.
-The map (2.7) induces an homeomorphism
+The map (2.6) induces a homeomorphism
 $\mathcal{M}^{\operatorname{Harm}}(G)\cong\mathcal{M}^{\operatorname{Flat}}(G).$
@@
 The non-generic complex structures on $\mathcal{M}^{\operatorname{Harm}}(G)$ correspond to the moduli space of $G$-Higgs bundles over $X$, induced by complex structure $\mathbf{I}$, and its conjugate, as we discuss in the next sections.
-2.4. $\boldsymbol{G}$-Higgs bundles and Hitchin equations
+2.4. $\boldsymbol{G}$-Higgs bundles and Hitchin’s equations
 Through this section we consider a connected semisimple complex Lie group $G$, with Lie algebra $\mathfrak{g}$, and fix an antiholomorphic involution $\tau$ of $G$ defining a maximal compact subgroup $K:=G^{\tau}\subset G$, with Lie algebra $\mathfrak{k}$.
@@
 A $G$-Higgs bundle over $X$ is a pair $(E,\varphi)$ consisting of a holomorphic principal $G$-bundle $E$ over $X$ and an element $\varphi\in H^{0}(X,E(\mathfrak{g})\otimes K_{X})$, that is a holomorphic section of $E(\mathfrak{g})\otimes K_{X}$, where $E(\mathfrak{g})$ is the adjoint bundle associated to $E$ via the adjoint representation of $G$ in $\mathfrak{g}$.
-There are appropriate notions of stability, semistability and polystability and we can consider $\mathcal{M}(G)$ to be the moduli space of isomorphism classes polystable $G$-Higgs bundles over $X$.
+There are appropriate notions of stability, semistability and polystability and we can consider $\mathcal{M}^{Higgs}(G)$ to be the moduli space of isomorphism classes polystable $G$-Higgs bundles over $X$.
 Let $h\in\Omega^{0}E(X,G/K)$ be a smooth reduction of structure group of $E$ to $K$ and let $E_{h}$ be the corresponding smooth principal $K$-bundle.
-Let $F_{h}\in\Omega^{2}(X,E_{h}(\mathfrak{k}))$ be the curvature of the unique connection compatible with $h$ and the holomorphic structure on $E$ (defined by the Chern–Singer correspondence).
+Let $F_{h}\in\Omega^{2}(X,E_{h}(\mathfrak{k}))$ be the curvature of the unique connection compatible with $h$ and the holomorphic structure on $E$, defined by the Chern–Singer correspondence [Ch, Sin].
 Here $E_{h}(\mathfrak{k})$ is the adjoint bundle of $E_{h}$.
 Abusing notation, let
-$\tau\colon\Omega^{1,0}(X,E(\mathfrak{g}))\to\Omega^{0,1}(X,E(\mathfrak{g}))$
+$\tau\colon\Omega^{1,0}(X,E(\mathfrak{g}))\longrightarrow\Omega^{0,1}(X,E(\mathfrak{g}))$
 be the conjugation defined by $h$ combined with the conjugation on complex $1$-forms on $X$.
@@
 Theorem 2.6.
-A reduction $h$ of structure group of $E$ to $K$ satisfies the Hitchin equation
+A reduction $h$ of structure group of $E$ to $K$ satisfies Hitchin’s equation
 $F_{h}-[\varphi,\tau(\varphi)]=0$
@@
 From the point of view of moduli spaces it is convenient to fix a $C^{\infty}$ principal $K$-bundle $E_{K}$ and study the moduli space of solutions to Hitchin’s equations for a pair $(A,\varphi)$ consisting of a connection $A$ on $E_{K}$ and $\varphi\in\Omega^{1,0}(X,E_{K}(\mathfrak{g}))$:
-(2.15) $\begin{array}[]{l}F_{A}-[\varphi,\tau(\varphi)]=0\\ \bar{\partial}_{A}\varphi=0.\end{array}$
+(2.12) $\begin{split}F_{A}-[\varphi,\tau(\varphi)]&=0,\\ \bar{\partial}_{A}\varphi&=0.\end{split}$
 Here $\bar{\partial}_{A}$ is the $(0,1)$ part of the covariant derivative $d_{A}$ defined by $A$.
 This is of course the Dolbeault operator defined by the holomorphic structure $J_{A}$ on $E_{G}$ corresponding to $A$ via de Chern–Singer correspondence.
-Here $E_{G}$ is the smooth principal $G$-bundle obtained frm $E_{K}$ by extension of structure group.
-The gauge group $\curly K$} of automorphisms of $E_{K}$ acts on the space of solutions and the moduli space of solutions is
+The bundle $E_{G}$ is the smooth principal $G$-bundle obtained from $E_{K}$ by extension of structure group.
+The gauge group $\curly K$} of automorphisms of $E_{K}$ acts on the space of solutions, and the moduli space of solutions is
 […]
@@
 There is a homeomorphism
-$\mathcal{M}(G)\cong\mathcal{M}^{\operatorname{Hit}}(G)$
+$\mathcal{M}^{Higgs}(G)\cong\mathcal{M}^{\operatorname{Hit}}(G).$
 To explain this correspondence we interpret the moduli space of $G$-Higgs bundles in terms of pairs $(J_{E},\varphi)$ consisting of a holomorphic structure $J_{E}$ on the smooth $G$-bundle $E_{G}$ obtained from $E_{K}$ by the extension of structure group, and $\varphi\in\Omega^{1,0}(X,E_{G}(\mathfrak{g}))$ satisfying $\bar{\partial}_{E}\varphi=0$.
-Such pairs are in correspondence with $G$-Higgs bundles $(E,\varphi)$, where $E$ is the holomorphic $G$-bundle defined by $J_{E}$ on $E_{G}$, and $\bar{\partial}_{E}\varphi=0$, that is $\varphi\in H^{0}(X,E(\mathfrak{g})\otimes K_{X})$.
-The moduli space of polystable $G$-Higgs bundles $\mathcal{M}(G)$ can now be identified with the orbit space
+Such pairs are in correspondence with $G$-Higgs bundles $(E,\varphi)$, where $E$ is the holomorphic $G$-bundle defined by $J_{E}$ on $E_{G}$, and $\bar{\partial}_{E}\varphi=0$, that is, $\varphi\in H^{0}(X,E(\mathfrak{g})\otimes K_{X})$.
+The moduli space of polystable $G$-Higgs bundles $\mathcal{M}^{Higgs}(G)$ can now be identified with the orbit space
 $\{(J_{E},\varphi)\;\;\mbox{with}\;\;\bar{\partial}_{E}\varphi=0\;\;\mbox{such that $(E,\varphi)$ is polystable}\}/\curly{G},$
@@
 Since, by the Chern–Singer correspondence, there is a bijection between connections on $E_{K}$ and holomorphic structures on $E_{G}$, the correspondence given in Theorem 2.7 can be interpreted by saying that in the $\curly{G}$-orbit of a polystable $G$-Higgs bundle $(J_{E_{0}},\varphi_{0})$ one can find another Higgs bundle $(J_{E},\varphi)$ whose corresponding pair $(A,\varphi)$ satisfies $F_{A}-[\varphi,\tau(\varphi)]=0$, and this is unique up to gauge transformations in $\curly K$}.
-2.5. Hitchin equations and hyperkähler structure
+2.5. Hitchin’s equations and hyperkähler structure
 Coming back to the setup of Section 2.4, let $E_{K}$ be a smooth principal $K$-bundle over $X$, and let $E_{G}$ the principal $G$-bundle obtained by extension of structure group.
@@
 $J_{\curly C}(\alpha)=i\alpha,\;\;\;\mbox{for}\;\;J_{E}\in{\curly C}\;\;\mbox{and}\;\;\alpha\in T_{J_{E}}{\curly C}=\Omega^{0,1}(X,E_{G}(\mathfrak{g})).$
-The Chern–Singer correspondence [Sin, W] establishes an isomorphism
-(2.16) ${\curly A}$ $\longrightarrow$ ${\curly C}$
-(2.17) $A$ $\longmapsto$ $J_{A}$
+As mentioned above, the Chern–Singer correspondence establishes an isomorphism
+(2.13) $\begin{array}[]{rcl}{\curly A}&\longrightarrow&{\curly C}\\ A&\longmapsto&J_{A}.\end{array}$
 The corresponding tangent spaces are in bijection under the map
-(2.18) $\Omega^{0,1}(X,E_{G}(\mathfrak{g}))$ $\longrightarrow$ $\Omega^{1}(X,E_{K}(\mathfrak{k}))$
-(2.19) $\alpha$ $\longmapsto$ $a:=\alpha-\tau(\alpha)$
+(2.14) $\begin{array}[]{rcl}\Omega^{0,1}(X,E_{G}(\mathfrak{g}))&\longrightarrow&\Omega^{1}(X,E_{K}(\mathfrak{k}))\\ \alpha&\longmapsto&a:=\alpha-\tau(\alpha).\end{array}$
 Under this identification $J_{\curly C}$ defines a complex structure $J_{\curly A}$ on $\curly A$}.
@@
 By Atiyah–Bott [AB] one has that the moment map for the action of $\curly K$} on $\curly A$} is given by
-(2.20) ${\curly A}$ $\longrightarrow$ $\Omega^{2}(X,E_{K}(\mathfrak{k}))$
-(2.21) $A$ $\longmapsto$ $F_{A}.$
+(2.15) $\begin{array}[]{rcl}{\curly A}&\longrightarrow&\Omega^{2}(X,E_{K}(\mathfrak{k}))\\ A&\longmapsto&F_{A}.\end{array}$
 Now, let us denote $\Omega=\Omega^{1,0}(X,E_{G}(\mathfrak{g}))$.
@@
 The action of $\curly K$} on ${\curly A}\times\Omega$ preserves these symplectic and complex structures and there is a moment map given by ([Hi1])
-(2.22) $\begin{array}[]{ccc}{\curly A}\times\Omega&\longrightarrow&\Omega^{2}(X,E_{K}(\mathfrak{k}))\\ (A,\varphi)&\mapsto&F_{A}-[\varphi,\tau(\varphi)].\end{array}$
+(2.16) $\begin{array}[]{rcl}{\curly A}\times\Omega&\longrightarrow&\Omega^{2}(X,E_{K}(\mathfrak{k}))\\ (A,\varphi)&\longmapsto&F_{A}-[\varphi,\tau(\varphi)].\end{array}$
 Let us denote $J_{1}:=J_{\curly A}+J_{\Omega}$.
 Via the identification ${\curly A}\cong{\curly C}$, we have for $\alpha\in\Omega^{0,1}(X,E_{G}(\mathfrak{g}))$ and $\eta\in\Omega^{1,0}(X,E_{G}(\mathfrak{g}))$ the following three complex structures on ${\curly A}\times\Omega$:
-$\begin{array}[]{ccc}J_{1}(\alpha,\eta)&=&(i\alpha,i\eta)\\ J_{2}(\alpha,\eta)&=&(-i\tau(\eta),i\tau(\alpha))\\ J_{3}(\alpha,\eta)&=&(\tau(\eta),-\tau(\alpha)),\end{array}$
-These complex structures correspond to the complex structures $\mathbf{I},\mathbf{J}$ and $\mathbf{K}$ in (2.8) via the the usual Chern–Singer correspondence.
+$J_{1}(\alpha,\eta)$ $=(i\alpha,i\eta),$
+$J_{2}(\alpha,\eta)$ $=(-i\tau(\eta),i\tau(\alpha)),$
+$J_{3}(\alpha,\eta)$ $=(\tau(\eta),-\tau(\alpha)).$
+These complex structures correspond to the complex structures $\mathbf{I},\mathbf{J}$ and $\mathbf{K}$ in (2.7) via the Chern–Singer correspondence.
 Consequently, $J_{i}$, $i=1,2,3$, satisfy the quaternion relations, and define a hyperkähler structure on ${\curly A}\times\Omega$, with symplectic structures $\omega_{i}$, $i=1,2,3$, where $\omega_{1}=\omega_{\curly A}+\omega_{\Omega}$.
@@
 $\mu_{1}(A,\varphi)=F_{A}-[\varphi,\tau(\varphi)],\;\;\;\mu_{2}(A,\varphi)=\mathop{{Re}}\nolimits(\bar{\partial}_{E}\varphi),\;\;\;\mu_{3}(A,\varphi)=\mathop{{Im}}\nolimits(\bar{\partial}_{E}\varphi).$
-We thus have that ${{\boldsymbol{\mu}}}^{-1}({0})/{\curly K}$ is the moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$ of solutions to Hitchin equations (2.15).
+We thus have that ${{\boldsymbol{\mu}}}^{-1}({0})/{\curly K}$ is the moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$ of solutions to Hitchin’s equations (2.12).
 In particular, if we consider the set ${{\boldsymbol{\mu}}}_{*}^{-1}({0})$ of irreducible solutions (equivalently, smooth) one has that
 ${{\boldsymbol{\mu}}}_{*}^{-1}({0})/{\curly K}$
-is a hyperkähler manifold which, by Theorem 2.7, is homeomorphic to the subvariety of smooth points of the moduli space $\mathcal{M}(G)$, consisting of stable and simple $G$-Higgs bundles on $E_{G}$.
+is a hyperkähler manifold which, by Theorem 2.7, is homeomorphic to the subvariety of smooth points of the moduli space $\mathcal{M}^{Higgs}(G)$, consisting of stable and simple $G$-Higgs bundles on $E_{G}$.
 2.6. Non-abelian Hodge correspondence
@@
 One can easily see that under the affine map
-$\begin{array}[]{ccc}{\curly A}\times\Omega&\longrightarrow&{\curly D}\\ (A,\varphi)&\longmapsto&A-i\varphi+i\tau(\varphi)\end{array}$
+$\begin{array}[]{rcl}{\curly A}\times\Omega&\longrightarrow&{\curly D},\\ (A,\varphi)&\longmapsto&A-i\varphi+i\tau(\varphi),\end{array}$
 ${\curly A}\times\Omega$ with complex structure $J_{2}$ corresponds to $\curly D$} with complex structure $\mathbf{J}$ (see [Hi1]).
@@
 The symplectic quotient construction can also be extended to the holomorphic situation (see e.g.
-[K]) to obtain the holomorphic symplectic quotient $\{(J_{E},\varphi)\;:\;\bar{\partial}_{E}\varphi=0\}/\curly{G}$.
+[Ko]) to obtain the holomorphic symplectic quotient $\{(J_{E},\varphi)\;:\;\bar{\partial}_{E}\varphi=0\}/\curly{G}$.
 What Theorem 2.7 says is that for a class $(J_{E},\varphi)]$ in this quotient to have a representative (unique up to $K$-gauge) satisfying $\mu_{1}=0$ it is necessary and sufficient that the pair $(J_{E},\varphi)$ be polystable.
@@
 3.1. Kähler fibrations
-In this section we recall some basic aspects of the theory of Kähler fibrations, which we will use later. We follow closely [GLS, M].
-A Kähler fibration is a holomorphic fibre bundle $\curly{X}\to\curly{J}$, with typical fiber a Kähler manifold $({\curly D},\omega_{\curly D})$, and such that transition functions between local holomorphic trivializations are contained in the group of Kähler isometries of $\curly D$}.
+In this section we recall some basic aspects of the theory of Kähler fibrations, which we will use later. We follow closely [GLS, Mu].
+A Kähler fibration is a holomorphic fibre bundle $\curly{X}\to\curly{J}$, with typical fibre a Kähler manifold $({\curly D},\omega_{\curly D})$, and such that transition functions between local holomorphic trivializations are contained in the group of Kähler isometries of $\curly D$}.
 Equivalently, $\curly{X}$ admits a smoothly varying Kähler structure on the fibres, that we shall denote $\widehat{\boldsymbol{\omega}}$.
@@
 […]
-The following basic result shows that the existence of closed $(1,1)$-form restricting on the fibres to $\widehat{\boldsymbol{\omega}}$ is indeed a sufficient condition for the existence of a Kähler connection.
-The proof follows similarly as in [GLS, Theorem 1.2.4] and is ommited.
+The following basic result shows that the existence of a closed $(1,1)$-form restricting on the fibres to $\widehat{\boldsymbol{\omega}}$ is indeed a sufficient condition for the existence of a Kähler connection.
+The proof follows similarly to [GLS, Theorem 1.2.4] and is omited.
 Theorem 3.2.
@@
 If this is the case, the horizontal subspace of the induced connection is given by $H^{\boldsymbol{\sigma}}$ in (3.1), and $\boldsymbol{\sigma}$ is a coupling form for $\Gamma^{\boldsymbol{\sigma}}$.
-The existence a coupling form is a non-trivial question, related to the topology on the fibration.
+The existence of a coupling form is a non-trivial question, related to the topology of the fibration.
 In the next result we recall a remarkable identity which relates the curvature of $\Gamma$ with the horizontal part of any coupling $\boldsymbol{\sigma}$ (see [GLS, Eq.
 (1.12)]).
-Recall that the curvature of an Ehresmann connection $\Gamma$ on $\curly{X}\to\curly{J}$ is the basic two-form $F_{\Gamma}\in\Omega^{2}(\curly{X},V\curly{X})$ defined by
+Recall that the curvature of an Ehresmann connection $\Gamma$ on $\curly{X}\to\curly{J}$ is the basic 2-form $F_{\Gamma}\in\Omega^{2}(\curly{X},V\curly{X})$ defined by
 $F_{\Gamma}(v_{1},v_{2})=-\Gamma[v_{1}^{\Gamma},v_{2}^{\Gamma}],$
@@
 Proposition 3.3.
-Assume the hypothesis of Theorem 3.2. Then,
+Assume the hypothesis of Theorem 3.2. Then
 (3.2) $\widehat{\boldsymbol{\omega}}(F_{\Gamma}(v_{1},v_{2}),\Gamma\cdot)=-d(\boldsymbol{\sigma}(v_{1}^{\Gamma},v_{2}^{\Gamma}))(\Gamma\cdot).$
@@
 The idea is that, under this hypothesis, the vertical vector field $F_{\Gamma}(v_{1},v_{2})$ is Hamiltonian, and hence it determines up to a constant the vertical variation of the function $\boldsymbol{\sigma}(v_{1}^{\Gamma},v_{2}^{\Gamma})$.
-More explicitly, provided that there exists a coherent choice of fibrewise Hamiltonian function $\mu(v_{1},v_{2})$ for $F_{\Gamma}(v_{1},v_{2})$, for any pair of horizontal vector fields $v_{1},v_{2}$, that is,
+More explicitly, provided that there exists a coherent choice of a fibrewise Hamiltonian function $\mu(v_{1},v_{2})$ for $F_{\Gamma}(v_{1},v_{2})$, for any pair of horizontal vector fields $v_{1},v_{2}$, that is,
 $d\mu(v_{1},v_{2})=\widehat{\boldsymbol{\omega}}(F_{\Gamma}(v_{1},v_{2}),)$
 the natural choice of coupling form is
-(3.3) $\boldsymbol{\sigma}=\widehat{\boldsymbol{\omega}}(\Gamma,\Gamma)-\mu(v_{1},v_{2}).$
-This is the case, for instance, if the connection $\Gamma$ has holonomy contained in a subgroup $K$ of Hamiltonian isommetries of the fibre $\curly D$}, and there is a $K$-equivariant moment map $\mu\colon{\curly D}\to\mathfrak{k}^{*}$.
+(3.3) $\boldsymbol{\sigma}=\widehat{\boldsymbol{\omega}}(\Gamma,\Gamma)-\mu,$
+where $\mu$ is regarded as a basic 2-form.
+This is the case, for instance, if the connection $\Gamma$ has holonomy contained in a subgroup $K$ of Hamiltonian isometries of the fibre $\curly D$}, and there is a $K$-equivariant moment map $\mu\colon{\curly D}\to\mathfrak{k}^{*}$.
 A fundamental question in the theory of Kähler fibrations is whether there exists a Kähler metric $\omega{X}$ on $\curly{X}$ which restricts to $\widehat{\boldsymbol{\omega}}$ on the fibres.
@@
 $\boldsymbol{\omega}_{\alpha}=\boldsymbol{\omega}{J}+\alpha\boldsymbol{\sigma}$
-for a small coupling constant $0<\alpha\ll 1$.
+for a small ‘coupling constant’ $0<\alpha\ll 1$.
 In the present paper, we are interested in the study of Kähler fibrations satisfying the hypothesis of Theorem 3.2.
-We will find situations in which the pre-symplectic manifold $(\curly{X},\boldsymbol{\sigma})$ admits a Hamiltonian action of a real Lie group $\widetilde{\curly{G}}$, preserving the holomorphic fibration structure.
-This will have an impact in the structure of the group of symmetries, which will typically appear as a non-trivial extension
+We will find situations in which the presymplectic manifold $(\curly{X},\boldsymbol{\sigma})$ admits a Hamiltonian action of a real Lie group $\widetilde{\curly{G}}$, preserving the holomorphic fibration structure.
+This will have an impact on the structure of the group of symmetries, which will typically appear as a non-trivial extension
 (3.4) $1\to\curly{G}\longrightarrow{}\widetilde{\curly{G}}\longrightarrow\curly{H}\to 1,$
 of a (real) subgroup $\curly{H}$ of the holomorphic automorphisms of the base $\curly{J}$ by a subgroup of Kähler isometries $\curly{G}$ of the typical fibre $\curly D$}.
-Previous work on this type of reductions, in the context of gauge theory, can be found in [AGG, AGG3, AGGP, AGGPY].
-3.2. Kähler connection and norm squared of the Higgs field
+Previous work on this type of reductions, in the context of gauge theory, can be found in [AGG1, AGG3, AGGP, AGGPY].
+3.2. The Kähler connection and the squared norm of the Higgs field
 We fix a smooth oriented compact surface $\Sigma$.
@@
 Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$.
-Let $\curly D$} be the space of connections on $E_{G}$, equipped with the constant complex structure (2.1).
-Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation.
+Let $\curly D$} be the space of connections on $E_{G}$, equipped with the constant complex structure (2.1), and $\curly{J}$ the space of complex structures on $\Sigma$ that are compatible with the given orientation.
 Consider the space
-(3.5) $\curly{X}=\curly{J}\times{\curly D}$
+(3.5) $\curly{X}=\curly{J}\times{\curly D},$
 endowed with the product complex structure
@@
 We should emphasize that, even though the holomorphic fibration structure on $\curly{X}$ is trivial, $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$ has a non-trivial dependence on the base $\curly{J}$.
-In order to give a more explicit description of $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$, we consider $h\in\Omega^{0}(E_{G}(G/K))$ a reduction of structure group of $E_{G}$ to $K$, and let $E_{K}$ be the corresponding principal $K$-bundle.
-From the decomposition $\mathfrak{g}=\mathfrak{k}\oplus i\mathfrak{k}$ one has the identification ${\curly D}\cong{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ in (2.7), which induces a biholomorphism
-(3.6) $\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k})),$
-where the right hand side is endowed with the product complex structure
+In order to give a more explicit description of $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$, we consider a reduction $h\in\Omega^{0}(E_{G}(G/K))$ of the structure group of $E_{G}$ to $K$, and let $E_{K}$ be the corresponding principal $K$-bundle.
+From the decomposition $\mathfrak{g}=\mathfrak{k}\oplus i\mathfrak{k}$ one has the identification ${\curly D}\cong{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$ in (2.6), which induces a biholomorphism
+(3.6) $\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k})),$
+where the right-hand side is endowed with the product complex structure
 (3.7) $\mathbb{J}(\dot{J},a,\psi)=(J\dot{J},-\dot{\psi},a).$
@@
 (3.8) $\widehat{\boldsymbol{\omega}}_{\mathbf{J}|(J,A,\psi)}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2}))=\int_{\Sigma}B(a_{1}\wedge J\dot{\psi}_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge Ja_{2}).$
-One can clearly see that $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$ has a non-trivial dependence on $J\in\curly{J}$, and hence it defines a non-trivial structure of Kähler fibration on $\curly{X}\to\curly{J}$.
-The aim of this section is to prove that $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}})_{\mathbf{J}}$ admits a Kähler Ehresmann connection, following Theorem 3.2.
+It is clear that $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$ has a non-trivial dependence on $J\in\curly{J}$, and hence defines a structure of a non-trivial Kähler fibration on $\curly{X}\to\curly{J}$.
+The aim of this section is to prove that $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ admits a Kähler Ehresmann connection, following Theorem 3.2.
 For this, we follow a suggestion by N.
@@
 [Hi1, Section 9]) and define a real $(1,1)$-form $\boldsymbol{\sigma}^{\mathbb{J}}\in\Omega^{1,1}(\curly{X},\mathbb{R})$ by
-(3.9) $\boldsymbol{\sigma}^{\mathbb{J}}=i\partial_{\mathbb{J}}\bar{\partial}_{\mathbb{J}}\|\psi\|^{2}_{L^{2}}$
-where $\|\psi\|^{2}_{L^{2}}$ denotes the $J$-dependent $L^{2}$-norm of the unitary Higgs field $\psi$, namely
+(3.9) $\boldsymbol{\sigma}^{\mathbb{J}}=i\partial_{\mathbb{J}}\bar{\partial}_{\mathbb{J}}\|\psi\|^{2}_{L^{2}},$
+where $\|\psi\|^{2}_{L^{2}}$ denotes the $J$-dependent $L^{2}$-norm of the ‘unitary Higgs field’ $\psi$, namely
 $\|\psi\|^{2}_{L^{2}}=\int_{\Sigma}B(\psi\wedge J\psi).$
-In the next result we calculate an explicit formula for $\boldsymbol{\sigma}^{\mathbb{J}}$.
+In the next result, we derive an explicit formula for $\boldsymbol{\sigma}^{\mathbb{J}}$.
 Lemma 3.4.
-(3.10) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}_{|(J,A,\psi)}((\dot{J}_{1},a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge(J\dot{\psi}_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B(a_{2}\wedge(J\dot{\psi}_{1}-\psi(\dot{J}_{1})))\\ &-\int_{\Sigma}B\left(\left(J\dot{\psi}_{1}-\tfrac{1}{2}\psi(\dot{J}_{1})\right)\wedge\psi(\dot{J}_{2})\right)\\ &+\int_{\Sigma}B\left(\left(J\dot{\psi}_{2}-\tfrac{1}{2}\psi(\dot{J}_{2})\right)\wedge\psi(\dot{J}_{1})\right)\end{split}$
+(3.10) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}_{|(J,A,\psi)}((\dot{J}_{1},a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge(J\dot{\psi}_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B(a_{2}\wedge(J\dot{\psi}_{1}-\psi(\dot{J}_{1})))\\ &-\int_{\Sigma}B((J\dot{\psi}_{1}-\tfrac{1}{2}\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2}))\\ &+\int_{\Sigma}B((J\dot{\psi}_{2}-\tfrac{1}{2}\psi(\dot{J}_{2}))\wedge\psi(\dot{J}_{1}))\end{split}$
 Proof.
@@
 Lemma 3.5.
-Given $C\in\operatorname{End}T\Sigma$ and $\psi_{1},\psi_{2}\in\Omega^{1}(X,E_{K}(\mathfrak{k}))$, the following pointwise identity holds:
+For $C\in\operatorname{End}T\Sigma$ and $\psi_{1},\psi_{2}\in\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$, the following pointwise identity holds:
 $B(\psi_{1}(C)\wedge\psi_{2})+B(\psi_{1}\wedge\psi_{2}(C))=0.$
 Proof.
-By direct calculation at a point $x\in\Sigma$, if $C=v\otimes\gamma\in T_{x}\Sigma\otimes T^{*}_{x}\Sigma$, by dimensional reasons we have
+By direct calculation at a point $x\in\Sigma$, if $C=v\otimes\gamma\in T_{x}\Sigma\otimes T^{*}_{x}\Sigma$, we see by dimensional reasons that
 $0$ $=\gamma\wedge i_{v}B(\psi_{1}\wedge\psi_{2})$
-$=\gamma\wedge(B(\psi_{1}(v),\psi_{2})-B(\psi_{1},\psi_{2}(v)))=B(\psi_{1}(v)\gamma\wedge\psi_{2})+B(\psi_{1}\wedge\psi_{2}(v)\gamma).$
-Our next result proves the existence of a natural Kähler Ehresmann connection in our Kähler fibration $\curly{X}\to\curly{J}$, with coupling form $\boldsymbol{\sigma}^{\mathbb{J}}$ and curvature given (essentially) by the unitary Higgs field $\psi$.
+$=\gamma\wedge(B(\psi_{1}(v),\psi_{2})-B(\psi_{1},\psi_{2}(v)))=B(\psi_{1}(v)\gamma\wedge\psi_{2})+B(\psi_{1}\wedge\psi_{2}(v)\gamma).\qed$
+Our next result proves the existence of a natural Kähler Ehresmann connection on the Kähler fibration $\curly{X}\to\curly{J}$, whose coupling form is $\boldsymbol{\sigma}^{\mathbb{J}}$ and whose curvature is essentially given by the ‘unitary Higgs field’ $\psi$.
 Proposition 3.6.
@@
 $H^{\mathbb{J}}=\{v\in T\curly{X}\;|\;i_{v}\boldsymbol{\sigma}^{\mathbb{J}}_{|V\curly{X}}=0\},$
-where $\sigma$ is defined by (3.9). More explicitly,
+where $\boldsymbol{\sigma}^{\mathbb{J}}$ is defined by (3.9). More explicitly,
 (3.11) $H^{\mathbb{J}}_{|(J,A,\psi)}=\{(\dot{J},\psi(\dot{J}),(J\psi)(\dot{J}))\;|\;\dot{J}\in T_{J}\curly{J}\}.$
-Furthermore, $H^{\mathbb{J}}$ is preserved by $\mathbb{J}$ and the curvature $F_{\mathbb{J}}:=F_{\Gamma^{\mathbb{J}}}\in\Omega^{2}(\curly{X},V\curly{X})$ of $\Gamma$ is of type $(1,1)$ and given explicitly by
-(3.12) $\begin{split}(F_{\mathbb{J}})_{|(J,A,\psi)}(v_{1},v_{2})&=(0,\psi([\dot{J}_{2},\dot{J}_{1}]),0).\end{split}$
+Furthermore, $H^{\mathbb{J}}$ is preserved by $\mathbb{J}$ and the curvature $F_{\mathbb{J}}:=F_{\Gamma^{\mathbb{J}}}\in\Omega^{2}(\curly{X},V\curly{X})$ of $\Gamma^{\mathbb{J}}$ is of type $(1,1)$, and given explicitly by
+(3.12) $(F_{\mathbb{J}})_{|(J,A,\psi)}(v_{1},v_{2})=(0,\psi([\dot{J}_{2},\dot{J}_{1}]),0),$
 for any pair of horizontal vector fields $v_{1},v_{2}\in H^{\mathbb{J}}$ covering $\dot{J}_{1},\dot{J}_{2}$, respectively.
@@
 The existence of $\Gamma$ follows from Theorem 3.2 combined with Lemma 3.4, which implies that
-(3.13) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge J\dot{\psi}_{2})+\int_{\Sigma}B(J^{2}\dot{\psi}_{1}\wedge Ja_{2})\\ &=\widehat{\boldsymbol{\omega}}_{\mathbf{J}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2})).\end{split}$
-where we ommit the evaluation at the point $(J,A,\psi)$ for simplicity in the notation. Contracting now $\boldsymbol{\sigma}^{\mathbb{J}}$ with a vertical vector field
-$\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}((0,a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge(J\dot{\psi}_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B\left(J\dot{\psi}_{1}\wedge(\psi(\dot{J}_{2})-a_{2})\right)\end{split}$
-and therefore $(\dot{J}_{2},a_{2},\dot{\psi}_{2}))\in H^{\mathbb{J}}$ if and only if
+(3.13) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge J\dot{\psi}_{2})+\int_{\Sigma}B(J^{2}\dot{\psi}_{1}\wedge Ja_{2})\\ &=\widehat{\boldsymbol{\omega}}_{\mathbf{J}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2})),\end{split}$
+where we omit the evaluation at the point $(J,A,\psi)$ for simplicity in the notation. Contracting now $\boldsymbol{\sigma}^{\mathbb{J}}$ with a vertical vector field
+$\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}((0,a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge(J\dot{\psi}_{2}-\psi(\dot{J}_{2})))-\int_{\Sigma}B(J\dot{\psi}_{1}\wedge(\psi(\dot{J}_{2})-a_{2})),\end{split}$
+and therefore $(\dot{J}_{2},a_{2},\dot{\psi}_{2})\in H^{\mathbb{J}}$ if and only if
 $a_{2}=\psi(\dot{J}_{2}),\qquad\dot{\psi}_{2}=(J\psi)(\dot{J}_{2}).$
@@
 Finally, evaluating $\boldsymbol{\sigma}^{\mathbb{J}}$ in a pair of horizontal vector fields $v_{j}=(\dot{J}_{j},\psi(\dot{J}_{j}),(J\psi)(\dot{J}_{j}))$, we have
-$\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2})&=-\int_{\Sigma}B\left(\left(\psi(\dot{J}_{1})-\tfrac{1}{2}\psi(\dot{J}_{1})\right)\wedge\psi(\dot{J}_{2})\right)+\int_{\Sigma}B\left(\left(\psi(\dot{J}_{2})-\tfrac{1}{2}\psi(\dot{J}_{2})\right)\wedge\psi(\dot{J}_{1})\right)\\ &=-\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2}))\end{split}$
+$\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2})&=-\int_{\Sigma}B((\psi(\dot{J}_{1})-\tfrac{1}{2}\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2}))+\int_{\Sigma}B((\psi(\dot{J}_{2})-\tfrac{1}{2}\psi(\dot{J}_{2}))\wedge\psi(\dot{J}_{1}))\\ &=-\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})),\end{split}$
 where we have used that
@@
 Consequently,
-$d(\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2}))((0,a,\dot{\psi}))=-\int_{\Sigma}B(\dot{\psi}(\dot{J}_{1}))\wedge\psi(\dot{J}_{2}))-\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\dot{\psi}(\dot{J}_{2})).$
-On the other hand, applying Lemma 3.5,
-$\widehat{\boldsymbol{\omega}}_{\mathbf{J}}((0,(J\psi)([\dot{J}_{1},\dot{J}_{2}],0),(0,a,\dot{\psi}))$ $=\int_{\Sigma}B((J\psi)([\dot{J}_{1},\dot{J}_{2}])\wedge J\dot{\psi})$
+$d(\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2}))((0,a,\dot{\psi}))=-\int_{\Sigma}B(\dot{\psi}(\dot{J}_{1})\wedge\psi(\dot{J}_{2}))-\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\dot{\psi}(\dot{J}_{2})).$
+On the other hand, applying Lemma 3.5, we obtain
+$\widehat{\boldsymbol{\omega}}_{\mathbf{J}}((0,(J\psi)([\dot{J}_{1},\dot{J}_{2}]),0),(0,a,\dot{\psi}))$ $=\int_{\Sigma}B((J\psi)([\dot{J}_{1},\dot{J}_{2}])\wedge J\dot{\psi})$
 $=\int_{\Sigma}B(J(\psi([\dot{J}_{1},\dot{J}_{2}]))\wedge J\dot{\psi})$
@@
 $=-\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\dot{\psi}(\dot{J}_{2}))+\int_{\Sigma}B(\psi(\dot{J}_{2})\wedge\dot{\psi}(\dot{J}_{1}))$
-$=d(\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2}))((0,a,\dot{\psi})),$
-which proves the last part of the statement, by application of Proposition 3.3. ∎
+$=d(\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2}))(0,a,\dot{\psi}),$
+which proves the last part of the statement, by Proposition 3.3. ∎
 To finish this section, we provide a formula for the coupling form $\boldsymbol{\sigma}^{\mathbb{J}}$ adapted to its associated connection $\Gamma^{\mathbb{J}}$.
@@
 For any tangent vectors $v_{j}=(\dot{J}_{j},a_{j},\dot{\psi}_{j}),v=(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$ one has
-(3.14) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2})&=\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}_{\mathbb{J}}(v,v)&=\int_{\Sigma}B((a-\psi(\dot{J}))\wedge J(a-\psi(\dot{J})))\\ &+\int_{\Sigma}B((\dot{\psi}-(J\psi)(\dot{J}))\wedge J(\dot{\psi}-(J\psi)(\dot{J}))\\ &-\int_{\Sigma}B(\psi(\dot{J}))\wedge J(\psi(\dot{J}))).\end{split}$
+(3.14) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2})&=\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}_{\mathbb{J}}(v,v)&=\int_{\Sigma}B((a-\psi(\dot{J}))\wedge J(a-\psi(\dot{J})))\\ &+\int_{\Sigma}B((\dot{\psi}-(J\psi)(\dot{J}))\wedge J(\dot{\psi}-(J\psi)(\dot{J})))\\ &-\int_{\Sigma}B(\psi(\dot{J})\wedge J(\psi(\dot{J}))).\end{split}$
 In particular, given a horizontal vector field $v\in H^{\mathbb{J}}$ at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$, one has
-(3.15) $\begin{split}\mathbf{g}_{\mathbb{J}}(v,v)=-\int_{\Sigma}B\left(\psi(\dot{J})\wedge J(\psi(\dot{J}))\right).\end{split}$
+(3.15) $\begin{split}\mathbf{g}_{\mathbb{J}}(v,v)=-\int_{\Sigma}B(\psi(\dot{J})\wedge J(\psi(\dot{J}))).\end{split}$
 Consequently, $\mathbf{g}_{\mathbb{J}}$ is negative semi-definite along the horizontal directions of $\Gamma^{\mathbb{J}}$.
-3.3. Minimal coupling and Hamiltonian action
+3.3. Minimal coupling and the Hamiltonian action
 Let $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ be the Kähler fibration defined in Section 3.2.
-In this section we undertake minimal couplings for the connection $\Gamma^{\mathbb{J}}$ constructed in Proposition 3.6.
+In this section we apply the minimal-coupling construction to the connection $\Gamma^{\mathbb{J}}$ introduced in Proposition 3.6.
 We will see that the corresponding (pre)symplectic structures on $\curly{X}$ admit a Hamiltonian action by a suitable extended gauge group $\widetilde{{\curly K}}$, naturally associated to a choice of symplectic structure $\omega$ on $\Sigma$.
-This will allow us, in Section 4, to construct pseudo-Kähler structures on the universal moduli space of solutions of the harmonicity equations (2.14) over Teichmuller space.
-We start by introducing the Kähler structure we shall consider, up to sign, in the base of of the fibration $\curly{X}\to\curly{J}$.
-Following Donaldson [Do2] and Fujiki [F], we fix a symplectic (volume) form $\omega$ on $\Sigma$ compatible with the given orientation.
-The space $\curly{J}$ of complex structures $J$ on $\Sigma$ compatible with the orientation (and hence with $\omega$) is an infinite dimensional Kähler manifold, with complex structure $\mathbb{J}{J}\colon T_{J}\curly{J}\to T_{J}\curly{J}$ and Kähler form $\boldsymbol{\omega}{J}$ given, respectively, by
+This will allow us, in Section 4, to construct pseudo-Kähler structures on the universal moduli space of solutions of the harmonicity equations (2.11) over the Teichmüller space.
+We start by introducing the Kähler structure that we will consider, up to sign, on the base of the fibration $\curly{X}\to\curly{J}$.
+Following Donaldson [Do2] and Fujiki [Fu], we fix a symplectic (volume) form $\omega$ on $\Sigma$ compatible with the given orientation.
+The space $\curly{J}$ of complex structures $J$ on $\Sigma$ compatible with the orientation (and hence with $\omega$) is an infinite-dimensional Kähler manifold, with complex structure $\mathbb{J}{J}\colon T_{J}\curly{J}\to T_{J}\curly{J}$ and Kähler form $\boldsymbol{\omega}{J}$ given, respectively, by
 (3.16) $\mathbb{J}{J}\dot{J}:=J\dot{J}\;\;\text{ and }\;\;\boldsymbol{\omega}{J}(\dot{J}_{1},\dot{J}_{2}):=\frac{1}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega,$
 for $\dot{J}_{1},\dot{J}_{2}\in T_{J}\curly{J}$.
-Note that, by dimensional reasons, any $\dot{J}\in T_{J}\curly{J}$, regarded as an endomorphism $\dot{J}\colon T\Sigma\to T\Sigma$, is symmetric with respect to the induced metric $g=\omega(\cdot,J\cdot)$
+Note that, for dimensional reasons, any $\dot{J}\in T_{J}\curly{J}$, regarded as an endomorphism $\dot{J}\colon T\Sigma\to T\Sigma$, is symmetric with respect to the induced metric $g=\omega(\cdot,J\cdot)$.
 Let $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ be the Kähler fibration defined in Section 3.2.
-Given $\alpha>0$ a real coupling constant and $\varepsilon\in\{-1,1\}$, the family of minimal coupling symplectic structures of our interest is defined by
+Given a real ‘coupling constant’ $\alpha>0$ and $\varepsilon\in\{-1,1\}$, the family of minimal coupling symplectic structures of our interest is defined by
 (3.17) $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}=\varepsilon\boldsymbol{\omega}{J}+\alpha\boldsymbol{\sigma}^{\mathbb{J}},$
@@
 (3.18) $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}:=\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(,\mathbb{J})=\varepsilon\mathbf{g}_{\curly{J}}+\alpha\mathbf{g}_{\mathbb{J}}.$
-Applying Corollary 3.7, $\boldsymbol{\omega}_{\alpha}^{-1}$ is negative-definite along the horizontal subspace $H^{\mathbb{J}}\subset T\curly{X}$.
+Applying Corollary 3.7, $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,-1}$ is negative-definite along the horizontal subspace $H^{\mathbb{J}}\subset T\curly{X}$.
 The analysis for the case $\varepsilon=1$ is much more subtle, as we can see from the next result.
 Lemma 3.8.
-Given a horizontal vector field $v\in H^{\mathbb{J}}$ at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$, one has
+Let $v\in H^{\mathbb{J}}$ be a horizontal vector field at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$. Then
 (3.19) $\begin{split}\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,v)=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}\dot{J})\omega-\alpha\int_{\Sigma}B\left(\psi(\dot{J})\wedge J(\psi(\dot{J}))\right).\end{split}$
-Consequently, for any $\alpha>0$ one has
+Consequently, for any $\alpha>0$:
   1. (1)
-    For $\varepsilon=-1$, $\mathbf{g}_{\alpha}^{\varepsilon}$ is negative definite along $H^{\mathbb{J}}$. In particular, $\boldsymbol{\omega}_{\alpha}^{-1}$ is a non-degenerate symplectic structure.
+    If $\varepsilon=-1$, $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is negative definite along $H^{\mathbb{J}}$.
+In particular, $\boldsymbol{\omega}_{\alpha}^{-1}$ is a non-degenerate symplectic structure.
   2. (2)
-    For $\varepsilon=1$ and $\psi\neq 0$, $\mathbf{g}_{\alpha}^{\varepsilon}$ changes signature along the line $(J,A,\lambda\psi)\in\curly{X}$, for $\lambda\in\mathbb{R}$.
+    If $\varepsilon=1$ and $\psi\neq 0$, $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ changes signature along the line $(J,A,\lambda\psi)\in\curly{X}$, for $\lambda\in\mathbb{R}$.
 Proof.
-The case $\varepsilon=-1$ is a direct consequence of Corollary 3.7.
-For the case $\varepsilon=1$, let $(J,A,\psi)\in\curly{X}$.
-Then, for any $\dot{J}\in T_{J}\curly{J}$, we have a pointwise equalities
+The case $\varepsilon=-1$ is a direct consequence of Corollary 3.7. For the case $\varepsilon=1$, let $(J,A,\psi)\in\curly{X}$. Then, for any $\dot{J}\in T_{J}\curly{J}$, we have pointwise identities
 $\operatorname{tr}(\dot{J}\dot{J})=C|\dot{J}|^{2},\qquad\Lambda_{\omega}B\left(\psi(\dot{J})\wedge J(\psi(\dot{J}))\right)=C^{\prime}|\psi|^{2}|\dot{J}|^{2},$
@@
   • null at the horizontal bundle $H^{\mathbb{J}}$ at the points $(J,A,\lambda_{0}\psi)\in\curly{X}$.
-The result follows now from the previous cases.
+The result follows now from the previous cases. ∎
 Our next goal is to prove that $(\curly{X},\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon})$ admits a Hamiltonian action by a suitable extended gauge group $\widetilde{{\curly K}}$, determined by $\omega$ and the reduction $E_{K}\subset E_{G}$.
@@
 (3.20) $1\to{\curly K}\longrightarrow{}\widetilde{{\curly K}}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\curly{H}\to 1,$
-where $p$ maps each $g\in\widetilde{{\curly K}}$ into the Hamiltonian symplectomorphism $p(g)\in\curly{H}$ that it covers, and so its kernel $\curly K$} is the gauge group of $E_{K}$, that is, the normal subgroup of $\widetilde{{\curly K}}$ consisting of automorphisms of $E_{K}$ covering the identity map on $\Sigma$.
-Similarly as in [AGG, Section 2.2], the group $\widetilde{{\curly K}}$ acts on $(\curly{X},\mathbb{J},\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon})$ preserving $\mathbb{J}$ and $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$, and covering the $\curly{H}$-action on $\curly{J}$ by push-forward.
-This last action preserves the Kähler structure on $\curly{J}$, and furthermore is Hamiltonian, with equivariant moment map $\mu\colon\curly{J}\to(\operatorname{Lie}\curly{H})^{*}$ given by
+where $p$ maps each $g\in\widetilde{{\curly K}}$ to the Hamiltonian symplectomorphism $p(g)\in\curly{H}$ that it covers, and so its kernel $\curly K$} is the gauge group of $E_{K}$, that is, the normal subgroup of $\widetilde{{\curly K}}$ consisting of automorphisms of $E_{K}$ covering the identity map on $\Sigma$.
+As in [AGG1, Section 2.2], the group $\widetilde{{\curly K}}$ acts on $(\curly{X},\mathbb{J},\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon})$, preserving both $\mathbb{J}$ and $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$, and covering the $\curly{H}$-action on $\curly{J}$ given by push-forward.
+This last action preserves the Kähler structure on $\curly{J}$ and is moreover Hamiltonian, with equivariant moment map $\mu\colon\curly{J}\to(\operatorname{Lie}\curly{H})^{*}$ given by
 (3.21) $\langle\mu(J),\eta_{f}\rangle=-\int_{\Sigma}fS_{J}\omega,$
-for $f\in C^{\infty}_{0}(\Sigma)$, identified with an element $\eta_{f}$ in $\operatorname{Lie}\curly{H}$, and $S_{J}$ is the scalar curvature of the metric $g=\omega(,J)$ defined by $\omega$ and $J$.
+where $f\in C^{\infty}_{0}(\Sigma)$ is identified with an element $\eta_{f}$ in $\operatorname{Lie}\curly{H}$, and $S_{J}$ is the scalar curvature of the metric $g=\omega(,J)$ determined by $\omega$ and $J$.
 Here, $C^{\infty}_{0}(\Sigma)$ denotes the space of smooth functions on $\Sigma$ such that $\int_{\Sigma}f\omega=0$.
-This fact was shown by Donaldson [Do2] and Fujiki [F] independently for symplectic manifolds of arbitrary dimensions, and it seems that was known to Quillen in the case of surfaces.
+This fact was established by Donaldson [Do2] and Fujiki [Fu] independently for symplectic manifolds of arbitrary dimension, and it appears to have been known to Quillen in the case of surfaces.
 Proposition 3.9.
@@
 (3.22) $\begin{split}\langle\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}(J,A,\psi),\zeta\rangle&=\alpha\int_{\Sigma}B(A\zeta+2\psi(Jp\zeta),d_{A}(J\psi))\\ &-\int_{\Sigma}f\left(\varepsilon S_{J}-\alpha\Lambda_{\omega}d\left(B\left(\Lambda_{\omega}F_{A},J\psi\right)\right)\right)\omega,\end{split}$
-where $A\zeta\in\Omega^{0}(X,E_{K}(\mathfrak{k}))$ denotes the vertical part of the $K$-equivariant vector field $\zeta$ on the total space of $E_{K}$, with respect to the connection $A$.
+where $A\zeta\in\Omega^{0}(\Sigma,E_{K}(\mathfrak{k}))$ denotes the vertical part of the $K$-equivariant vector field $\zeta$ on the total space of $E_{K}$, with respect to the connection $A$.
 Proof.
-Since $\sigma=\tfrac{1}{2}dd^{c}_{\mathbb{J}}\nu$, with $\nu(J,A,\psi):=\|\psi\|^{2}_{L^{2}}$ (see Lemma 3.4), and $d^{c}_{\mathbb{J}}\nu$ is preserved by the $\widetilde{{\curly K}}-$action, there exists an equivariant moment map given by
+Since $\sigma=\tfrac{1}{2}dd^{c}_{\mathbb{J}}\nu$, with $\nu(J,A,\psi):=\|\psi\|^{2}_{L^{2}}$ (see Lemma 3.4), and because $d^{c}_{\mathbb{J}}\nu$ is preserved by the $\widetilde{{\curly K}}-$action, there exists an equivariant moment map given by
 $\langle\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}(J,A,\psi),\zeta\rangle=-\varepsilon\int_{\Sigma}fS_{J}\omega-\frac{\alpha}{2}d^{c}_{\mathbb{J}}\nu(\zeta\cdot(J,A,\psi)),$
-where $\zeta\cdot(J,A,\psi)$ is the the infinitesimal action of $\zeta$ on $(J,A,\psi)$. Using the complex connection $D=A+i\psi$ it follows that
+where $\zeta\cdot(J,A,\psi)$ denotes the infinitesimal action of $\zeta$ on $(J,A,\psi)$. Using the complex connection $D=A+i\psi$, it follows that
 $\zeta\cdot D$ $=-d_{D}(D\zeta)-i_{y}F_{D}$
@@
 $=-2\int_{\Sigma}B(A\zeta+2\psi(Jy),d_{A}(J\psi))-2\int_{\Sigma}fd\left(B\left(\Lambda_{\omega}F_{A},J\psi\right)\right).$
-For the second equality, we have used that, for any vector field $y$ on $\sigma$, one has $L_{y}J=2i\bar{\partial}y^{1,0}-2i\partial y^{0,1}$ and therefore
-$B(J\psi\wedge\psi(L_{y}J))$ $=2B(\psi^{1,0}\wedge\psi(\bar{\partial}y^{1,0}))+cc.$
-$=2B(\psi^{1,0}\wedge(i_{y^{1,0}}(\bar{\partial}_{A}\psi^{1,0})+\bar{\partial}_{A}(i_{y^{1,0}}\psi)))+cc.$
-$=-2\bar{\partial}(B(\psi^{1,0},\psi(y^{1,0})))+4B(\bar{\partial}_{A}\psi^{1,0},\psi(y^{1,0}))+cc.$
-The $\widetilde{{\curly K}}-$action produces a ‘coupling term’ in the base $\Sigma$ (cf.
-Remark 4.2), which combines with the scalar curvature of the metric $g=\omega(,J)$.
-In particular, similarly as in [AGG], zeros of the moment map $\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}$ are given by solutions of the coupled system of equations
-(3.23) $\begin{split}d_{A}^{*}\psi&=0\\ \varepsilon S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
+For the second equality, we use the identity $L_{y}J=2i\bar{\partial}y^{1,0}-2i\partial y^{0,1}$ for any vector field $y$ on $\Sigma$, and therefore
+$B(J\psi\wedge\psi(L_{y}J))$ $=2B(\psi^{1,0}\wedge\psi(\bar{\partial}y^{1,0}))+\textnormal{c.c.}$
+$=2B(\psi^{1,0}\wedge(i_{y^{1,0}}(\bar{\partial}_{A}\psi^{1,0})+\bar{\partial}_{A}(i_{y^{1,0}}\psi)))+\textnormal{c.c.}$
+$=-2\bar{\partial}(B(\psi^{1,0},\psi(y^{1,0})))+4B(\bar{\partial}_{A}\psi^{1,0},\psi(y^{1,0}))+\textnormal{c.c.},$
+where as usual “c.c.” stands for complex conjugate. ∎
+The $\widetilde{{\curly K}}-$action produces a ‘coupling term’ on the base $\Sigma$ (cf.
+Remark 4.2), which interacts with the scalar curvature of the metric $g=\omega(,J)$.
+In particular, as in [AGG1], the zeros of the moment map $\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}$ correspond to the solutions of the coupled system of equations
+(3.23) $\begin{split}d_{A}^{*}\psi&=0,\\ \varepsilon S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
 where $\chi(\Sigma)$ is the Euler characteristic of $\Sigma$ and $V=\int_{\Sigma}\omega$ is the total volume.
-In the next section we combine these equations with the flatness condition $F_{D}=0$ for the complex connection, in order to introduce the universal Hitchin moduli space.
+In the next section we combine these equations with the flatness condition $F_{D}=0$ for the complex connection to introduce the universal Hitchin moduli space.
 4. Universal moduli spaces
@@
 Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$.
-Let $\curly D$} be the space of connection on $E_{G}$, equipped with the constant complex structure (2.1).
+Let $\curly D$} be the space of connections on $E_{G}$, equipped with the constant complex structure $\mathbf{J}$ given by (2.1).
 Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation.
 Consider the Kähler fibration $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ defined in Section 3.2.
-Our aim in this section is to investigate the existence of a Kähler structure on $\curly{X}$ and quotients thereof, combining the minimal coupling for the connection $\Gamma^{\mathbb{J}}$ constructed in Proposition 3.6, as considered in Section 3.3, with symplectic reduction.
-Difficulties will arise from Corollary 3.7, which shows that the associated symmetric tensor $\mathbf{g}_{\mathbb{J}}:=\sigma(,\mathbb{J})$ is negative semi-definite along horizontal directions.
+Our aim in this section is to investigate the existence of a Kähler structure on $\curly{X}$ and on its quotients, by combining the minimal coupling associated with the connection $\Gamma^{\mathbb{J}}$ constructed in Proposition 3.6, as discussed in Section 3.3, with symplectic reduction.
+Difficulties arise from Corollary 3.7, which shows that the associated symmetric tensor $\mathbf{g}_{\mathbb{J}}:=\sigma(,\mathbb{J})$ is negative semi-definite along the horizontal directions.
 In order to have some control in our construction, we need to impose the integrability condition $F_{D}=0$.
@@
 $\pi_{1}(g(J,D))=p(g)_{*}J.$
-In different words, the extension (4.1) is given by restriction of the natural map from holomorphic automorphisms of $\curly{X}$ which preserve the fibration structure, to holomorphic automorphisms of the base $\curly{J}$.
+In other words, the extension (4.1) is given by restriction of the natural map from holomorphic automorphisms of $\curly{X}$ which preserve the fibration structure, to holomorphic automorphisms of the base $\curly{J}$.
 It is important to observe that even though $\operatorname{Diff}_{0}(\Sigma)$ acts by holomorphic automorphisms of the base $\curly{J}$, this is a real infinite-dimensional Lie group, which does not admit a complexification (see e.g.
@@
 $\Omega_{\mathbb{J}}=\pi_{2}^{*}\Omega_{\mathbf{J}}.$
-Similarly as in the previous section, $\Omega_{\mathbb{J}}$ induces a structure of holomorphic symplectic fibration on $\curly{X}\to\curly{J}$ and a holomorphic Ehresmann connection, which in this case is simply the trivial connection.
+As in the previous section, $\Omega_{\mathbb{J}}$ induces a structure of holomorphic symplectic fibration on $\curly{X}\to\curly{J}$ and a holomorphic Ehresmann connection, which in this case is simply the trivial connection.
 We will return to this structure when studying the universal Higgs bundle moduli space in Section 5.
-The following result is an immediate consequence of the proof of [AGG, Proposition 1.6], but we sketch the proof in the present setup for the convenience of the reader.
+The following result is an immediate consequence of the proof of [AGG1, Proposition 1.6], but we sketch the proof in the present setup for the convenience of the reader.
 Proposition 4.1.
@@
 $\langle\mu_{\widetilde{\curly{G}}}(J,D),\zeta\rangle=\int_{\Sigma}B(F_{D},D\zeta),$
-where $D\zeta\in\Omega^{0}(X,E_{G}(\mathfrak{g}))$ denotes the vertical part of the $G$-equivariant vector field $\zeta$ on the total space of $E_{G}$, with respect to the connection $D$.
+where $D\zeta\in\Omega^{0}(\Sigma,E_{G}(\mathfrak{g}))$ denotes the vertical part of the $G$-equivariant vector field $\zeta$ on the total space of $E_{G}$, with respect to the connection $D$.
 Proof.
@@
 $=-\int_{\Sigma}B((d_{D}(D\zeta)+i_{p\zeta}F_{D})\wedge\dot{D}).$
-On the other hand, the infinitesimal action of $\zeta$ on $D$ is given by
-$\zeta\cdot D=-d_{D}(D\zeta)-i_{p\zeta}F_{D}$
-and the result follows. ∎
-It is remarkable that, unlike in the moment calculation in Proposition 3.9, in the present setup the extension of the complex gauge group by diffeomorphisms does not produce any additional ‘coupling term’ in the base $\Sigma$ (cf.
+The result now follows because the infinitesimal action of $\zeta$ on $D$ is given by
+$\zeta\cdot D=-d_{D}(D\zeta)-i_{p\zeta}F_{D}.\qed$
+It is remarkable that, unlike in the moment-map calculation in Proposition 3.9, in the present setup the extension of the complex gauge group by diffeomorphisms does not produce any additional ‘coupling term’ in the base $\Sigma$ (cf.
 Remark 4.2).
-One can intepret this fact as a functorial property of flat connections under the action of diffeomorphisms on a surface.
+One can interpret this fact as a functorial property of flat connections under the action of diffeomorphisms on a surface.
 This allows us to define the universal moduli space of flat $G$-connections, as follows, and to impose the flatness condition for $D$ as an integrability condition in the next sections.
@@
 Therefore, considering the pull-back of $\Omega_{\mathbb{J}}$ to $\curly{X}^{*}$, there is an induced moment map away from the singularities of $\curly{X}^{*}$, which, by abuse of notation, we denote simply by
-$\mu_{\widetilde{\curly{G}}|\curly{X}^{*}}\colon\curly{X}^{*}\to\operatorname{Lie}\widetilde{\curly{G}}^{*}$.
+$\mu_{\widetilde{\curly{G}}|\curly{X}^{*}}\colon\curly{X}^{*}\longrightarrow\operatorname{Lie}\widetilde{\curly{G}}^{*}$.
 From the previous discussion, it is natural to define the universal moduli space of flat $G$-connections on $E_{G}$ as the complex symplectic quotient
 $\mathcal{U}^{\operatorname{Flat}}(G):=\{(J,D)\;\in\;\curly{X}^{*}\;\;\mbox{with}\;\;F_{D}=0\}/\widetilde{\curly{G}}=\mu_{\widetilde{\curly{G}}|\curly{X}^{*}}^{-1}(0)/\widetilde{\curly{G}}$.
-By construction, $\mathcal{U}^{\operatorname{Flat}}(G)$ fibers over the Teichmüller space $\mathcal{T}$ of complex structures modulo diffeomorphisms isotopic to the identity
-(4.2) $\mathcal{U}^{\operatorname{Flat}}(G)\to\mathcal{T}:=\curly{J}/\operatorname{Diff}_{0}(\Sigma).$
+By construction, $\mathcal{U}^{\operatorname{Flat}}(G)$ fibres over the Teichmüller space $\mathcal{T}$ of complex structures modulo diffeomorphisms isotopic to the identity
+(4.2) $\mathcal{U}^{\operatorname{Flat}}(G)\longrightarrow\mathcal{T}:=\curly{J}/\operatorname{Diff}_{0}(\Sigma).$
 This fibration is naturally holomorphic, and $\Omega_{\mathbb{J}}$ induces a structure of holomorphic symplectic fibration with flat Ehresmann connection.
@@
 For a $G$-bundle on a higher dimensional symplectic manifold $(M^{2n},\omega)$, consider the subgroup $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$ given by the preimage by $p$ of the group of Hamiltonian symplectomorphisms on $M$.
-Then, the proof of [AGG, Proposition 1.6] shows that the induced $\widetilde{{\curly K}}-$action on $\curly D$} is Hamiltonian, with equivariant moment map
+Then the proof of [AGG1, Proposition 1.6] shows that the induced $\widetilde{{\curly K}}-$action on $\curly D$} is Hamiltonian, with equivariant moment map
 $\langle\boldsymbol{\mu}_{\widetilde{{\curly K}}}(J,D),\zeta\rangle=\frac{1}{(n-1)!}\int_{M}B(F_{D},D\zeta)\wedge\omega^{n-1}-\frac{1}{2(n-2)!}\int_{M}fB(F_{D}\wedge F_{D})\wedge\omega^{n-2},$
-where $f\in C^{\infty}(M,\mathbb{R})$ is the unique smooth function satisfying $i_{y}\omega=df$ and $\int_{M}f\omega^{n}=0$, for $p(\zeta)=y$.
-In this case, there is a ‘coupling term’ in the base corresponding to the function
+where $f\in C^{\infty}(M,\mathbb{R})$ is the unique smooth function satisfying $i_{y}\omega=df$ and $\int_{M}f\omega^{n}=0$, with $p(\zeta)=y$.
+In this case, there is a ‘coupling term’ on the base corresponding to the function
 $\frac{(n-1)}{2}\frac{B(F_{D}\wedge F_{D})\wedge\omega^{n-2}}{\omega^{n}}.$
 4.2. Universal moduli space of the coupled harmonic equations
-In this section we combine the results of the previous two sections in order to define universal moduli spaces of solutions of the harmonicity equations (2.14), varying over Teichmuller space.
+In this section we combine the results of the previous two sections in order to define universal moduli spaces of solutions of the harmonicity equations (2.11), varying over Teichmüller space.
 As in the previous sections, we consider a fixed principal $K$-bundle $E_{K}$ over a smooth compact oriented surface $\Sigma$ with fixed symplectic form $\omega$.
 Definition 4.3.
-We say that a triple $(J,A,\psi)\in\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ is a solution of the coupled harmonic equations with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, if the following conditions are satisfied
+A solution of the coupled harmonic equations with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$ is a triple $(J,A,\psi)\in\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$ that satisfies the equations
 (4.3) $\begin{split}F_{A}-\frac{1}{2}[\psi,\psi]&=0,\\ d_{A}\psi&=0,\\ d_{A}^{\ast}\psi&=0,\\ S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi\right)\right)&=\frac{2\pi\chi(\Sigma)}{V},\end{split}$
 where $g=\omega(,J)$ and * is the corresponding Hodge star operator.
-As in Section 4.1, consider ${\curly D}^{*}\subset{\curly D}$ the complex subspace of reductive flat connections.
-Then, $\curly{X}^{*}:=\curly{J}\times{\curly D}^{*}$ is formally a complex submanifold of $\curly{X}$ preserved by the $\widetilde{{\curly K}}-$action (see Section 3.3), and inherits a minimal coupling structure $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ and moment map $\mu_{\widetilde{{\curly K}}|\curly{X}^{*}}$.
+As in Section 4.1, we consider the complex subspace ${\curly D}^{*}\subset{\curly D}$ of reductive flat connections.
+Then $\curly{X}^{*}:=\curly{J}\times{\curly D}^{*}$ is formally a complex submanifold of $\curly{X}$ preserved by the $\widetilde{{\curly K}}-$action (see Section 3.3), and inherits a minimal coupling structure $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ and moment map $\mu_{\widetilde{{\curly K}}|\curly{X}^{*}}$.
 We define the universal Hitchin moduli space as the symplectic quotient
-$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}:=\mu_{\widetilde{{\curly K}}|\curly{X}^{*}}^{-1}(0)/\widetilde{{\curly K}}$,
-By construction, $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ parametrizes solutions of the coupled harmonic equations (4.3) with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, modulo the $\widetilde{{\curly K}}-$action, and inherits a natural pre-symplectic structure on its smooth locus.
+$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}:=\mu_{\widetilde{{\curly K}}|\curly{X}^{*}}^{-1}(0)/\widetilde{{\curly K}}$.
+By construction, $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ parametrizes solutions of the coupled harmonic equations (4.3) with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, modulo the $\widetilde{{\curly K}}-$action, and inherits a natural presymplectic structure on its smooth locus.
 This structure is furthermore symplectic in the case $\varepsilon=-1$, by Lemma 3.8.
@@
 (5.3))
-$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)\to\mathcal{T}.$
+$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\longrightarrow\mathcal{U}^{\operatorname{Flat}}(G)\longrightarrow\mathcal{T}.$
 Hence, in particular, $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ can be regarded as a fibration over the Teichmüller space $\mathcal{T}$.
@@
 An infinitesimal variation of the triple $(J,A,\psi)$ is given by
-(4.4) $(\dot{J},a,\dot{\psi})\in\mathcal{S}^{1}:=T_{J}\curly{J}\oplus\Omega^{1}(X,E_{K}(\mathfrak{k}))\oplus\Omega^{1}(X,E_{K}(\mathfrak{k})).$
+(4.4) $(\dot{J},a,\dot{\psi})\in\mathcal{S}^{1}:=T_{J}\curly{J}\oplus\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))\oplus\Omega^{1}(\Sigma,E_{K}(\mathfrak{k})).$
 As above, we identify $T_{J}\curly{J}$ with the space of endomorphisms $\dot{J}\colon T\Sigma\to T\Sigma$ such that $\dot{J}J=-J\dot{J}$.
@@
 We denote by ${\bf L}_{\alpha}^{\varepsilon}(\dot{J},a,\dot{\psi})$ the differential operator defined by the left-hand side of equations (4.5).
-We turn next to the study of the infinitesimal action, in order to define a complex.
+We turn next to the study of the infinitesimal action, in order to define a complex of differential operators.
 From the proof of Proposition 3.9, using the connection $A$ we can identify elements $\zeta\in\operatorname{Lie}\widetilde{{\curly K}}$ with pairs
-$\zeta\cong(f,u)\in\mathcal{S}^{0}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(X,E_{K}(\mathfrak{k})),$
+$\zeta\cong(f,u)\in\mathcal{S}^{0}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(\Sigma,E_{K}(\mathfrak{k})),$
 and the infinitesimal action $\mathbf{P}(f,u):=(f,u)\cdot(J,A,\psi)$ at $(J,A,\psi)$ is
@@
 Define the vector space
-$\mathcal{S}^{2}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus(\Omega^{2}(X,E_{K}(\mathfrak{k}))^{\oplus 3},$
+$\mathcal{S}^{2}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus(\Omega^{2}(\Sigma,E_{K}(\mathfrak{k})))^{\oplus 3},$
 so that ${\bf L}_{\alpha}^{\varepsilon}(\dot{J},a,\dot{\psi})\in\mathcal{S}^{2}$ (where we read equations (4.5) from bottom to top), and consider the complex of linear differential operators
-(4.7) ${(\mathcal{S}^{*})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{S}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}}$${\mathcal{S}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L}_{\alpha}^{\varepsilon}}$${\mathcal{S}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
-The cohomology $H^{1}(\mathcal{S}^{*}):=\frac{\operatorname{Ker}{\bf L}_{\alpha}^{\varepsilon}}{\operatorname{Im}\,\mathbf{P}}$ can be formally identified with the tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$.
+[…]
+The cohomology $H^{1}(\mathcal{S}^{*}):=\frac{\ker{\bf L}_{\alpha}^{\varepsilon}}{\operatorname{Im}\,\mathbf{P}}$ can be formally identified with the tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$.
 Our next result shows that the moduli space $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ is finite dimensional.
-Notice that both $\mathbf{P}$ and ${\bf L}_{\alpha}^{\varepsilon}$ are multi-degree differential operators, and hence we shall use the use the generalized notion of ellipticity provided by Douglis and Nirenberg [DN].
+Notice that both $\mathbf{P}$ and ${\bf L}_{\alpha}^{\varepsilon}$ are multi-degree differential operators, and hence we shall use the generalized notion of ellipticity provided by Douglis and Nirenberg [DN].
 For the general theory of linear multi-degree elliptic differential operators we refer to [LM].
@@
 where to clarify the notation, we notice e.g. that
-$\mathbf{P}_{11}(f)$ $=-L_{\eta_{f}}J,$ $\mathbf{P}_{12}$ $=0,$ $\mathbf{P}_{21}(f)$ $=-i_{\eta_{f}}F_{A}$
+$\mathbf{P}_{11}(f)$ $=-L_{\eta_{f}}J,$ $\mathbf{P}_{12}$ $=0,$ $\mathbf{P}_{21}(f)$ $=-i_{\eta_{f}}F_{A},$
 $\mathbf{P}_{22}(u)$ $=-d_{A}u,$ $\mathbf{P}_{31}(f)$ $=-d_{A}(i_{\eta_{f}}\psi),$ $\mathbf{P}_{32}(u)$ $=-[\psi,u].$
 Then, the tuples $\mathbf{t}=(2,2)$ and $\mathbf{s}=(0,1,0)$ form a system of orders for $\mathbf{P}$, since
-$o(\mathbf{P}_{11})$ $=2\leqslant 2-0,$ $o(\mathbf{P}_{12})$ $=0\leqslant 2-0,$ $o(\mathbf{P}_{21})$ $=1\leqslant 2-1$
+$o(\mathbf{P}_{11})$ $=2\leqslant 2-0,$ $o(\mathbf{P}_{12})$ $=0\leqslant 2-0,$ $o(\mathbf{P}_{21})$ $=1\leqslant 2-1,$
 $o(\mathbf{P}_{22})$ $=1\leqslant 2-1,$ $o(\mathbf{P}_{31})$ $=2\leqslant 2-0,$ $o(\mathbf{P}_{32})$ $=0\leqslant 2-0.$
@@
 Similarly, $\mathbf{t}=(2,2,2)$ and $\mathbf{s}=(0,1,1,1)$ form a system of orders for ${\bf L}_{\alpha}^{\varepsilon}$, since
-$o(\mathbf{L}_{11})$ $=2\leqslant 2-0,$ $o(\mathbf{L}_{12})$ $=2\leqslant 2-0,$ $o(\mathbf{L}_{13})$ $=1\leqslant 2-0$
+$o(\mathbf{L}_{11})$ $=2\leqslant 2-0,$ $o(\mathbf{L}_{12})$ $=2\leqslant 2-0,$ $o(\mathbf{L}_{13})$ $=1\leqslant 2-0,$
 $o(\mathbf{L}_{21})$ $=1\leqslant 2-1,$ $o(\mathbf{L}_{22})$ $=0\leqslant 2-1,$ $o(\mathbf{L}_{23})$ $=1\leqslant 2-1,$
@@
 By dimensional reasons $v\wedge\omega=0$, and hence
-$u_{4}\omega=v\wedge a,\qquad\textrm{ for }a\in\Omega^{1}(X,E_{K}(\mathfrak{k}).$
+$u_{4}\omega=v\wedge a,\qquad\textrm{ for }a\in\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}).$
 We can choose $\dot{J}$ in the line spanned by $g^{-1}v\otimes Jv$ such that
@@
 To finish, assume that $\sigma_{{\bf L}_{\alpha}^{\varepsilon}}(v)(\dot{J},a,\dot{\psi})=0$.
-Then, $\sigma_{\mathbf{L}_{4}}(v)=0$ implies that there exists $x\in\Omega^{0}(X,E_{K}(\mathfrak{k})$ such that
+Then, $\sigma_{\mathbf{L}_{4}}(v)=0$ implies that there exists $x\in\Omega^{0}(\Sigma,E_{K}(\mathfrak{k})$ such that
 $a=v\otimes x.$
@@
 $B(\Lambda_{\omega}(v\wedge a),*\psi)=0$
-and therefore $\sigma_{\mathbf{L}_{1}}(v)=0$ implies $v\wedge v(\dot{J})=0$. By [F], there exists a smooth function $f\in C^{\infty}_{0}(\Sigma,\mathbb{R})$ such that
-$\dot{J}=-fJv\otimes\omega^{-1}v$
+and therefore $\sigma_{\mathbf{L}_{1}}(v)=0$ implies $v\wedge v(\dot{J})=0$. By [Fu], there exists a smooth function $f\in C^{\infty}_{0}(\Sigma,\mathbb{R})$ such that
+$\dot{J}=-fJv\otimes\omega^{-1}v.$
 Combined with $\sigma_{\mathbf{L}_{2}}(v)=0$, we obtain
@@
 $\dot{\psi}=fv\otimes\psi(Jg^{-1}v).$
-Finally, writting $\omega=\lambda v\wedge Jv$ and
-$F_{A}=tv\wedge Jv\textrm{ for }t\in\Omega^{0}(X,E_{K}(\mathfrak{k}),$
+Finally, writing $\omega=\lambda v\wedge Jv$ and
+$F_{A}=tv\wedge Jv\textrm{ for }t\in\Omega^{0}(\Sigma,E_{K}(\mathfrak{k}),$
 it follows that
@@
 and therefore we conclude
-$(\dot{J},a,\dot{\psi})=\sigma_{{\bf P}}(v)(f,-x-\lambda^{-1}ft).$
+$(\dot{J},a,\dot{\psi})=\sigma_{{\bf P}}(v)(f,-x-\lambda^{-1}ft).\qed$
 Our strategy to build a complex structure induced by (3.7) on the moduli space is to work orthogonally to the image of the infinitesimal action operator $\mathbf{P}$ in (4.7) with respect to the indefinite pairing $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ in (3.18).
-The existence of this complex structure will automatically yield a symmetric tensor of type $(1,1)$, since the two-form $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ in (3.17) is well defined on the cohomology $H^{1}(\mathcal{S}^{*})$ by Proposition 3.9.
-Consider the $L^{2}$-pairing on $\mathcal{S}^{0}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(X,E_{K}(\mathfrak{k})$, the domain of the operator $\mathbf{P}$, induced by $\omega$ and $B$:
+The existence of this complex structure will automatically yield a symmetric tensor of type $(1,1)$, since the 2-form $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ in (3.17) is well defined on the cohomology $H^{1}(\mathcal{S}^{*})$ by Proposition 3.9.
+Consider the $L^{2}$-pairing on $\mathcal{S}^{0}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(\Sigma,E_{K}(\mathfrak{k})$, the domain of the operator $\mathbf{P}$, induced by $\omega$ and $B$:
 (4.8) $\begin{split}\langle(f,u),(f,u)\rangle&{}=\int_{\Sigma}f^{2}\omega+\int_{\Sigma}B(u,u)\omega.\end{split}$
@@
 The following operator provides a formal adjoint of $\widetilde{\mathbf{P}}$ for the pairings (4.8) and (3.18)
-$\widetilde{\mathbf{P}}^{*}=\delta\mu\circ\mathbb{J}\colon\mathcal{S}^{1}\to\mathcal{S}^{0}.$
+$\widetilde{\mathbf{P}}^{*}=\delta\mu\circ\mathbb{J}\colon\mathcal{S}^{1}\longrightarrow\mathcal{S}^{0}.$
 More explicitly, setting $\widetilde{\mathbf{P}}^{*}=\widetilde{\mathbf{P}}^{*}_{1}\oplus\widetilde{\mathbf{P}}^{*}_{2}$, we have:
@@
 Then, setting $v=(\dot{J},a,\dot{\psi})$, we have
-$\begin{split}\langle\widetilde{\mathbf{P}}^{*}v,(f,u)\rangle=\langle d\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}(\mathbb{J}v),\zeta\rangle=\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(\zeta\cdot(J,A,\psi),\mathbb{J}v)=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(\widetilde{\mathbf{P}}(f,u),v)=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,\widetilde{\mathbf{P}}(f,u)).\end{split}$
+$\langle\widetilde{\mathbf{P}}^{*}v,(f,u)\rangle$ $=\langle d\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}(\mathbb{J}v),\zeta\rangle=\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(\zeta\cdot(J,A,\psi),\mathbb{J}v)$
+$=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(\widetilde{\mathbf{P}}(f,u),v)=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,\widetilde{\mathbf{P}}(f,u)).\qed$
 Consider now the differential operator
-[…]
+(4.11) $\begin{array}[]{cccl}\mathcal{L}\colon&\mathcal{S}^{0}&\longrightarrow&\mathcal{S}^{0}\\ &(f,u)&\longmapsto&\widetilde{\mathbf{P}}^{*}\circ\widetilde{\mathbf{P}}(f,u).\end{array}$
 The key condition on the solution $(J,A,\psi)$ of (4.3) which we need to assume in order to construct the complex structure on the moduli space is the vanishing of the kernel of $\mathcal{L}$.
-Notice that, unlike in the standard cases in gauge theory in which the parameter space metric is positive definite, $\operatorname{ker}\mathcal{L}$ does not relate in general to automorphisms of the triple $(J,\psi,A)$, but rather to null vectors with respect to $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$.
+Notice that, unlike in the standard cases in gauge theory in which the parameter space metric is positive definite, $\ker\mathcal{L}$ does not relate in general to automorphisms of the triple $(J,\psi,A)$, but rather to null vectors with respect to $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$.
 We build on the following technical result, which shows in particular that the subspace of null vectors in $\operatorname{Im}\mathbf{P}$ is finite-dimensional.
@@
 Proposition 4.7.
-The operator $\mathcal{L}$ is Fredholm with zero index. Furthermore, elements $(f,u)\in\operatorname{ker}\mathcal{L}$ are smooth.
+The operator $\mathcal{L}$ is Fredholm with zero index. Furthermore, elements $(f,u)\in\ker\mathcal{L}$ are smooth.
 Proof.
@@
 Then, the tuples $\mathbf{t}=(4,4)$ and $\mathbf{s}=(0,2)$ form a system of orders for $\widetilde{\mathcal{L}}$, since
-$o(\widetilde{\mathcal{L}}_{11})$ $=4\leqslant 4-0,$ $o(\widetilde{\mathcal{L}}_{12})$ $=2\leqslant 4-0,$
-$o(\widetilde{\mathcal{L}}_{21})$ $=2\leqslant 4-2,$ $o(\widetilde{\mathcal{L}}_{22})$ $=2\leqslant 4-2.$
+$o(\widetilde{\mathcal{L}}_{11})$ $=4\leqslant 4-0,\quad$ $\quad o(\widetilde{\mathcal{L}}_{12})$ $=2\leqslant 4-0,$
+$o(\widetilde{\mathcal{L}}_{21})$ $=2\leqslant 4-2,$ $o(\widetilde{\mathcal{L}}_{22})$ $=2\leqslant 4-2,$
 and the associated leading symbol is
@@
 which is clearly invertible.
-Consequently, by the general theory of linear multi-degree elliptic differential operators (see [LM]) it follows that $\widetilde{\mathcal{L}}$ is Fredholm and elements in $\operatorname{ker}\widetilde{\mathcal{L}}$ are smooth.
-The result follows now from $\operatorname{ker}\mathcal{L}=\mathbf{T}^{-1}\operatorname{ker}\widetilde{\mathcal{L}}$ and the fact that $\mathcal{L}$ is self-adjoint.
+Consequently, by the general theory of linear multi-degree elliptic differential operators (see [LM]) it follows that $\widetilde{\mathcal{L}}$ is Fredholm and elements in $\ker\widetilde{\mathcal{L}}$ are smooth.
+The result follows now from $\ker\mathcal{L}=\mathbf{T}^{-1}\ker\widetilde{\mathcal{L}}$ and the fact that $\mathcal{L}$ is self-adjoint.
 ∎
-Assuming that $\operatorname{ker}\mathcal{L}$ is trivial, in the next result we obtain a natural gauge fixing via a $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$-orthogonal decomposition
+Assuming that $\ker\mathcal{L}$ is trivial, in the next result we obtain a natural gauge fixing via a $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$-orthogonal decomposition
 (4.13) $\mathcal{S}^{1}=\operatorname{Im}\;\widetilde{\mathbf{P}}\oplus(\operatorname{Im}\;\widetilde{\mathbf{P}})^{\perp_{\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}}}.$
 Lemma 4.8.
-Assume that $\operatorname{ker}\mathcal{L}=\{0\}$.
+Assume that $\ker\mathcal{L}=\{0\}$.
 Then, there exists an orthogonal decomposition (5.20) for the pairing $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$.
 Consequently, for any element $v\in\mathcal{S}^{1}$ there exists a unique $\Pi v\in\operatorname{Im}\;\widetilde{\mathbf{P}}$ such that $(\dot{J},a,\dot{\psi})=v-\Pi v$ solves the linear equations
-$\begin{split}\alpha d_{A}(Ja)-\alpha[\dot{\psi},J\psi]-\alpha d_{A}(\psi(J\dot{J}))&=0\\ \varepsilon\delta S(J\dot{J})+\alpha*d\left(B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)-B(\Lambda_{\omega}F_{A},*a)+B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right)&=0.\end{split}$
+$\begin{split}\alpha d_{A}(Ja)-\alpha[\dot{\psi},J\psi]-\alpha d_{A}(\psi(J\dot{J}))&=0,\\ \varepsilon\delta S(J\dot{J})+\alpha*d\left(B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)-B(\Lambda_{\omega}F_{A},*a)+B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right)&=0.\end{split}$
 Proof.
@@
 If $v\in\operatorname{Im}\;\widetilde{\mathbf{P}}\cap(\operatorname{Im}\;\widetilde{\mathbf{P}})^{\perp_{\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}}}$, then $v=\widetilde{\mathbf{P}}(y)$ for $y\in\mathcal{S}^{0}$.
-But then $\widetilde{\mathbf{P}}^{*}\circ\widetilde{\mathbf{P}}(y)=0$ and, by $\operatorname{ker}\mathcal{L}=\{0\}$, $v=0$.
+But then $\widetilde{\mathbf{P}}^{*}\circ\widetilde{\mathbf{P}}(y)=0$ and, by $\ker\mathcal{L}=\{0\}$, $v=0$.
 Thus
@@
 (4.15) ${\bf\widetilde{P}}^{*}(v)={\bf{\widetilde{P}}}^{*}\circ\widetilde{\mathbf{P}}(y).$
-But by Proposition 4.7 and condition $\operatorname{ker}\mathcal{L}=\{0\}$, $\widetilde{\mathbf{P}}^{*}\circ{\bf\widetilde{P}}$ is surjective.
+But by Proposition 4.7 and condition $\ker\mathcal{L}=\{0\}$, $\widetilde{\mathbf{P}}^{*}\circ{\bf\widetilde{P}}$ is surjective.
 Then, by elliptic regularity, one can solve (4.15) for $y\in\mathcal{S}^{0}$.
@@
 Proposition 4.9.
-Assume $\operatorname{ker}\mathcal{L}=\{0\}$ and $\alpha\neq 0$. Then, the inclusion $\mathcal{H}^{1}(\mathcal{S}^{*})\subset\ker\bf{L}^{\varepsilon}_{\alpha}$ induces an isomorphism
-$\mathcal{H}^{1}(\mathcal{S}^{*})\simeq H^{1}(\mathcal{S}^{*}).$
+Assume $\ker\mathcal{L}=\{0\}$ and $\alpha\neq 0$. Then, the inclusion $\mathcal{H}^{1}(\mathcal{S}^{*})\subset\ker\bf{L}^{\varepsilon}_{\alpha}$ induces an isomorphism
+$\mathcal{H}^{1}(\mathcal{S}^{*})\cong H^{1}(\mathcal{S}^{*}).$
 More precisely, any class in the cohomology $H^{1}(\mathcal{S}^{*})$ of the complex (4.7) admits a unique representative $(\dot{J},a,\dot{\psi})$ solving the linear equations
-(4.16) $\begin{split}d_{A}a-[\dot{\psi},\psi]&=0,\\ d_{A}\dot{\psi}+[a,\psi]&=0,\\ d_{A}(J\dot{\psi})+[a,J\psi]-d_{A}(\psi(\dot{J}))&=0,\\ d_{A}(Ja)-[\dot{\psi},J\psi]-d_{A}(\psi(J\dot{J}))&=0\\ \varepsilon\delta S(\dot{J})-\alpha*d\left(B(\Lambda_{\omega}d_{A}a,*\psi)+B(\Lambda_{\omega}F_{A},*\dot{\psi})-B(\Lambda_{\omega}F_{A},\psi(\dot{J}))\right)&=0,\\ \varepsilon\delta S(J\dot{J})-\alpha*d\left(-B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)+B(\Lambda_{\omega}F_{A},*a)-B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right)&=0,\end{split}$
+(4.16) $\begin{split}d_{A}a-[\dot{\psi},\psi]&=0,\\ d_{A}\dot{\psi}+[a,\psi]&=0,\\ d_{A}(J\dot{\psi})+[a,J\psi]-d_{A}(\psi(\dot{J}))&=0,\\ d_{A}(Ja)-[\dot{\psi},J\psi]-d_{A}(\psi(J\dot{J}))&=0\\ \varepsilon\delta S(\dot{J})-\alpha*d\left(B(\Lambda_{\omega}d_{A}a,*\psi)+B(\Lambda_{\omega}F_{A},*\dot{\psi})-B(\Lambda_{\omega}F_{A},\psi(\dot{J}))\right)&=0,\\ \varepsilon\delta S(J\dot{J})-\alpha*d\left(-B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)+B(\Lambda_{\omega}F_{A},*a)-B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right)&=0.\end{split}$
 Proof.
@@
 Define
-$\mathcal{U}^{*}_{\alpha,\varepsilon}=\{[(J,A,\psi)]\;|\;\operatorname{ker}\mathcal{L}=\{0\}\}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}.$
+$\mathcal{U}^{*}_{\alpha,\varepsilon}=\{[(J,A,\psi)]\;|\;\ker\mathcal{L}=\{0\}\}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}.$
 Theorem 4.10.
@@
 In either case, $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ admits a global Kähler potential, that is, $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}=dd^{c}_{\mathbb{J}}\Phi$, where
-$\Phi=\varepsilon\nu{J}+\frac{\alpha}{2}\|\psi\|^{2}_{L^{2}}$
-and $\nu{J}$ is induced by the global Kähler potential in the space of complex structures compatible with the orientation $\curly{J}$ (see [F, Section 4]).
+$\Phi=\varepsilon\nu{J}+\frac{\alpha}{2}\|\psi\|^{2}_{L^{2}},$
+and $\nu{J}$ is induced by the global Kähler potential in the space of complex structures compatible with the orientation $\curly{J}$ (see [Fu, Section 4]).
 Proof.
@@
 Given now a smooth point $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$, the tangent space is identified with $H^{1}(\mathcal{S}^{*})$ and inherits a complex structure by Proposition 4.9, using that $\mathbb{J}$ in (3.7) preserves (4.16).
-The existence the of the symmetric tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is a direct consequence of Lemma 3.4 and Proposition 3.9, while the listed signature properties follow from Lemma 3.8.
-The formula for the Kähler potential is a direct consequence of (3.13), while a explicit formula for $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ follows from Lemma 3.4.
+The existence of the symmetric tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is a direct consequence of Lemma 3.4 and Proposition 3.9, while the listed signature properties follow from Lemma 3.8.
+The formula for the Kähler potential is a direct consequence of (3.9), while an explicit formula for $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ follows from Lemma 3.4.
 ∎
@@
 Then, one has
-(4.17) $\begin{split}\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2}))\\ &+\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge J(\psi(\dot{J}_{2}))),\end{split}$
+(4.17) $\begin{split}\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &+\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge J(\psi(\dot{J}_{2}))).\end{split}$
 4.3. Comparison with $\boldsymbol{\mathcal{U}^{\operatorname{Flat}}(G)}$ and existence
-This aim of this section is twofold.
+The aim of this section is twofold.
 Firstly, we establish a comparison between the moduli spaces $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ and the universal moduli space of flat $G$-connections $\mathcal{U}^{\operatorname{Flat}}(G)$ constructed in Section 4.1.
 As we will see, for any $\alpha>0$ and $\varepsilon\neq 0$, Theorem 4.10 induces a natural holomorphic map
-$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)$
-and hence a holomorphic map into Teichmuller space $\mathcal{T}$.
+$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\longrightarrow\mathcal{U}^{\operatorname{Flat}}(G),$
+and hence a holomorphic map into Teichmüller space $\mathcal{T}$.
 Secondly, we will prove that for genus of the surface $\Sigma$ bigger than zero, the open $\mathcal{U}^{*}_{\alpha,\varepsilon}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ is non-empty for sufficiently small values of $\alpha$.
@@
 (4.18) $\begin{split}F_{A}-\frac{1}{2}[\psi,\psi]&=0,\\ d_{A}\psi&=0.\end{split}$
-The tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{\operatorname{Flat}}(G)$ can be formally identified with the cohomology of the complex of linear differential operators
-(4.19) ${(\mathcal{C}^{*})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{C}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}^{c}}$${\mathcal{C}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L^{c}}}$${\mathcal{C}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
+The tangent space $T_{\left[(J,A,\psi)\right]}\mathcal{U}^{\operatorname{Flat}}(G)$ can be formally identified with the cohomology of the complex of linear differential operators
+[…]
 where we have
-$\mathcal{C}^{0}=\operatorname{Lie}\widetilde{\curly{G}}\cong\Omega^{0}(T\Sigma)\oplus\Omega^{0}(X,E_{G}(\mathfrak{g})),\qquad\mathcal{C}^{1}=\mathcal{S}^{1},\qquad\mathcal{C}^{2}=\Omega^{2}(X,E_{G}(\mathfrak{g}))$
+$\mathcal{C}^{0}=\operatorname{Lie}\widetilde{\curly{G}}\cong\Omega^{0}(T\Sigma)\oplus\Omega^{0}(\Sigma,E_{G}(\mathfrak{g})),\qquad\mathcal{C}^{1}=\mathcal{S}^{1},\qquad\mathcal{C}^{2}=\Omega^{2}(\Sigma,E_{G}(\mathfrak{g})),$
 and
@@
 Again, by dimensional reasons $\dot{J}=-Jv\otimes y$ for some vector $y$, and hence
-$(\dot{J},a,\dot{\psi})=\sigma_{\mathbf{P}^{c}}(v)(y,u_{0}+i(u_{1}^{\prime}-\psi(y))).$
+$(\dot{J},a,\dot{\psi})=\sigma_{\mathbf{P}^{c}}(v)(y,u_{0}+i(u_{1}^{\prime}-\psi(y))).\qed$
 Applying Theorem 2.5, a solution $(J,A,\psi)$ of the coupled harmonic equations (4.3) induces a reductive flat $G$-connection $D=A+i\psi\in{\curly D}^{*}$.
 This fact, jointly with the natural inclusion $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$, leads to a continuous map
-(4.20) $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)\colon[(J,A,\psi)]\mapsto[(J,A+i\Psi)].$
+(4.20) $\begin{array}[]{rcl}\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}&\longrightarrow&\mathcal{U}^{\operatorname{Flat}}(G)\\ \left[(J,A,\psi)\right]&\longmapsto&\left[(J,A+i\psi)\right].\end{array}$
 Building on Theorem 4.10, our next goal is to prove that this induces a holomorphic map $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)$.
@@
 Let $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$. Then, (4.20) induces a complex linear map
-(4.21) ${H^{1}(\mathcal{S}^{*})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathcal{C}^{*}).}$
+(4.21) $H^{1}(\mathcal{S}^{*})$$H^{1}(\mathcal{C}^{*}),$}
 where the complex structure on $H^{1}(\mathcal{S}^{*})$ is the one induced by Proposition 4.9.
@@
 Using that $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$ we can identify
-$H^{1}(\mathcal{S}^{*})\simeq\mathcal{H}^{1}(\mathcal{S}^{*}),$
-where the right hand side is given by solutions of the gauge-fixed system of linear equations (4.16). Then, the map
-$\mathcal{H}^{1}(\mathcal{S}^{*})\to H^{1}(\mathcal{C}^{*})\colon[(\dot{J},a,\dot{\psi})]\mapsto[(\dot{J},a+i\dot{\psi})]$
+$H^{1}(\mathcal{S}^{*})\cong\mathcal{H}^{1}(\mathcal{S}^{*}),$
+where the right-hand side is given by solutions of the gauge-fixed system of linear equations (4.16). Then, the map
+$\begin{array}[]{ccl}\mathcal{H}^{1}(\mathcal{S}^{*})&\longrightarrow&H^{1}(\mathcal{C}^{*})\\ \left[(\dot{J},a,\dot{\psi})\right]&\longmapsto&\left[(\dot{J},a+i\dot{\psi})\right]\end{array}$
 is complex $\mathbb{C}$-linear, since both complex structures are induced by $\mathbb{I}$ in (3.7). ∎
@@
 In this setup, consider the coupled equations
-(4.22) $\begin{split}d_{A_{h}}^{*}\psi_{h}&=0\\ \varepsilon S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi_{h}\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V}\end{split}$
+(4.22) $\begin{split}d_{A_{h}}^{*}\psi_{h}&=0,\\ \varepsilon S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi_{h}\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
 for pairs $(g,h)$, where $g$ is a Kähler metric on $X$ with total volume $V$ and $h\in\Omega^{0}(E_{G}(G/K))$ is a reduction of structure group of $E_{G}$ to $K$. Here, we denote
@@
 Consider the non-linear operator
-$\mathbf{Q}_{\alpha}\colon C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(X,E_{K}(\mathfrak{k}))\to C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{2}(X,E_{K}(\mathfrak{k}))$
+$\mathbf{Q}_{\alpha}\colon C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(\Sigma,E_{K}(\mathfrak{k}))\longrightarrow C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{2}(\Sigma,E_{K}(\mathfrak{k}))$
 defined by
-$\mathbf{Q}_{\alpha}(f,u)=(S_{g_{f}}-\varepsilon^{-1}\alpha*d\left(B\left(\Lambda_{\omega_{f}}F_{A_{h_{u}}},*\psi_{h_{u}}\right)\right)-2\pi\chi(\Sigma)V^{-1},\operatorname{Ad}(e^{-iu})d_{A_{h_{u}}}J\psi_{h_{u}})$
+$\mathbf{Q}_{\alpha}(f,u)=(S_{g_{f}}-\varepsilon^{-1}\alpha*d\left(B\left(\Lambda_{\omega_{f}}F_{A_{h_{u}}},*\psi_{h_{u}}\right)\right)-2\pi\chi(\Sigma)V^{-1},\operatorname{Ad}(e^{-iu})d_{A_{h_{u}}}J\psi_{h_{u}}),$
 where
@@
 Following the notation in Lemma 5.14, for $(g_{\alpha},A_{h_{\alpha}},\psi_{h_{\alpha}})$ define the one-parameter family of multidegree differential operators
-$\mathcal{N}^{\alpha}=(\varepsilon^{-1}\widetilde{\mathbf{P}}_{1}^{*},\alpha^{-1}\widetilde{\mathbf{P}}_{2}^{*})\circ\widetilde{\mathbf{P}}\colon\mathcal{S}^{0}\to\mathcal{S}^{0}.$
+$\mathcal{N}^{\alpha}=(\varepsilon^{-1}\widetilde{\mathbf{P}}_{1}^{*},\alpha^{-1}\widetilde{\mathbf{P}}_{2}^{*})\circ\widetilde{\mathbf{P}}\colon\mathcal{S}^{0}\longrightarrow\mathcal{S}^{0}.$
 Notice that $\mathcal{N}^{0}$ is well-defined, and furthermore
@@
 Decomposing
-$\mathcal{N}^{0}=\mathcal{N}^{0}_{0}\oplus\mathcal{N}^{0}_{1}$
+$\mathcal{N}^{0}=\mathcal{N}^{0}_{0}\oplus\mathcal{N}^{0}_{1},$
 it follows that
@@
 $\ker\mathcal{N}^{\alpha}=\{0\}.$
-The last part of the statement follows from that fact that (5.25) is induced by a $C^{\infty}$-Frechet map, combined with Lemma 5.21.
+The last part of the statement follows from that fact that (4.20) is induced by a $C^{\infty}$-Frechet map, combined with Lemma 5.21. ∎
 Remark 4.15.
@@
 We also consider a connected semisimple complex Lie group $G$, with Lie algebra $\mathfrak{g}$, and fix an antiholomorphic involution $\tau$ of $G$ defining a maximal compact subgroup $K:=G^{\tau}\subset G$, with Lie algebra $\mathfrak{k}$.
-Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$.
-Consider $h\in\Omega^{0}(E_{G}(G/K))$ a reduction of structure group of $E_{G}$ to $K$, and let $E_{K}$ be the corresponding principal $K$-bundle.
+Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$, $h\in\Omega^{0}(E_{G}(G/K))$ a reduction of the structure group of $E_{G}$ to $K$, and $E_{K}$ the corresponding principal $K$-bundle.
 Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation.
 Consider the space
-$\curly{X}=\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k})).$
+$\curly{X}=\curly{J}\times{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k})).$
 In this section we are interested in the geometry of the submanifold of $\curly{X}$ given by universal Higgs fields,
-$\curly{X}^{Higgs}=\{(J,A,\psi)\in\curly{X}\;|\;\bar{\partial}_{J,A}\varphi=0\}$
+$\curly{X}^{Higgs}=\{(J,A,\psi)\in\curly{X}\;|\;\bar{\partial}_{J,A}\varphi=0\},$
 where the Higgs field
@@
 In particular, the map $\pi_{1}\colon\curly{X}\to\curly{J}$ is holomorphic, and the complex structure along the fibres coincides with the $J$-dependent complex structure $\mathbf{I}$ in (2.7).
-Using the Chern-Singer correspondence, it is not difficult to give holomorphic coordinates on the space $\curly{J}\times{\curly A}\cong\curly{J}\times{\curly C}$, and hence on $(\curly{X},\mathbb{I})$.
+Using the Chern–Singer correspondence, it is not difficult to give holomorphic coordinates on the space $\curly{J}\times{\curly A}\cong\curly{J}\times{\curly C}$, and hence on $(\curly{X},\mathbb{I})$.
 In the next result we give an independent proof of the vanishing of the Nijenhuis tensor $N_{\mathbb{I}}$ of $\mathbb{I}$.
 Lemma 5.1.
-The complex structure $\mathbb{I}$ is formally integrable, that is, its Nijenhuis tensor vanishes
+The complex structure $\mathbb{I}$ is formally integrable, that is, its Nijenhuis tensor vanishes:
 $N_{\mathbb{I}}=0.$
 Proof.
-The complex structures on the fibers and base of the fibration $(\curly{X},\mathbb{I})\to(\curly{J},\mathbb{J}{J})$ are well-known to be integrable.
+The complex structures on the fibres and base of the fibration $(\curly{X},\mathbb{I})\to(\curly{J},\mathbb{J}{J})$ are well-known to be integrable.
 Hence, it suffices to calculate the mixed component of the Nijenhuis tensor
@@
 $=(0,-a(J\dot{J}),\dot{\psi}(J\dot{J}))-\mathbb{I}(0,-a(\dot{J}),\dot{\psi}(\dot{J}))$
-$=(0,-a(J\dot{J}+\dot{J}J),\dot{\psi}(J\dot{J}+\dot{J}J))=0.$
-In our next result we show that $\curly{X}^{Higgs}\subset(\curly{X},\mathbb{I})$ is fomally a complex submanifold.
+$=(0,-a(J\dot{J}+\dot{J}J),\dot{\psi}(J\dot{J}+\dot{J}J))=0.\qed$
+In our next result we show that $\curly{X}^{Higgs}\subset(\curly{X},\mathbb{I})$ is formally a complex submanifold.
 Proposition 5.2.
-The tangent space of $\curly{X}^{Higgs}$ at $(J,A,\psi)$, is given by triples $(\dot{J},a,\dot{\psi})\in T_{J}\curly{J}\oplus\Omega^{1}(X,E_{K}(\mathfrak{k}))\oplus\Omega^{1}(X,E_{K}(\mathfrak{k}))$ satisfying
+The tangent space of $\curly{X}^{Higgs}$ at $(J,A,\psi)$, is given by triples $(\dot{J},a,\dot{\psi})\in T_{J}\curly{J}\oplus\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))\oplus\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$ satisfying
 (5.2) $d_{A}\dot{\psi}^{1,0_{J}}+[a,\varphi]-\frac{i}{2}d_{A}(\psi(\dot{J}))=0.$
-Consequently, $\curly{X}^{Higgs}$ is fomally a complex submanifold of $(\curly{X},\mathbb{I})$.
+Consequently, $\curly{X}^{Higgs}$ is formally a complex submanifold of $(\curly{X},\mathbb{I})$.
 Proof.
@@
 $-\frac{1}{2}d_{A}\left(\psi^{0,1_{J}}(\dot{J})\right)+\frac{1}{2}d_{A}\left(\psi^{1,0_{J}}(\dot{J})\right)$
-$=i\left(d_{A}\dot{\psi}^{1,0_{J}}+[a,\varphi]-\frac{i}{2}d_{A}\left(\psi(\dot{J})\right)\right)=0$
+$=i\left(d_{A}\dot{\psi}^{1,0_{J}}+[a,\varphi]-\frac{i}{2}d_{A}\left(\psi(\dot{J})\right)\right)=0,$
 and hence the tangent space to $\curly{X}^{Higgs}$ is preserved by $\mathbb{I}$. ∎
@@
 The first step is to show that the integrability condition for $\varphi$ is compatible with the group $\widetilde{\curly{G}}$ defined in (4.1).
-We observe first that there is a holomorhic action of $\widetilde{\curly{G}}$ on $\curly{J}\times{\curly A}$, given by
+We observe first that there is a holomorphic action of $\widetilde{\curly{G}}$ on $\curly{J}\times{\curly A}$, given by
 $g(J,A)=(p(g)_{*}J,A_{g}),$
-where $A_{g}$ is the defined via the Chern-Singer correspondence
+where $A_{g}$ is the defined via the Chern–Singer correspondence:
 $A_{g}=A_{h,g_{*}\bar{\partial}_{J,A}}.$
@@
 The $\widetilde{\curly{G}}-$action on $\curly{X}$ preserves the subsets
-$\curly{X}^{ps}\subset\curly{X}^{Higgs}\subset\curly{X}$
+$\curly{X}^{ps}\subset\curly{X}^{Higgs}\subset\curly{X},$
 and we define the universal moduli space of $G$-Higgs bundles on $E_{G}$ as the quotient
 $\mathcal{U}^{Higgs}(G):=\curly{X}^{ps}/\widetilde{\curly{G}}$.
-By construction, $\mathcal{U}^{Higgs}(G)$ fibers over the Teichmüller space $\mathcal{T}$ of complex structures modulo diffeomorphisms isotopic to the identity
-(5.3) $\mathcal{U}^{Higgs}(G)\to\mathcal{T}:=\curly{J}/\operatorname{Diff}_{0}(\Sigma).$
+By construction, $\mathcal{U}^{Higgs}(G)$ fibres over the Teichmüller space $\mathcal{T}$ of complex structures modulo diffeomorphisms isotopic to the identity
+(5.3) $\mathcal{U}^{Higgs}(G)\longrightarrow\mathcal{T}:=\curly{J}/\operatorname{Diff}_{0}(\Sigma).$
 This fibration is naturally holomorphic with respect to the complex structure $\mathbb{I}$.
@@
 5.2. Kähler fibration for the universal Higgs field
-Consider the structure of Kähler fibration $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ on $(\curly{X},\mathbb{I})\to\curly{J}$ defined fibrewise by the symplectic structure $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ on ${\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$.
+Consider the structure of Kähler fibration $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ on $(\curly{X},\mathbb{I})\to\curly{J}$ defined fibrewise by the symplectic structure $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ on ${\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$.
 More explicitly,
 (5.4) $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2}))=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2}).$
-The aim of this section is to prove that such structure admits a Kähler Ehresmann connection compatible with Hitchin’s equations (2.15).
-The situation here is opposite to that on Section 3.2: the holomorphic fibration defined by $\mathbb{I}$ is non-trivial, whereas $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ defines a trivial symplectic fibration, constant along $\curly{X}$.
+The aim of this section is to show that such a structure admits a Kähler Ehresmann connection compatible with Hitchin’s equations (2.12).
+The situation here is opposite to that in Section 3.2: the holomorphic fibration defined by $\mathbb{I}$ is non-trivial, whereas $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ defines a trivial symplectic fibration, constant along $\curly{X}$.
 Nonetheless, the combination of these two structures defines a non-trivial structure of Kähler fibration on $\curly{X}\to\curly{J}$, different from the one considered in Section 3.2.
-In order to achieve our goal in this section, we will apply Theorem 3.2. As we have seen in Lemma 5.2, the integrability condition
+In order to achieve our goal in this section, we will apply Theorem 3.2. As we have seen in Proposition 5.2, the integrability condition
 $\bar{\partial}_{J,A}\psi^{1,0_{J}}=0$
 cuts a holomorphic submanifold $\curly{X}^{Higgs}$ of $(\curly{X},\mathbb{I})$.
-Given that $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ is constant on $\curly{X}$, it provides a candidate for coupling two-form $\boldsymbol{\sigma}$ on $(\curly{X},\widehat{\boldsymbol{\omega}}_{\mathbf{I}})$, inducing the trivial Ehresmann connection on $\curly{X}\to\curly{J}$, regarded as the product $\curly{X}=\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k})$.
-However, one can readily check that the horizontal subspace of this connection is not tangent to the submanifold $\curly{X}^{Higgs}$, and hence is not appropiate for our goals.
+Given that $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ is constant on $\curly{X}$, it provides a candidate for a coupling 2-form $\boldsymbol{\sigma}$ on $(\curly{X},\widehat{\boldsymbol{\omega}}_{\mathbf{I}})$, inducing the trivial Ehresmann connection on $\curly{X}\to\curly{J}$, regarded as the product $\curly{X}=\curly{J}\times{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$.
+However, one can readily check that the horizontal subspace of this connection is not tangent to the submanifold $\curly{X}^{Higgs}$, and hence is not appropriate for our goals.
 Motivated by this, we consider the Ehresmann connection $\Gamma^{\mathbb{I}}$ with horizontal subspace
-$H^{\mathbb{I}}=\Big\{\left(\dot{J},0,-\tfrac{1}{2}\psi(J\dot{J})\right)\;|\;\dot{J}\in T_{J}\curly{J}\Big\}.$
+$H^{\mathbb{I}}=\Big\{\left(\dot{J},0,-\tfrac{1}{2}\psi(J\dot{J})\right)\;\Big|\;\dot{J}\in T_{J}\curly{J}\Big\}.$
 Lemma 5.4.
@@
 $\Gamma^{\mathbb{I}}\circ\mathbb{I}=\mathbf{I}\circ\Gamma^{\mathbb{I}},$
-where $\mathbf{I}$ denotes the complex structure along the fibres. Furthermore, its curvature $F_{\mathbb{I}}\in\Omega^{2}(\curly{X},V\curly{X})$ is of type $(1,1)$ and given explicitly by
+where $\mathbf{I}$ denotes the complex structure along the fibres. Furthermore, its curvature $F_{\mathbb{I}}\in\Omega^{2}(\curly{X},V\curly{X})$ is of type $(1,1)$, and given explicitly by
 (5.5) $\begin{split}(F_{\mathbb{I}})_{|(J,A,\psi)}(v_{1},v_{2})&=\frac{1}{4}\left(0,0,\psi([\dot{J}_{1},\dot{J}_{2}])\right),\end{split}$
@@
 Given $(J,A,\psi)\in\curly{X}$ and any pair of tangent vectors $v_{j}=(\dot{J}_{j},a_{j},\dot{\psi}_{j})\in T_{(J,A,\psi)}\curly{X}$, one has
-(5.6) $\begin{split}\widehat{\boldsymbol{\omega}}_{\mathbf{I}}(\Gamma^{\mathbb{I}}v_{1},\Gamma^{\mathbb{I}}v_{2})&=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{1}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{1}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right)\end{split}$
+(5.6) $\begin{split}\widehat{\boldsymbol{\omega}}_{\mathbf{I}}(\Gamma^{\mathbb{I}}v_{1},\Gamma^{\mathbb{I}}v_{2})&=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{1}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{1}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right).\end{split}$
 Our next result proves that the Ehresmann connection $\Gamma^{\mathbb{I}}$ is Kähler, by calculation of an explicit coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$ via equation (3.3).
-Define a basic two-from $\mu\in\Omega^{2}(\curly{X})$ by
+Define a basic 2-form $\mu\in\Omega^{2}(\curly{X})$ by
 $\mu\left((\dot{J}_{1},a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2})\right)=\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right).$
 Proposition 5.6.
-The Ehresmann connection $\Gamma^{\mathbb{I}}$ is Kähler. Furthermore, the following formula gives a coupling form for $\Gamma^{\mathbb{I}}$
+The Ehresmann connection $\Gamma^{\mathbb{I}}$ is Kähler. Furthermore, the following formula gives a coupling form for $\Gamma^{\mathbb{I}}$:
 $\boldsymbol{\sigma}^{\mathbb{I}}:=\widehat{\boldsymbol{\omega}}_{\mathbf{I}}(\Gamma^{\mathbb{I}},\Gamma^{\mathbb{I}})-\mu.$
@@
 $=\frac{1}{4}\int_{\Sigma}B\left(\dot{\psi}\wedge\psi([\dot{J}_{1},\dot{J}_{2}])\right)$
-$=-\widehat{\boldsymbol{\omega}}_{\mathbf{I}}((0,a,\dot{\psi}),F_{\mathbb{I}}(v_{1},v_{2})).$
-To finish this section, we provide a formula for the the symmetric tensor on $\curly{X}$ associated to the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$, explicitly given by
+$=-\widehat{\boldsymbol{\omega}}_{\mathbf{I}}((0,a,\dot{\psi}),F_{\mathbb{I}}(v_{1},v_{2})).\qed$
+To finish this section, we provide a formula for the symmetric tensor on $\curly{X}$ associated to the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$, explicitly given by
 $\mathbf{g}_{\mathbb{I}}:=\boldsymbol{\sigma}^{\mathbb{I}}(,\mathbb{I}).$
 By construction, this coincides with the flat hyperkähler metric (2.8) along the fibres of $\curly{X}\to\curly{J}$.
-Similarly as in Corollary 3.7, $\mathbf{g}_{\mathbb{I}}$ is negative semi-definite along the horizontal directions of the connection $\Gamma^{\mathbb{I}}$.
+As in Corollary 3.7, $\mathbf{g}_{\mathbb{I}}$ is negative semi-definite along the horizontal directions of the connection $\Gamma^{\mathbb{I}}$.
 Corollary 5.7.
-For any tangent vector $v=(\dot{J},a,\dot{\psi}),\in T_{(J,A,\psi)}\curly{X}$ one has
-(5.8) $\begin{split}\mathbf{g}_{\mathbb{I}}(v,v)&=\int_{\Sigma}B(a\wedge Ja)+\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{1}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{1}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right)\end{split}$
+For any tangent vector $v=(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$ one has
+(5.8) $\begin{split}\mathbf{g}_{\mathbb{I}}(v,v)&=\int_{\Sigma}B(a\wedge Ja)+\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{1}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}+\tfrac{1}{2}\psi(J\dot{J})\right)\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right).\end{split}$
 In particular, given a horizontal vector field $v\in H^{\mathbb{I}}$ at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$, one has
@@
 Proof.
-Formula (5.8) follows from (5.7), using that $\Gamma^{\mathbb{I}}\circ\mathbb{I}=\mathbf{I}\circ\Gamma^{\mathbb{I}}$ (see Lemma 5.4).
-5.3. Hamiltonian action and coupled Hitchin’s equations
+Formula (5.8) follows from (5.7), using that $\Gamma^{\mathbb{I}}\circ\mathbb{I}=\mathbf{I}\circ\Gamma^{\mathbb{I}}$ (see Lemma 5.4). ∎
+5.3. Hamiltonian action and the coupled Hitchin equations
 We follow the notation of the previous section.
-Our aim in this section is to investigate a natural Hamiltonian action for minimal couplings associated to the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$ for the connection $\Gamma^{\mathbb{I}}$, in the Kähler fibration $(\curly{X},\mathbb{I},\widehat{\boldsymbol{\omega}}_{\mathbf{I}})$ over the space of complex structure $\curly{J}$.
+Our aim now is to investigate a natural Hamiltonian action for the minimal couplings defined by the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$ associated to the connection $\Gamma^{\mathbb{I}}$, on the Kähler fibration $(\curly{X},\mathbb{I},\widehat{\boldsymbol{\omega}}_{\mathbf{I}})$ over the space of complex structures $\curly{J}$.
 We fix a symplectic form $\omega$ on $\Sigma$, compatible with the orientation.
-Given $\alpha>0$ a real coupling constant and $\varepsilon\in\{-1,1\}$, the family of minimal coupling symplectic structures of our interest is defined by
+Given $\alpha>0$ a real ‘coupling constant’ and $\varepsilon\in\{-1,1\}$, the family of minimal coupling symplectic structures of our interest is defined by
 (5.10) $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}=\varepsilon\boldsymbol{\omega}{J}+\alpha\boldsymbol{\sigma}^{\mathbb{I}},$
@@
 (5.11) $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}:=\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}(,\mathbb{I})=\varepsilon\mathbf{g}_{\curly{J}}+\alpha\mathbf{g}_{\mathbb{I}}.$
-The following result is the analogue of Lemma 3.8 in the present setup. The proof is analogue, by application in this case of Corollary 5.7, and is ommitted.
+The following result is the analogue of Lemma 3.8 in the present setup. Its proof is analogous, by application of Corollary 5.7 in this case, and is omitted.
 Lemma 5.8.
-Given a horizontal vector field $v\in H^{\mathbb{I}}$ at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$, one has
+Let $v\in H^{\mathbb{I}}$ be a horizontal vector field at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$. Then
 (5.12) $\begin{split}\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(v,v)=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}\dot{J})\omega-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J(\psi(\dot{J}))\right).\end{split}$
-Consequently, for any $\alpha>0$ one has
+Consequently, for any $\alpha>0$:
   1. (1)
-    For $\varepsilon=-1$, $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is negative definite along $H^{\mathbb{I}}$.
+    If $\varepsilon=-1$, $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is negative definite along $H^{\mathbb{I}}$.
 In particular, $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,-1}$ is a non-degenerate symplectic structure.
   2. (2)
-    For $\varepsilon=1$ and $\psi\neq 0$, $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ changes signature along the line $(J,A,\lambda\psi)\in\curly{X}$, for $\lambda\in\mathbb{R}$.
+    If $\varepsilon=1$ and $\psi\neq 0$, $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ changes signature along the line $(J,A,\lambda\psi)\in\curly{X}$, for $\lambda\in\mathbb{R}$.
 Our next goal is to prove that $(\curly{X},\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon})$ admits a Hamiltonian action by the extended gauge group $\widetilde{{\curly K}}$, determined by $\omega$ and the reduction $E_{K}\subset E_{G}$.
 For this, we will need some refined information about the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$.
-Consider the closed two form $\pi^{*}_{2}\omega_{\mathbf{I}}$ on $\curly{X}$, given by the pull-back of $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ via the canonical projection $\pi_{2}\colon\curly{X}\to{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$.
+Consider the closed 2-form $\pi^{*}_{2}\omega_{\mathbf{I}}$ on $\curly{X}$, given by the pull-back of $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ via the canonical projection $\pi_{2}\colon\curly{X}\to{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$.
 Explicitly, for $v_{j}=(\dot{J}_{j},a_{j},\dot{\psi}_{j})\in T_{(J,A,\psi)}\curly{X}$, $j=1,2$, one has
 $\pi^{*}_{2}\omega_{\mathbf{I}}(v_{1},v_{2})=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2}).$
-The two-form $\pi^{*}_{2}\omega_{\mathbf{I}}$ restricts on the fibres to $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$, and hence, by the general theory of symplectic fibrations (see [GLS, Section 1.6]), one expects that there exists a one-form $\lambda\in\Omega^{1}(\curly{X})$ such that
+The 2-form $\pi^{*}_{2}\omega_{\mathbf{I}}$ restricts on the fibres to $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$, and hence, by the general theory of symplectic fibrations (see [GLS, Section 1.6]), one expects that there exists a 1-form $\lambda\in\Omega^{1}(\curly{X})$ such that
 (5.13) $\boldsymbol{\sigma}^{\mathbb{I}}=\pi^{*}_{2}\omega_{\mathbf{I}}+d\lambda.$
@@
 Lemma 5.9.
-Consider the one-form $\lambda\in\Omega^{1}(\curly{X})$ defined by
+Consider the 1-form $\lambda\in\Omega^{1}(\curly{X})$ defined by
 $\lambda(v)=\frac{1}{4}\int_{\Sigma}B\left(\psi(J\dot{J})\wedge\psi\right)$
-for $v=(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$. Then, formula (5.13) holds. Furthemore, considering the $L^{2}$-norm of the unitary Higgs field
+for $v=(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$. Then formula (5.13) holds. Furthermore, considering the $L^{2}$-norm of the ‘unitary Higgs field’,
 $\nu(J,A,\psi)=\|\psi\|^{2}_{L^{2}}:=\int_{\Sigma}B(\psi\wedge J\psi),$
 one has
-(5.14) $\boldsymbol{\sigma}^{\mathbb{I}}=\pi^{*}_{2}\omega_{{\curly A}}+\tfrac{i}{2}\partial_{\mathbb{I}}\bar{\partial}_{\mathbb{I}}\nu$
+(5.14) $\boldsymbol{\sigma}^{\mathbb{I}}=\pi^{*}_{2}\omega_{{\curly A}}+\tfrac{i}{2}\partial_{\mathbb{I}}\bar{\partial}_{\mathbb{I}}\nu,$
 where
@@
 $=-4\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})+4\boldsymbol{\sigma}^{\mathbb{I}}-4\pi^{*}_{2}\omega_{\mathbf{I}}$
-$=-4\pi^{*}_{2}\omega_{{\curly A}}+4\boldsymbol{\sigma}^{\mathbb{I}}.$
+$=-4\pi^{*}_{2}\omega_{{\curly A}}+4\boldsymbol{\sigma}^{\mathbb{I}}.\qed$
 Consider the extended gauge group $\widetilde{{\curly K}}$ of $h$ and $\omega$.
-Similarly as in [AGG, Section 2.2], the group $\widetilde{{\curly K}}$ acts on $(\curly{X},\mathbb{I},\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon})$ preserving $\mathbb{I}$ and $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$, and covering the $\curly{H}$-action on $\curly{J}$ by push-forward.
+As in [AGG1, Section 2.2], the group $\widetilde{{\curly K}}$ acts on $(\curly{X},\mathbb{I},\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon})$ preserving $\mathbb{I}$ and $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$, and covering the $\curly{H}$-action on $\curly{J}$ by push-forward.
 We are ready to prove the main result of this section.
@@
 $\langle\mu_{\widetilde{{\curly K}}}^{\mathbb{I}}(J,A,\psi),\zeta\rangle$ $=\alpha\int_{\Sigma}B\left(F_{A}-\tfrac{1}{2}[\psi\wedge\psi],A\zeta\right)$
-$-\int_{\Sigma}f\left(\varepsilon S_{J}\omega+\alpha d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-\alpha d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)$
+$-\int_{\Sigma}f\left(\varepsilon S_{J}\omega+\alpha d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-\alpha d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right).$
 Proof.
@@
 $=\int_{\Sigma}d^{c}f\wedge B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi)))\right)$
-$=-\int_{\Sigma}fd^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)$
-Similarly as in Proposition 3.9, the $\widetilde{{\curly K}}-$action produces a coupling term in the base $\Sigma$ (cf.
-Remark 4.2), which combines with the scalar curvature of the metric $g=\omega(,J)$.
-In particular, similarly as in [AGG], zeros of the moment map $\mu_{\widetilde{{\curly K}}}^{\mathbb{I}}$ are given by solutions of the coupled system of equations
-(5.15) $\begin{split}F_{A}-\tfrac{1}{2}[\psi\wedge\psi]&=0\\ \varepsilon S_{J}+\alpha\Lambda_{\omega}\left(d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
-In the next section we combine these equations with the integrability condition $\bar{\partial}_{A}\varphi=0$ for the Higgs field, in order to introduce the universal moduli space of Hitchin’s equations.
+$=-\int_{\Sigma}fd^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right).\qed$
+As in Proposition 3.9, the $\widetilde{{\curly K}}-$action produces a ‘coupling term’ on the base $\Sigma$ (cf.
+Remark 4.2), which interacts with the scalar curvature of the metric $g=\omega(,J)$.
+In particular, as in [AGG1], zeros of the moment map $\mu_{\widetilde{{\curly K}}}^{\mathbb{I}}$ are given by solutions to the coupled system of equations
+(5.15) $\begin{split}F_{A}-\tfrac{1}{2}[\psi\wedge\psi]&=0,\\ \varepsilon S_{J}+\alpha\Lambda_{\omega}\left(d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V}.\end{split}$
+In the next section, we combine these equations with the integrability condition $\bar{\partial}_{A}\varphi=0$ for the Higgs field, in order to introduce the universal moduli space of Hitchin’s equations.
 5.4. Metric structure of the universal moduli of Hitchin’s equations
-In this section we combine the results of the previous two sections in order to define a universal moduli spaces of solutions of Hitchin’s equations (2.15), varying over (a cover of) Teichmuller space.
+In this section we combine the results of the previous two sections to define a universal moduli space of solutions of Hitchin’s equations (2.12), varying over (a cover of) the Teichmüller space.
 As in the previous sections, we consider a fixed principal $K$-bundle $E_{K}$ over a smooth compact oriented surface $\Sigma$ with fixed symplectic form $\omega$.
 Definition 5.11.
-We say that a triple $(J,A,\psi)\in\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ is a solution of the coupled Hitchin’s equations, if the following conditions are satisfied
+A triple $(J,A,\psi)\in\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$ is a solution of the coupled Hitchin equations, if the following conditions are satisfied
 (5.16) $\begin{split}F_{A}-[\varphi,\tau(\varphi)]&=0,\\ \bar{\partial}_{J,A}\varphi&=0,\\ S_{g}&=\frac{2\pi\chi(\Sigma)}{V},\end{split}$
@@
 $\mathcal{U}^{Hit}(G)_{\alpha}^{\varepsilon}:=(\mu_{\widetilde{{\curly K}}^{\mathbb{I}}|\curly{X}^{Higgs}})^{-1}(0)/\widetilde{{\curly K}}$.
-By the explicit formula for the moment in Proposition 5.10, the underlying space to $\mathcal{U}^{Hit}(G)_{\alpha}^{\varepsilon}$ is independent of the choice of parameters, as the coupling terms in the scalar equation in (5.15) vanishes identically upon imposing the integrability condition $\bar{\partial}_{J,A}\varphi=0$.
+By the explicit formula for the moment map in Proposition 5.10, the underlying space of $\mathcal{U}^{Hit}(G)_{\alpha}^{\varepsilon}$ is independent of the choice of parameters, since the ‘coupling term’ in the scalar equation in (5.15) vanishes identically when the integrability condition $\bar{\partial}_{J,A}\varphi=0$ is imposed.
 We shall denote this moduli space simply by $\mathcal{U}^{Hit}(G)$.
@@
 (5.3))
-$\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)\to\mathcal{T}.$
+$\mathcal{U}^{Hit}(G)\longrightarrow\mathcal{U}^{Higgs}(G)\longrightarrow\mathcal{T}.$
 Hence, in particular, $\mathcal{U}^{Hit}(G)$ can be regarded as a fibration over the Teichmüller space $\mathcal{T}$.
-Similarly as in Section 4.2, since the symmetric $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is not positive definite (see Lemma 5.8), it is not obvious a priori that $\mathcal{U}^{Hit}(G)$ inherits a complex structure compatible with the (pre)symplectic structure induced by $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$.
-The main goal of this section is to study sufficient conditions under which this natural condition for the moduli space holds, furthermore proving that the map $\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)$ is holomorphic.
-The analysis is analogue to that in Sections 4.2 and 3.3, and hence we just outline the argument.
-The first step is to undertake a gauge fixing for solutions of the coupled Hitchin’s equations (5.16), whereby the complex structure (5.1) and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ descend to the moduli space.
+As in Section 4.2, since the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is not positive definite (see Lemma 5.8), it is not obvious a priori whether $\mathcal{U}^{Hit}(G)$ inherits a complex structure compatible with the (pre)symplectic structure induced by $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$.
+The main goal of this section is to study sufficient conditions under which this natural condition holds for the moduli space, and to prove that the map $\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)$ is holomorphic.
+The analysis is analogous to that in Sections 4.2 and 3.3, so we only outline the argument.
+The first step is to undertake a gauge fixing for solutions of the coupled Hitchin equations (5.16), whereby the complex structure (5.1) and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ descend to the moduli space.
 We observe that this will produce a priori different complex structures on $\mathcal{U}^{Hit}(G)$.
-We fix a solution $(J,A,\psi)$ of the coupled Hitchin’s equations (5.16).
+We fix a solution $(J,A,\psi)$ of the coupled Hitchin equations (5.16).
 Lemma 5.12.
-The linearization of the coupled Hitchin’s equations (5.16) at $(J,A,\psi)$ is given by
+The linearization of the coupled Hitchin equations (5.16) at $(J,A,\psi)$ is given by
 (5.17) $\begin{split}d_{A}a-[\dot{\psi},\psi]&=0,\\ d_{A}\dot{\psi}+[a,\psi]&=0,\\ d_{A}(J\dot{\psi})+[a,J\psi]-d_{A}(\psi(\dot{J}))&=0,\\ \varepsilon\delta S(\dot{J})&=0,\end{split}$
@@
 Using the same notation as in (4.7), we define a complex of linear differential operators
-(5.18) ${(\mathcal{S}^{*}_{0})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{S}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}}$${\mathcal{S}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L}}$${\mathcal{S}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
-The cohomology $H^{1}(\mathcal{S}^{*}_{0}):=\frac{\operatorname{Ker}{\bf L}}{\operatorname{Im}\,\mathbf{P}}$ can be formally identified with the tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{Hit}(G)$.
+[…]
+The cohomology $H^{1}(\mathcal{S}^{*}_{0}):=\frac{\ker{\bf L}}{\operatorname{Im}\,\mathbf{P}}$ can be formally identified with the tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{Hit}(G)$.
 Our next result shows that the moduli space $\mathcal{U}^{Hit}(G)$ is finite dimensional.
@@
 In order to construct a complex structure induced by (5.1) on the moduli space, we work orthogonal to the image of the infinitesimal action operator $\mathbf{P}$ in (5.18) with respect to the indefinite pairing $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$.
-The existence of this complex structure will automatically yield a symmetric tensor of type $(1,1)$, since the two-form $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$ is well defined on the cohomology $H^{1}(\mathcal{S}^{*}_{0})$ by Proposition 5.10.
+The existence of this complex structure will automatically yield a symmetric tensor of type $(1,1)$, since the 2-form $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$ is well defined on the cohomology $H^{1}(\mathcal{S}^{*}_{0})$ by Proposition 5.10.
 Consider the $L^{2}$-pairing on $\mathcal{S}^{0}$ define by (4.8).
@@
 Proposition 5.10)
-$\mu^{\mathbb{I}}(J,A,\psi)=\left(\mu^{\mathbb{I}}_{0},\mu^{\mathbb{I}}_{1}\right).$
+$\mu^{\mathbb{I}}(J,A,\psi)=\left(\mu^{\mathbb{I}}_{0},\mu^{\mathbb{I}}_{1}\right),$
 where
-$\mu^{\mathbb{I}}_{0}$ $=-\left(\varepsilon S_{J}\omega+\alpha d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-\alpha d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)$
+$\mu^{\mathbb{I}}_{0}$ $=-\left(\varepsilon S_{J}\omega+\alpha d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-\alpha d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right),$
 $\mu^{\mathbb{I}}_{0}$ $=\alpha\left(F_{A}-\tfrac{1}{2}[\psi\wedge\psi]\right).$
 Lemma 5.14.
-The following operator provides a formal adjoint of the infinitesimal action $\mathbf{P}$ for the pairings (4.8) and (5.11)
-$\mathbf{P}^{*}_{\alpha,\varepsilon}=\delta\mu^{\mathbb{I}}\circ\mathbb{I}\colon\mathcal{S}^{1}\to\mathcal{S}^{0}.$
+The following operator provides a formal adjoint of the infinitesimal action $\mathbf{P}$ for the pairings (4.8) and (5.11):
+$\mathbf{P}^{*}_{\alpha,\varepsilon}=\delta\mu^{\mathbb{I}}\circ\mathbb{I}\colon\mathcal{S}^{1}\longrightarrow\mathcal{S}^{0}.$
 Consider now the differential operator
@@
 The key condition on the solution $(J,A,\psi)$ of (5.16) which we need to assume in order to construct the complex structure on the moduli space is the vanishing of the kernel of $\mathcal{L}_{0}^{\alpha,\varepsilon}$.
-Notice that, unlike in the standard cases in gauge theory in which the parameter space metric is positive definite, $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}$ does not relate in general to automorphisms of the triple $(J,\psi,A)$, but rather to null vectors with respect to $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$.
+Notice that, unlike in the standard cases in gauge theory in which the parameter space metric is positive definite, $\ker\mathcal{L}_{0}^{\alpha,\varepsilon}$ does not relate in general to automorphisms of the triple $(J,\psi,A)$, but rather to null vectors with respect to $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$.
 Proposition 5.15.
-The operator $\mathcal{L}_{0}^{\alpha,\varepsilon}$ is Fredholm with zero index. Furthermore, elements $(f,u)\in\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}$ are smooth.
-Assuming that $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}$ is trivial, in the next result we obtain a natural gauge fixing via a $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$-orthogonal decomposition
+The operator $\mathcal{L}_{0}^{\alpha,\varepsilon}$ is Fredholm with zero index. Furthermore, elements $(f,u)\in\ker\mathcal{L}_{0}^{\alpha,\varepsilon}$ are smooth.
+Assuming that $\ker\mathcal{L}_{0}^{\alpha,\varepsilon}$ is trivial, in the next result we obtain a natural gauge fixing via a $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$-orthogonal decomposition
 (5.20) $\mathcal{S}^{1}=\operatorname{Im}\;\mathbf{P}\oplus(\operatorname{Im}\;\mathbf{P})^{\perp_{\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}}}.$
 Lemma 5.16.
-Assume that $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}$.
+Assume that $\ker\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}$.
 Then, there exists an orthogonal decomposition (5.20) for the pairing $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$.
@@
 Consequently, we have
-$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})=\ker\bf{L}\cap\ker{\bf L}\circ\mathbb{I}.$
-Our next result provides our gauge fixing mechanism for the linearization of the coupled Hitchin’s equations (5.16).
+$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})=\ker\bf{L}\cap\ker({\bf L}\circ\mathbb{I}).$
+Our next result provides our gauge fixing mechanism for the linearization of the coupled Hitchin equations (5.16).
 Proposition 5.17.
-Assume $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}$. Then, the inclusion $\mathcal{H}^{1}(\mathcal{S}^{*}_{0})\subset\ker\bf{L}$ induces an isomorphism
-$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})\simeq H^{1}(\mathcal{S}^{*}_{0}).$
+Assume $\ker\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}$. Then, the inclusion $\mathcal{H}^{1}(\mathcal{S}^{*}_{0})\subset\ker\bf{L}$ induces an isomorphism
+$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})\cong H^{1}(\mathcal{S}^{*}_{0}).$
 More precisely, any class in the cohomology $H^{1}(\mathcal{S}^{*}_{0})$ of the complex (4.7) admits a unique representative $(\dot{J},a,\dot{\psi})$ solving the linear equations
@@
 Define
-$\mathcal{U}^{*}_{\alpha,\varepsilon}=\{[(J,A,\psi)]\;|\;\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}\}\subset\mathcal{U}^{Hit}(G).$
+$\mathcal{U}^{*}_{\alpha,\varepsilon}=\{[(J,A,\psi)]\;|\;\ker\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}\}\subset\mathcal{U}^{Hit}(G).$
 Theorem 5.18.
@@
 Then, one has
-(5.22) $\begin{split}\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a_{1}\wedge a_{2})-\alpha\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{\alpha}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{\alpha}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{\alpha}{2}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right),\\ \mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a\wedge Ja)+\alpha\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{\alpha}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right),\end{split}$
+(5.22) $\begin{split}\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a_{1}\wedge a_{2})-\alpha\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{\alpha}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{\alpha}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{\alpha}{2}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right),\\ \mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a\wedge Ja)+\alpha\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{\alpha}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right).\end{split}$
 5.5. Comparison with $\boldsymbol{\mathcal{U}^{Higgs}(G)}$ and existence
@@
 As we will see, Theorem 5.18 induces a natural holomorphic map
-$\mathcal{U}^{Hit}(G)\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{Higgs}(G)$
-and hence a holomorphic map into Teichmuller space $\mathcal{T}$.
+$\mathcal{U}^{Hit}(G)\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\longrightarrow\mathcal{U}^{Higgs}(G),$
+and hence a holomorphic map into Teichmüller space $\mathcal{T}$.
 We will also prove that for genus of the surface $\Sigma$ bigger than zero, the open $\mathcal{U}^{*}\subset\mathcal{U}^{Hit}(G)$ is non-empty for sufficiently small values of $\alpha$.
@@
 Consider a point $(J,A,\psi)]\in\mathcal{U}^{Higgs}(G)$, regarded as the $\widetilde{\curly{G}}-$orbit of $(J,A,\psi)\in\curly{X}$ solving the equations
-(5.23) $\begin{split}\bar{\partial}_{J,A}\psi^{1,0_{J}}=0\end{split}$
+(5.23) $\begin{split}\bar{\partial}_{J,A}\psi^{1,0_{J}}=0.\end{split}$
 The tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{Higgs}(G)$ can be formally identified with the cohomology of the complex of linear differential operators
-(5.24) ${(\mathcal{C}_{0}^{*})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{C}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}^{c}_{0}}$${\mathcal{C}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L^{c}_{0}}}$${\mathcal{C}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
+[…]
 where we have
-$\mathcal{C}^{0}=\operatorname{Lie}\widetilde{\curly{G}}\cong\Omega^{0}(T\Sigma)\oplus\Omega^{0}(X,E_{G}(\mathfrak{g})),\qquad\mathcal{C}^{1}=\mathcal{S}^{1},\qquad\mathcal{C}^{2}=\Omega^{2}(X,E_{G}(\mathfrak{g}))$
+$\mathcal{C}^{0}=\operatorname{Lie}\widetilde{\curly{G}}\cong\Omega^{0}(T\Sigma)\oplus\Omega^{0}(\Sigma,E_{G}(\mathfrak{g})),\qquad\mathcal{C}^{1}=\mathcal{S}^{1},\qquad\mathcal{C}^{2}=\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))$
 and
@@
 The sequence (5.24) is an elliptic complex of degree-one linear differential operators. Consequently, the cohomology groups $H^{j}(\mathcal{C}^{*})$, with $j=0,1,2$, are finite-dimensional.
-Applying Theorem 2.7, a solution $(J,A,\psi)$ of the coupled Hitchin’s equations (5.16) induces a polystable $G$-Higgs bundle $(E,\varphi)\cong(\bar{\partial}_{J,A},\psi^{1,0_{J}})$.
+Applying Theorem 2.7, a solution $(J,A,\psi)$ of the coupled Hitchin equations (5.16) induces a polystable $G$-Higgs bundle $(E,\varphi)\cong(\bar{\partial}_{J,A},\psi^{1,0_{J}})$.
 This fact, jointly with the natural inclusion $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$, leads to a continuous map
-(5.25) $\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)\colon[(J,A,\psi)]\mapsto[(J,A,\psi)].$
+(5.25) $\begin{array}[]{rcl}\mathcal{U}^{Hit}(G)&\longrightarrow&\mathcal{U}^{Higgs}(G)\\ \left[(J,A,\psi)\right]&\longmapsto&\left[(J,A,\psi)\right].\end{array}$
 By definition of $\mathcal{U}^{Hit}(G)$, there is a continuous diagram of moduli spaces
-(5.26) ${\mathcal{U}^{Hit}(G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{U}^{Higgs}(G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mu_{\curly{H}}^{-1}(0)/\curly{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}}$
+(5.26) $\mathcal{U}^{Hit}(G)$$\mathcal{U}^{Higgs}(G)$$\mu_{\curly{H}}^{-1}(0)/\curly{H}$$\mathcal{T}$}
 where $\mu_{\curly{H}}^{-1}(0)/\curly{H}$ is the moduli space of constant scalar curvature Kähler metrics with total volume $V$, modulo $\omega$-Hamiltonian diffeomorphisms.
-The fibre of $\mu_{\curly{H}}^{-1}(0)/\curly{H}\to\mathcal{T}$ can be identified with $H^{1}(\Sigma,\mathbb{R})$ [F].
+The fibre of $\mu_{\curly{H}}^{-1}(0)/\curly{H}\to\mathcal{T}$ can be identified with $H^{1}(\Sigma,\mathbb{R})$ [Fu].
 Building on Theorem 5.18, our next goal is to prove that this induces a holomorphic map $\mathcal{U}^{Hit}(G)\supset\mathcal{U}^{*}\to\mathcal{U}^{Higgs}(G)$.
@@
 Let $(J,A,\psi)]\in\mathcal{U}^{*}_{0}$. Then, (5.25) induces a complex linear map
-(5.27) ${H^{1}(\mathcal{S}^{*}_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathcal{C}^{*}_{0}).}$
+(5.27) $H^{1}(\mathcal{S}^{*}_{0})$$H^{1}(\mathcal{C}^{*}_{0}),$}
 where the complex structure on $H^{1}(\mathcal{S}^{*}_{0})$ is the one induced by Proposition 5.17.
@@
 For this, we fix a compact Riemann surface $X=(\Sigma,J)$ and consider a $G$-Higgs bundle $(E,\varphi)$.
-In this setup, consider the coupled equations
-(5.28) $\begin{split}F_{h}-[\varphi\wedge\tau_{h}\varphi]&=0\\ \varepsilon S_{g}&=\varepsilon\frac{2\pi\chi(\Sigma)}{V}\end{split}$
-for pairs $(g,h)$, where $g$ is a Kähler metric on $X$ with total volume $V$ and $h\in\Omega^{0}(E_{G}(G/K))$ is a reduction of structure group of $E$ to $K$.
+In this setup, we consider the coupled equations
+(5.28) $\begin{split}F_{h}-[\varphi\wedge\tau_{h}\varphi]&=0,\\ \varepsilon S_{g}&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
+for pairs $(g,h)$ consisting of a Kähler metric $g$ on $X$ with total volume $V$ and a reduction $h\in\Omega^{0}(E_{G}(G/K))$ of the structure group of $E$ to $K$.
 A solution $(g,h)$ of the equations (5.28) determines then a solution of (5.16), given by the triple $(J,A_{h},-i(\varphi-\tau_{h}\varphi))$.
@@
 Furthermore, the induced maps $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{Higgs}(G)$ and
-(5.29) $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mu_{\curly{H}}^{-1}(0)/\curly{H}$
+(5.29) $\mathcal{U}^{*}_{\alpha,\varepsilon}\longrightarrow\mu_{\curly{H}}^{-1}(0)/\curly{H}$
 are holomorphic, where $\mu_{\curly{H}}^{-1}(0)/\curly{H}$ is the moduli space of constant scalar curvature Kähler metrics with total volume $V$.
-The restriction of the $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ in Corollary 5.19 to the fibres of (5.29) is $\alpha\mathbf{g}$, where $\mathbf{g}$ denotes the hyperkähler metric on the moduli space of solutions of Hitchin’s equations.
+The restriction of the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ in Corollary 5.19 to the fibres of (5.29) is $\alpha\mathbf{g}$, where $\mathbf{g}$ denotes the hyperkähler metric on the moduli space of solutions of Hitchin’s equations.
 Consequently, (5.29) has a natural structure of Kähler fibration with coupling form $\boldsymbol{\omega}_{1,0}$ (see (5.22)) and Kähler Ehresmann connection.
 Proof.
-The proof is analogue of that of Theorem 4.14, and is therefore omitted.
+The proof is analogous to that of Theorem 4.14 and is therefore omitted.
 The only difference is that we can control the non-degeneracy of $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ along the fibres over the moduli space of constant scalar curvature Kähler metrics $\mathcal{N}=\mu_{\curly{H}}^{-1}(0)/\curly{H}$.